### v1

# A Visual Case Study of the Training Dynamics in Neural Networks

Abstract
This paper introduces a visual sandbox designed to explore the training dynamics of a small-scale transformer model, with the embedding dimension constrained to $d=2$. This restriction allows for a comprehensive two-dimensional visualization of each layer’s dynamics. Through this approach, we gain insights into training dynamics, circuit transferability, and the causes of loss spikes, including those induced by the high curvature of normalization layers. We propose strategies to mitigate these spikes, demonstrating how good visualization facilitates the design of innovative ideas of practical interest. Additionally, we believe our sandbox could assist theoreticians in assessing essential training dynamics mechanisms and integrating them into future theories. The code is available at https://github.com/facebookresearch/pal.
1 Introduction
The scaling hypothesis in deep learning states that a model’s performance will continue to improve as the size of the model and its training set increase (Hoffmann et al., 2022; Kaplan et al., 2020). This idea has led to a focus on developing very large models, with significant resources dedicated to their training. While this approach has yielded great empirical successes, it requires a substantial computational budget. This makes it challenging to conduct meaningful ablation studies to analyze the importance of different components or factors in a system. As a result, there is still a limited understanding of the training dynamics of these large models and much of the knowledge in this area is based on unpublished tricks and techniques rather than rigorous scientific investigation.
We believe that improving the understanding of this matter is of great importance not only to enhance performance but also to limit the failure cases that naturally occur when scaling the size of the models (Chowdhery et al., 2024; Dehghani et al., 2023; Molybog et al., 2023). This would notably contribute to the reduction of memory and budget costs, along with carbon emissions associated with training large models (Faiz et al., 2024). In this paper, we deliberately choose to focus on a small model to gain a deeper understanding of its training dynamics. In particular, the crux of our approach is to consider a model small enough to visualize it fully. Specifically, we focus on a transformer architecture with an embedding dimension $d=2$, which allows us to plot the dynamics of each layer in a two-dimensional plane. This approach enables us to analyze the model’s internal workings and visualize the underlying mechanisms that drive its behavior, which may be obscured in larger models. We hope that insights gained from small-scale studies could provide a foundation for understanding the complex dynamics at play in larger models, ultimately shedding light on their behavior and performance.
Our main contribution is to provide a visual sandbox which can be of great interest for several reasons. Theoreticians can use it to build intuition on key mechanisms underlying training dynamics, potentially leading to new theorems. Practitioners can use it to test new ideas, such as optimizer modifications. In particular, modifications to the architecture, to the training setting, or to the data can be seamlessly integrated into our pipeline. By providing this sandbox, we aim to facilitate the understanding of the complex dynamics at play in larger models.
Summary of our contributions.
Our contributions are as follows:
  • We develop a visual sandbox to visualize the training dynamics of transformers. Our code is made available to the reviewers and will be open-sourced. It is designed to create videos that help comprehend the model’s internal dynamics comprehensively.
  • Using our visual sandbox, we provide insights into the training dynamics of our model. We classify the types of circuits learned for our problem and illustrate the typical two-phase learning process, consisting of representation learning followed by classifier fitting.
  • We detail the transferability of circuits to showcase the usefulness of curriculum learning and data curation.
  • We investigate loss spikes, suggesting potential strategies for mitigation, which could lead to more stable training processes.
Overall, we hope that our study will guide theorists towards new theorems that capture important aspects beyond the training dynamics of transformers and help practitioners build intuition on potentially promising changes to current training pipelines.
Related Work.
Our work bridges explainability and training dynamics studies.
Opening the Black Box. As neural networks increasingly permeate civil society, concerns about their opacity intensify. This has fueled the ongoing quest to "open the black box." To tackle this issue, researchers have developed various methods, such as extracting meaningful features from neural network activations (Fel et al., 2023), and assessing the impact of perturbations on model inputs (Koh & Liang, 2017; Fel et al., 2021), among others. Recently, the field of mechanistic interpretability has advocated for exposing the internal mechanisms of transformers to provide novel insights into their capabilities (Elhage et al., 2021; Olsson et al., 2022). While some methodologies may apply to large language models (Templeton et al., 2024), precise ablation studies often focus on controlled environments and simplified architectures (Rauker et al., 2023; Charton, 2022; Liu et al., 2022; Nanda et al., 2023; Geva et al., 2023; Meng et al., 2022; Wang et al., 2023; Cabannes et al., 2024a). Our work aligns with this field, as we aim to make the internal behavior of transformers more explicit through carefully selected visualizations in controlled settings. In particular, by utilizing low-dimensional embeddings, we effectively circumvent the curse of dimensionality as defined by (Olah, 2023). Although low-dimensional embeddings make learning the task more challenging, they enable comprehensive visualization of every model component. However, unlike mechanistic interpretability, our goal is to provide insights into the training dynamics of these models, in the spirit of Wortsman et al. (2024).
Training Dynamics of Neural Networks. With the advent of deep learning, new optimization challenges have emerged. Training loss functions are no longer necessarily convex or smooth, the optimization landscape has expanded enormously, and computations are distributed to manage this scale. As the scale of models increases, the absence of theoretical results and closed-form solutions for optimizing neural networks underscores the need to better understand their behavior in practice during training, which is increasingly becoming a computational bottleneck. One approach to address this issue is to employ mathematical abstractions, as seen with neural tangent kernels (NTK) Jacot et al. (2018); Chizat et al. (2020) and mean-field analysis (Mei et al., 2018; Chen et al., 2024), among others (Abbe et al., 2022; Ahn et al., 2024; Cabannes et al., 2024b). Unfortunately, obtaining formal, rigorous results often requires simplifications that deviate from practical implementations. This type of analysis can overlook crucial details that significantly alter the training dynamics.
2 Sandbox Setup
In this section, we describe our controlled setup, focusing on both the data and the architecture.
2.1 Sparse Modular Addition
To probe our model’s decision-making process, we concentrate on a straightforward mathematical task. We draw inspiration from the sparse parity problem (Barak et al., 2023; Abbe et al., 2023) and the modular addition problem (Power et al., 2022; Nanda et al., 2023) to formulate the sparse modular addition problem. This problem is characterized by the following parameters:
  • Input length $L\in\mathbb{N}$,
  • Vocabulary size $p\in\mathbb{N}$,
  • Sparsity index $k\in[L]$, and a set of indices $I\subset[L]$ with cardinality $|I|=k$.
Our default configuration is set to $L=12$, $k=5$, and $p\in\{2,3\}$. Without loss of generality, we assume $I=[k]:=\{1,2,\ldots,k\}$. Inputs are sequences of $L$ tokens $x_{t}$ in $\mathbb{F}_{p}=\mathbb{Z}/p\mathbb{Z}\simeq[p]$, and the corresponding targets are the sum of the first $k$ terms modulo $p$. Formally, we aim to learn a mapping:
$\begin{array}[]{cccc}f^{*}:&\mathbb{F}_{p}^{L}&\to&\mathbb{F}_{p}\\ &x=(x_{1},x_{2},\ldots,x_{L})&\mapsto&y=\sum_{t\in[k]}x_{t}.\end{array}$
This mapping defines deterministic conditional distributions linking input and output data through the formula $p(Y=y|X=x)=\mathbf{1}_{\{f^{*}(x)=y\}}$.
The sparse modular addition task offers the benefits of simplicity while allowing for the observation of a variety of training behaviors.
2.2 One-layer Transformer Architecture
This paper focuses on a specific architecture designed to address our problem.
Sequence Embeddings.
The sparse modular addition problem is inherently discrete. To handle it with a differentiable architecture, we need to embed it in a continuous space ${\mathbb{R}}^{d}$. We first embed each token with both semantic and positional information. Given a learnable token embedding $E:{\mathbb{F}}_{p}\to{\mathbb{R}}^{d}$, and a learnable position embedding $P:[L]\to{\mathbb{R}}^{d}$, a sentence in token space is lifted to a sentence in embedding space through the following embedding operation
$\forall\,t\in[L],\qquad z_{t}:=Z(x_{t},t):=\frac{E(x_{t})+P(t)}{\left\|E(x_{t})+P(t)\right\|}.$ (embedding)
This type of embedding, known as absolute position embedding, incorporates both the semantic meaning of the token and its position within the sequence. The normalization, which projects each element of the sentence onto the sphere, is known as RMS-norm (Zhang & Sennrich, 2019).
Next, an attention mechanism is applied to the sequence of normalized embeddings $(z_{t})$ to aggregate them into a single sentence embedding $\xi\in{\mathbb{R}}^{d}$. It utilizes a query vector $q\in{\mathbb{R}}^{d}$, and a value matrix $V\in{\mathbb{R}}^{d\times d}$. Denoting $z=(z_{t})\in{\mathbb{R}}^{d\times L}$, the sentence embedding can be expressed in matrix form as
$\xi:=(Vz)\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\in{\mathbb{R}}^{d},$ (sentence embedding)
where the softmax operation is defined as a mapping from ${\mathbb{R}}^{L}$ to the simplex $\Delta_{L}$ in ${\mathbb{R}}^{L}$. Specifically, the $s$-th component of $\operatorname{softmax}((a_{t}))$ is proportional to $\exp(a_{s})$, formally expressed as $\operatorname{softmax}((a_{t})_{t\in[L]})=(\exp(a_{s})/\sum_{t\in[L]}\exp(a_{t}))_{s\in[L]}$. Compared to the attention mechanism in Vaswani et al. (2023), we omit both the key and output matrices, which would act as extra parameters that do not increase the expressivity of our model.
Feedforward Neural Network.
Finally, we transform the sentence embeddings using a neural network to cluster them according to the desired output classes. We employ a two-layer multi-layer perceptron (MLP) with pre-norm (Xiong et al., 2020) and residual connection (He et al., 2015). This network is parameterized by $h\in\mathbb{N}$receptors" weights $w_{i}\in\mathbb{R}^{d}$, $h$bias" terms $b_{i}\in\mathbb{R}$, and $h$assemblers" vectors $u_{i}\in\mathbb{R}^{d}$ for $i\in[h]$. It implements the following transformation:
$\zeta:=\xi+\sum_{i\in[h]}\sigma\left(\frac{w_{i}^{\top}\xi}{\|\xi\|}+b_{i}\right)\cdot u_{i}\in\mathbb{R}^{d},$
where $\sigma:\mathbb{R}\to\mathbb{R}$ is a non-linear function, chosen to be $\sigma(x)=x\varphi(x)$ with $\varphi$ being the Gaussian cumulative function. This function $\sigma$ is known as the GELU activation. The final embedding vector $\zeta$ is decoded back to token space based on how it aligns with the respective token embeddings. A softmax layer converts these alignments into a probability vector over the different output classes:
$p_{\zeta}(y)=\operatorname{softmax}((E(y)^{\top}\zeta)_{y\in\mathbb{F}_{p}}).$
Abstracting all the learnable weights into a single parameter $\theta$, our architecture takes as input a sentence $x\in\mathbb{F}_{p}^{L}$ and outputs a probability vector over the classes $p_{\theta}(y|x)$. The parameters of our model are optimized by minimizing the cross-entropy loss defined as:
$\mathcal{L}(\theta):=\mathbb{E}_{(X,Y)\sim p}\left[-\log(p_{\theta}(Y|X))\right],$
where $p$ is a training distribution, typically chosen as the counting measure over some training data. This loss is a proxy for the measure we aim to optimize, which is accuracy, defined as:
$\mathcal{L}(\theta):=\mathbb{E}_{(X,Y)\sim p}\left[\mathbf{1}_{\{Y=\operatorname*{arg\,max}_{y}p_{\theta}(y|X)\}}\right],$
with $p$ denoting a data distribution, typically chosen as the counting measure over some testing data.
Default Configuration.
In our experiments, we trained our networks using $n=2048=2^{11}$ data points, which were sampled uniformly with replacement from the $p^{L}$ possible sentences. We utilized the Adam optimizer with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$ (Kingma & Ba, 2017), and we initialized the network weights using the default schemes provided by PyTorch (Paszke et al., 2019).
2.3 Visualization Tools
The primary motivation of this paper is to embed all computations in dimension $d=2$, enabling us to visualize everything occurring within our model in a relatively straightforward manner. Our codebase is designed to generate videos of the training dynamics, tracking several key aspects. We notably track the following.
Position Embeddings. We visualize $P(t)$ for $t\in[L]$ as a point cloud. In this visualization, "spurious" embeddings, $P(t)$ for $t\notin[k]$, are represented by squares, and "non-spurious" ones, $P(t)$ for $t\in[k]$, by circles. Ideally, the transformer would collapse all spurious (resp. non-spurious) position embeddings into a single point, learning invariance of $y$ to sentence suffixes (resp. to non-spurious token position permutation).
(Normalized) Embeddings. We visualize the normalized version $Z(x,t)$ of $E(x)+P(t)$ for $(x,t)\in\mathbb{F}_{p}\times[L]$. We maintain the same circle and square distinction and use the same color for both $E(x)+P(t)$ and $E(x^{\prime})+P(t)$, for $x^{\prime}\in\mathbb{F}_{p}\setminus\{x\}$. On the normalized plot, we also plot the query $q$ as an arrow in $\mathbb{R}^{2}$, helping us understand how the attention learned where to focus.
Attention Map. A concatenation of attention vectors for different sentences is represented as a vectorized image. This visualization enables us to follow the change in activation patterns, even though it is a pure function of the normed embedding visualization.
Value transform. We visualize $VZ(x,t)$ as a point cloud. This allows us to understand how sequence embeddings are built and how the value matrix may overcome faulty attention patterns.
Sequence embeddings/transforms. We visualize the sequence embeddings $\xi$ (or their transforms $\zeta$) for a set of predefined sentences. These sentences are built by iterating over prefixes $(x_{t})_{t\in[k]}$ and suffixes $(x_{t})_{t\notin[k]}$. Sentences that share the same prefix have the same color. Sentences whose prefixes are equivalent up to a token position permutation have similar colors. Squares, circles, and triangles are used to distinguish between the classes of the sentences. The sequence embeddings visualization is a direct function of the value transformation and the normalized embeddings.
Transform level lines. We visualize the mapping from sentence embedding $\xi$ to their associated learned probabilities $p_{\zeta}(y=1)$. We also plot the sentence embedding on the same plot to better understand the level line changes.
MLP receptors (and assemblers). We visualize the $w_{i}\in\mathbb{R}^{2}$ (and $u_{i}\in\mathbb{R}^{2}$) defining the MLP transform as a point cloud in $\mathbb{R}^{2}$. A consistent color scheme is used to link receptors with the corresponding assemblers.
Loss and accuracy. We visualize the current train/test loss and accuracy. These are classical quantities to track. It is interesting to put them in relation to the other visualizations to better understand the loss spikes, loss plateaus, and phase transitions.
3 Family of Circuits
In this section, we describe idealized circuits designed to solve the sparse modular addition problem and discuss concrete solutions implemented by transformers.
Figure 1: Idealized circuit capturing the invariants of the problem
Figure 2: Faulty attention corrected by value collapse.
Figure 3: Embeddings having only learned some invariants.
Figure 4: Fuzzy construction.
3.1 Idealized Circuits
In the sparse modular addition problem, the output $y$ is invariant to two sets of transformations of the inputs $x=(x_{1},\ldots,x_{L})$.
  • “Permutation invariance:” $y$ does not depend on permutation of non-spurious tokens, i.e., if $\pi\in\mathfrak{S}_{k}$ is a permutation, then $y=f^{*}((x_{i}))=f^{*}((x_{\pi(1)},\ldots,x_{\pi(k)},x_{k+1},\ldots,x_{L}))$.
  • “Suffix invariance:” $y$ does not depend on the suffix $(x_{k+1},\ldots,x_{L})$.
[…]
Figure 4 illustrates this ideal model with our visualization. From left to right, we first plot the two token embeddings $E(x)$ for $x\in[p]$ with $p=2$. We then plot the twelve position embeddings $P(t)$. The spurious (resp. non-spurious) positions are represented with squares (resp. circles), they all collapse to a single point. The normalized embeddings $Z(x,t)$ are plotted in the third frame, annotated with $(x,t)$, with the query vector $q$ represented as a red arrow. This arrow points in the direction of the embedding $Z(x,t)$ for $t\in[k]$, allowing the attention mechanism to focus exclusively on non-spurious tokens. Finally, we plot some sequence embeddings, annotated with $(x_{t})$, and the output of the feedforward transform. The feedforward layer is then able to map each cluster to the appropriate output class. Its output is a probability vector over the classes, which we map to a color according to the RGB color wheel.
3.2 Concrete Realizations
In practice, the model weights learned through gradient descent deviate from the idealized model. We observe several variations of the idealized model, which we categorize as follows:
Faulty Attention, Corrected by Value.
In many instances, we find that the attention scores are not fully concentrated on the first five tokens. However, this faulty attention pattern is compensated by the value matrix, which effectively collapses the embeddings of the spurious tokens that are attended to. This adjustment allows the sequence embeddings to remain invariant to the suffix of the sequence. An example of such a configuration is depicted in Figure 4. We also observe examples where the attention focuses solely on non-spurious tokens that are not, for example, zeros, which still allows the sequence embedding to encode for the prefix sums.
Partially Learned Invariants.
Frequently, we observe that the sequence embeddings have not fully learned all the suffix and prefix invariants, resulting in more than six clusters of sequence embeddings. Specifically, Figure 4 illustrates a scenario where the sequence embeddings are not invariant to the value $x_{6}$, as evidenced by the positions of the blue squares $(0,6)$ and $(1,6)$ on the plot. They also lack invariance to permutations of the token in the first or fifth positions with another of the non-spurious tokens. This results in a sequence embedding that presents more clusters than the idealized model, leading to a greater number of connected decision regions in the feedforward layer.
Fuzzy Constructions.
Occasionally, we encounter fuzzy constructions where the sequence embeddings are clustered according to unconventional patterns that nonetheless generalize to unseen data. Such a construction is presented at the bottom of Figure 4.
To conclude, we observe numerous variations from the idealized construction. The networks discover a variety of weight configurations to effectively solve our task. While these configurations do not exhibit particularly striking geometric structures, they often capture the invariants of our problem.
4 Training Insights
This section discusses several training insights we obtained from our sandbox. It focuses on training dynamics, the transferability of circuits, and loss spikes.
4.1 Training dynamics
Our toolbox enables us to precisely track the training dynamics of the model. We notably observe that the loss curves typically present drops, corresponding to the learning of different parts of the network, hierarchically from the first to the last layers. They are illustrated in Figure 7.
Figure 5: Two-phase learning. Right: Loss profile featuring two significant drops in loss, marked by four red dashed lines at key snapshots. From left to right: (1) During the first snapshot, the sequence embeddings lack any clear structure. (2) They suddenly become clustered after the first loss drops, as seen in the second snapshot. (3) At this point, the MLP already classifies some clusters correctly (third snapshot). (4) A second loss drop occurs as the MLP gets fitted (last snapshot).
Figure 6: Ablation studies regarding test accuracy as a function of batch size, hidden neurons, learning rates, and MLP learning rates discount factor for a single run with 1000 epochs. We ensured consistency in initial weights and batch designs when changing hyperparameters.
Figure 7: This study, akin to Figure 7, yet averaged over $100$ runs, highlights some regularity in the effect of hyperparameters on the resulting test accuracy. The solid color indicates standard deviations.
First loss drop: learning of the sequence embeddings.
The first loss drop coincides with the learning of the sequence embeddings. It really corresponds to a phase change in the dynamics of the weights. Before the first phase change, the weights seem to wander as if trapped in a saddle point, waiting for a clear signal to escape it. At one point, they all move quite rapidly to create a relatively definitive structure for the sequence embeddings. Interestingly, we also notice that when changing the training hyperparameters, the time for this phase change to occur can vary quite a lot, reflecting the highly unpredictable time needed to escape from the saddle point. Figure 7 shows the high variability of the test accuracy after 1000 epochs when slightly varying one hyperparameter (Figure 7 shows more regularity when averaging over the runs). The exit from a saddle point can also be understood via the gradient norms as illustrated in Figures 16 and 16 of Appendix A.2. Our findings resonate with some theoretical arguments found in the deep learning theory literature (Dauphin et al., 2014; Du et al., 2017; Chi et al., 2019), results that we hope our toolbox could help theorists strengthen.
Second loss drop: fitting of the MLP.
The second loss drop is due to the learning of the feedforward network. This change is about fitting the MLP weights to assign the correct classes to the different clusters created during the learning of the sequence embeddings. Interestingly, this second loss drop does not correspond to a clear phase change in the dynamics of the weights. The MLP weights seem to evolve at a continuous speed, although the corresponding decision frontiers change relatively strongly. We notice that this second loss drop appears soon after the first one, if not simultaneously. Once again, we hope that theorists could use our visual toolbox to shed more light on this hierarchical learning of the weights and strengthen previous results on the matter (see, e.g., Abbe et al., 2021).
The influence of initialization.
As we have seen in the previous subsection, the final configuration can vary quite a lot from one run to another. From our manual inspection, the final configuration seems to be highly correlated with the initial weight configurations. We notably see a strong similarity between the attention patterns at the start and at the end of the training, as illustrated by Figure 8. We also notice that modifying the training hyperparameters (batch size, learning rates, etc.) without changing the initial weights does not significantly alter the final weight configurations.
(a) Pretraining with $p=2$.
(b) Finetuning with $p=3$.
Figure 8: Influence of the initialization. Visualization of attention maps: for each sequence on the left, we plot the corresponding attention pattern at the start and the end of training. The final attention maps are correlated with the original one, illustrating that final variations of the original circuit depend on the original weight configuration.
4.2 Transfer Experiments
Among the lessons learned from training very large models is the importance of careful data engineering (Touvron et al., 2023b; Team, 2024). Cabannes et al. (2024a) has highlighted that certain sources of data may facilitate the learning of invariances, while Abbe et al. (2023) discusses how data curation enables models to escape saddle points more quickly. These insights are consistent with the observations we made regarding our problem. In the sparse modular addition problem, the number of unique sequences is equal to $p^{T}$, which increases rapidly with both $p$ and $T$ and makes the problem quite hard to learn. In particular, when setting $T=12$ and limiting the training set to $n=2048$ data points, training for 1000 epochs does not result in any learning for $p>3$.
(a) Mean
(b) Maximum
Figure 9: Accuracies obtained from pretraining only with $p\in[2,10]$ and finetuning with $p\in[3,10]$ starting from $p=2$ for various vocabulary sizes. Left: Averaged accuracy; Right: Maximum accuracy. Finetuned models display better performances than pretrained-only models.
As previously seen in Figure 4, when training with $p=2$, we often find circuits that capture both permutation and suffix invariants. These invariants generalize for any $p\in{\mathbb{N}}$ when $k$ and $T$ are fixed. Consequently, initializing models with these invariants makes learning the sparse modular addition problem much easier. This observation was made after conducting the following experiments: we first trained a model with sequences in ${\mathbb{F}}_{p}$ for $p=2$ over 1000 epochs, before switching the dataset to sequences in ${\mathbb{F}}_{p}$ for $p=3$ for another 1000 epochs. We found that this procedure significantly facilitates the learning of the sparse modular addition problem for $p=3$, which we summarize in Figure 9 where each bar plot is obtained by over $250$ runs. Remarkably, we found that the only models that achieved 100% test accuracy were those that captured both the token and permutation invariances after the first 1000 epochs. Specifically, these were the models that created six sequence embedding clusters, as shown in Figures 4 and 4, rather than those depicted in Figures 4 or 4.
Figure 11 shows the final circuit found in one of our finetuning experiments. The training was initialized with the circuit in Figure 4, after adding a token embedding to encode for $x=2$, resulting in Figure 11. The final embeddings are not that far from the initial one, with the transformer having learned to mainly pay attention to the non-spurious tokens that are not equal to 2. It also pays some attention to 0 and 1 in positions $t=7$ and $t=10$. However, this faulty attention is corrected by the value matrix. Once again, the final configuration seems somewhat close to the initial one, as shown by the attention pattern reported in Figure 8. This is consistent with the observations made in the previous subsections.
Figure 10: Same circuit as in Figure 4 with an additional token embedding for $p=3$.
Figure 11: Circuit learned after $1000$ epochs of finetuning with $p=3$.
Overall, our transfer experiment highlights the transferability of circuits and the usefulness of a curriculum in facilitating the learning of challenging tasks by inducing effective circuits through tasks that are easier to solve. Although our experiments are performed in two stages, we hypothesize that the same type of mechanism can explain the importance of data curation.
4.3 Loss Spikes
One interesting aspect of our sandbox is that it generates loss spikes that we can study quite precisely. Our visual inspection showcases two aspects beyond loss spikes. The high curvature of multi-layer perceptrons with heavy weights, or with numerous small correlated weights, as well as the high curvature of the RMS normalization layer near the origin.
Figure 12: Loss spikes are linked to the high curvature of internal network functions. A small update to an element $x$ can result in a substantial change to its normalized version $x/\left\|x\right\|$, significantly altering the network’s subsequent behavior.
Figure 13: Loss spikes (both left) resulting from a small change from one iteration to another in sequence embeddings that are close to the decision boundaries of the subsequent feedforward layer (both right).
Figure 14: A small change in sequence embeddings (left) can lead to a big change in MLP response (right). This is due to heavy, or small but heavily correlated assemblers (middle).
Figure 12 illustrates that a small modification to an element can result in a disproportionately large change in its normalized version. In theory, gradient points towards directions that would reduce the training loss. However, considering a large step size in these directions could be counterproductive. This is especially true for functions with high curvature, such as the normalization layer $f(x)=x/\left\|x\right\|$ near the origin. At any point $x_{t}$, the gradient descent update rule suggests that one can update $x_{t+1}$ as $x_{t}-\eta_{t}u_{t}$ without changing $f(x_{t})$, where $u_{t}=x_{t}/\|x_{t}\|$ and $\eta_{t}$ is the learning rate. This holds true only if the learning is small enough, $\eta_{t}<\|x_{t}\|$. When $x_{t}$ is close to zero, ensuring $\eta_{t}<\|x_{t}\|$ becomes challenging, particularly if the step size $\eta_{t}$ was predetermined by some scheduler. This behavior is related to the "edge-of-stability" phenomenon highlighted by Cohen et al. (2022), further theoretical insights being provided by Cabannes et al. (2024b). Interestingly, this analysis suggests removing some loss spikes by smoothing out the normalization layer. For example, consider using $f(x)=\sigma(\left\|x\right\|)x/\left\|x\right\|$ where $\sigma$ is a smooth function with $\sigma(0)=0$ and $\sigma([1,\infty))=\{1\}$. This demonstrates the usefulness of our visual sandbox in gaining insights and building intuition, which can then be validated on a larger scale in separate projects.
Another source of loss spikes is illustrated in Figure 14. They are due to the decision boundaries of the feedforward layer being quite close to the sequence embeddings, meaning that a small change in the sequence embedding can lead to a high change in their classification. This is again due to the high curvature of the MLP layer, as illustrated in Figure 14. In particular, the heavy, or the small but heavily correlated, weights in the MLP cause the response $\zeta$ to vary highly as a function of $\zeta$. Once again, one can imagine different ways to regularize these types of loss spikes, with various regularization measures, or by ensuring that the capacity of the MLP is large enough for the MLP to avoid creating these heavy or highly correlated weights.
To conclude, our visual sandbox enabled us to peek into some reasons behind loss spikes in large neural networks and to build intuition on how to circumvent them, giving us ideas for large-scale experiments, which, if successful, could help increase the amount of intelligence learned for a given amount of training compute. Connection to gradient norms and sparsity is discussed in Appendix A.2 and A.3.
Conclusion
This paper aims to advance our understanding of neural network training dynamics through a detailed study of a small-scale transformer model, facilitated by a novel visual sandbox. This tool allows us to observe and analyze the model’s internal mechanisms vividly, providing both theoretical and practical insights that could lead to more efficient training strategies.
Our findings highlight a hierarchical learning process where low-level features are developed first, followed by a more predictable refinement of the model. We also emphasize the impact of initial configurations on the final model outcomes and the seemingly unpredictable nature of some feature learning phase transitions during training. Additionally, we explored the transferability of learned behaviors, which relates to the importance of data curation and curriculum learning in enhancing model performance. We also addressed loss spikes caused by high curvature in the model’s internal functions and proposed potential solutions to mitigate these issues. Looking ahead, future work will aim to apply these insights to larger and more complex models, assess the scalability of observed phenomena, and enhance our visualization tools for higher-dimensional data.
Societal Impact.
This paper aims to understand the complex dynamics at play in neural networks. The insights gained from our research stream could help in designing more efficient and robust neural networks, ultimately aiding in reducing the computational cost and environmental impact of training large models.
Appendix A Appendix
A.1 Reproducibility Statement
We open-source the code at https://github.com/facebookresearch/pal and provide the implementation details in Section 2.2 to ensure reproducibility.
A.2 Understanding Saddle points in Loss Optimization via Gradient Norms
When the model’s training stagnates, the loss plateau for several epochs before it decreases again. This behavior may be understood by the optimizer reaching a saddle point in the loss function w.r.t. the model’s parameter, that is the gradient of the loss is (almost) zero, but the optimization has not reached yet a local minimum. We illustrate this phenomenon in the full-batch and mini-batch setup, respectively in Figures 16 and 16. The connection between gradient norms and learning phases is salient in Figure 16. We can see that loss drops occur in tandem with high gradient norms for each layer. As studied in Section 4.1, these drops correspond to successive learning phases. This is even more salient in the last subfigure of Figure 16 where the pics precisely match the drops. This shows the connection between the exit of a saddle point and high gradient norms. Similarly, in Figure 16, we can see that the first drop in the loss (around epoch $250$ in the first plot) appears when the first increase in gradient norm occurs and the second point of inflection (around epoch $700$) also match an infection point on the gradient norms. It should be noted that this is less salient compared to the full-batch setup.
Figure 15: Connection between loss spikes and gradient norm in full-batch setup. From left to right: Evolution of train and test losses, the corresponding accuracies, the evolution of gradient norms for each layer in log-scale, and the similar evolution in linear scale. We see that the learning phases of Section 4.1 appear in tandem with high gradient norms. This can be seen in the last subfigure where the three pics correspond to the loss drops and their corresponding plateaus
Figure 16: Mini-batch setup. This study, akin to Figure 16, makes the connection between gradient norm and loss profile even more salient for our problem. The first loss drop appears simultaneously with the first increase in gradient norms. Loss spikes occur in tandem with spikes in gradient norms.
However, this setup enables us to study another training behavior mentioned in Sectin 4.3. Indeed, we observe in Figure 16 that the loss spikes appear in tandem with high gradient norms, indicating that when a too-large step size deviates the model from its current small loss region, it is taken back to where it was with large updates. It has been shown in the literature (Ilbert et al., 2024; Foret et al., 2021; Zhang et al., 2024) that studying the gradient and the hessian of the loss could provide valuable insights on the neural network optimization both from a theoretical and experimental point of view. We believe using our visual sandbox could help in future work on such an investigation.
A.3 Impact of Activation Sparsity on Models Performance
Sparsity is a phenomenon of interest in many fields such as signal processing, neuroscience, and machine learning (Mairal et al., 2009; Chen et al., 1998; Barth & Poulet, 2012). Recent studies focused on the sparsity in deep neural network activations. In particular, Li et al. (2023) showed that trained transformers have sparse activations and concluded that it was caused by the training dynamics rather than by a compact representation of the training data as commonly thought in computer vision and NLP. Mirzadeh et al. (2024) observed a similar phenomenon and showed how to leverage sparsity to reduce the inference cost of large language models. Inspired by this line of work, we study the activation sparsity of our model from a performance viewpoint. It should be noted that those works study deep transformers and identified that the sparsity increases with the depth while we only consider a one-layer transformer.
Following the framework from Li et al. (2023), we recall that the activation sparsity corresponds to the percentage of non-zero entries of the feed-forward activation map. Without loss of generality, the feed-forward block is an MLP with weights $W_{1},W_{2}$ and a non-linear activation $\sigma$ that outputs for any input $x$ a vector $z=W_{1}\sigma(W_{2}x)$. Formally, the activation sparsity is the percentage of non-zero entries in $\sigma(W_{2}x)$ and takes values in $[0,1]$. In the classical setting with ReLU activation (Fukushima, 1969), this is equivalent to computing the percentage of non-negative neurons before the activation. However, some activations do not have non-negative outputs. This is the case of the SiLU (Elfwing et al., 2018), used in Llama models (Touvron et al., 2023a), and of the GeLU Hendrycks & Gimpel (2023) used in Falcon (Almazrouei et al., 2023), PaLM (Chowdhery et al., 2024), and in our transformer implementation. Instead of replacing such activations by a ReLU (Mirzadeh et al., 2024; Li et al., 2023), we compute a smoothed sparsity [^1] This is akin to using the $\ell_{1}$-norm, respectively the nuclear norm, instead of the $\ell_{0}$ quasi-norm, respectively the rank (Ilbert et al., 2024; Gribonval & Nielsen, 2003). as the percentage of entries with an absolute value lower than $\varepsilon>0$. A sparsity of $1$ means that all entries are $\varepsilon$-close to 00 (i.e., sparse activations) and a sparsity of 00 means that all entries are at least $\varepsilon$-away from 00 (i.e., dense activations).
Figure 17: Connection between performance and sparsity. We display the evolution of the activation sparsity of $20$ trained models with $\varepsilon\in[10^{-5},10^{2}]$. Left: Successful models (i.e., with test accuracy above $0.9$) in blue have less sparse activation than failed models in red. Right: The color indicates the models’ test accuracy (the lighter, the better). The performance increases as the activation sparsity decreases.
To better understand the impact of sparsity, we train $20$ independent models and display in Figure 17 their sparsity after training for $\varepsilon\in[10^{-5},10^{2}]$ (the range is chosen such that the sparsity reaches its extremal values 00 and $1$). Given the task’s difficulty, achieving an accuracy above $0.9$ is a success; otherwise, it is a failure. On the left, we plot successful models in blue and failed ones in red. We observe a striking separation between successful and failed training. In the admissible range $[10^{-3},10]$, successful models tend to have less sparse activation than failed ones. To further study this phenomenon, we plot in the right subplot of Figure 17 the evolution of the sparsity with $\varepsilon$, and here, the color indicates the models’ test accuracy (the lighter the color, the better the model). We can see that the sparsity decreases as the performance increases. This explains the sharp transition between failure and success in the left subplot. This experiment seems to indicate that, contrary to images and textual data (Li et al., 2023; Mirzadeh et al., 2024), the sparse modular addition problem needs the involvement of many neurons during inference, and hence requires non-sparse activations.

### v2

@@
-A Visual Case Study of the Training Dynamics in Neural Networks
+Clustering Head: A Visual Case Study of the Training Dynamics in Transformers
 Abstract
-This paper introduces a visual sandbox designed to explore the training dynamics of a small-scale transformer model, with the embedding dimension constrained to $d=2$.
-This restriction allows for a comprehensive two-dimensional visualization of each layer’s dynamics.
-Through this approach, we gain insights into training dynamics, circuit transferability, and the causes of loss spikes, including those induced by the high curvature of normalization layers.
-We propose strategies to mitigate these spikes, demonstrating how good visualization facilitates the design of innovative ideas of practical interest.
-Additionally, we believe our sandbox could assist theoreticians in assessing essential training dynamics mechanisms and integrating them into future theories.
-The code is available at https://github.com/facebookresearch/pal.
+This paper introduces the sparse modular addition task and examines how transformers learn it.
+We focus on transformers with embeddings in ${\mathbb{R}}^{2}$ and introduce a visual sandbox that provides comprehensive visualizations of each layer throughout the training process.
+We reveal a type of circuit, called "clustering heads," which learns the problem’s invariants.
+We analyze the training dynamics of these circuits, highlighting two-stage learning, loss spikes due to high curvature or normalization layers, and the effects of initialization and curriculum learning.
+Mechanistic Interpretability, Circuits, Clustering Heads, Transformers
+[…]
+[…]
 1 Introduction
@@
 As a result, there is still a limited understanding of the training dynamics of these large models and much of the knowledge in this area is based on unpublished tricks and techniques rather than rigorous scientific investigation.
-We believe that improving the understanding of this matter is of great importance not only to enhance performance but also to limit the failure cases that naturally occur when scaling the size of the models (Chowdhery et al., 2024; Dehghani et al., 2023; Molybog et al., 2023).
-This would notably contribute to the reduction of memory and budget costs, along with carbon emissions associated with training large models (Faiz et al., 2024).
-In this paper, we deliberately choose to focus on a small model to gain a deeper understanding of its training dynamics.
+In this paper, we have deliberately chosen to focus on a small model to gain a deeper understanding of its training dynamics.
 In particular, the crux of our approach is to consider a model small enough to visualize it fully.
-Specifically, we focus on a transformer architecture with an embedding dimension $d=2$, which allows us to plot the dynamics of each layer in a two-dimensional plane.
+Specifically, we first focus on a transformer architecture with an embedding dimension $d=2$, which allows us to plot the dynamics of each layer in a two-dimensional plane.
+We then generalize our insights to larger embedding dimensions.
 This approach enables us to analyze the model’s internal workings and visualize the underlying mechanisms that drive its behavior, which may be obscured in larger models.
 We hope that insights gained from small-scale studies could provide a foundation for understanding the complex dynamics at play in larger models, ultimately shedding light on their behavior and performance.
-Our main contribution is to provide a visual sandbox which can be of great interest for several reasons.
-Theoreticians can use it to build intuition on key mechanisms underlying training dynamics, potentially leading to new theorems.
-Practitioners can use it to test new ideas, such as optimizer modifications.
-In particular, modifications to the architecture, to the training setting, or to the data can be seamlessly integrated into our pipeline.
-By providing this sandbox, we aim to facilitate the understanding of the complex dynamics at play in larger models.
-Summary of our contributions.
-Our contributions are as follows:
-  • We develop a visual sandbox to visualize the training dynamics of transformers.
-Our code is made available to the reviewers and will be open-sourced.
-It is designed to create videos that help comprehend the model’s internal dynamics comprehensively.
-  • Using our visual sandbox, we provide insights into the training dynamics of our model.
-We classify the types of circuits learned for our problem and illustrate the typical two-phase learning process, consisting of representation learning followed by classifier fitting.
-  • We detail the transferability of circuits to showcase the usefulness of curriculum learning and data curation.
+Our contributions can be summarized as follows:
+  • We introduce the sparse modular addition problem and develop a visual sandbox to visualize the training dynamics of transformers.
+Our code is available at https://github.com/facebookresearch/pal.
+It is designed to create videos that help thoroughly comprehend the model’s internal dynamics.
+  • Using our visual sandbox, we unveil Clustering Heads, circuits that capture the invariants of the task at hand, and we classify concrete realizations that can be seen as variations of these idealized circuits.
+  • We study the training dynamics behind clustering heads, illustrating a two-phase learning process consisting of representation learning followed by classifier fitting.
 We investigate loss spikes, suggesting potential strategies for mitigation, which could lead to more stable training processes.
+We also peek into the transferability of circuits to showcase the usefulness of curriculum learning and data curation.
+  • Our theoretical insights show that clustering heads can be learned via gradient descent and provide an understanding of the two-stage learning observed in practice.
 Overall, we hope that our study will guide theorists towards new theorems that capture important aspects beyond the training dynamics of transformers and help practitioners build intuition on potentially promising changes to current training pipelines.
@@
 While some methodologies may apply to large language models (Templeton et al., 2024), precise ablation studies often focus on controlled environments and simplified architectures (Rauker et al., 2023; Charton, 2022; Liu et al., 2022; Nanda et al., 2023; Geva et al., 2023; Meng et al., 2022; Wang et al., 2023; Cabannes et al., 2024a).
-Our work aligns with this field, as we aim to make the internal behavior of transformers more explicit through carefully selected visualizations in controlled settings.
-In particular, by utilizing low-dimensional embeddings, we effectively circumvent the curse of dimensionality as defined by (Olah, 2023).
+Our work aligns with this field, since we aim to make the internal behavior of transformers more explicit through carefully selected visualizations in controlled settings.
+In particular, by utilizing low-dimensional embeddings, we effectively circumvent the curse of dimensionality as defined by Olah (2023).
 Although low-dimensional embeddings make learning the task more challenging, they enable comprehensive visualization of every model component.
@@
 Training loss functions are no longer necessarily convex or smooth, the optimization landscape has expanded enormously, and computations are distributed to manage this scale.
-As the scale of models increases, the absence of theoretical results and closed-form solutions for optimizing neural networks underscores the need to better understand their behavior in practice during training, which is increasingly becoming a computational bottleneck.
-One approach to address this issue is to employ mathematical abstractions, as seen with neural tangent kernels (NTK) Jacot et al.
-(2018); Chizat et al.
-(2020) and mean-field analysis (Mei et al., 2018; Chen et al., 2024), among others (Abbe et al., 2022; Ahn et al., 2024; Cabannes et al., 2024b).
-Unfortunately, obtaining formal, rigorous results often requires simplifications that deviate from practical implementations.
+As the scale of models increases, the absence of theoretical results and closed-form solutions for optimizing neural networks underscores the need to better understand their behavior in practice during training, which is rapidly becoming a computational bottleneck.
+One approach to address this issue is to employ mathematical abstractions, as seen with neural tangent kernels (NTK) (Jacot et al., 2018; Chizat et al., 2020) and mean-field analysis (Mei et al., 2018; Chen et al., 2024), among others (Abbe et al., 2022; Ahn et al., 2024; Cabannes et al., 2024c).
+Unfortunately, obtaining formal, rigorous results often requires simplifications that deviate from practical implementations (Bietti et al., 2023; Cabannes et al., 2024c; Boix-Adsera et al., 2023; Tian et al., 2023; Nichani et al., 2024; Edelman et al., 2024; Cabannes et al., 2024b; Ahn et al., 2023; Geshkovski et al., 2023; Jelassi et al., 2022; Mahankali et al., 2024).
 This type of analysis can overlook crucial details that significantly alter the training dynamics.
-2 Sandbox Setup
-In this section, we describe our controlled setup, focusing on both the data and the architecture.
+In our work, we try to remain as faithful as possible to the original Transformer architecture to ensure that our insights bear real-world implications.
+2 Controlled Setup
+In this section, we describe our controlled setup, focusing on the data and architecture, along with our visualization tools to better understand the training dynamics.
 2.1 Sparse Modular Addition
-To probe our model’s decision-making process, we concentrate on a straightforward mathematical task.
+In terms of data, we concentrate on a straightforward mathematical task.
 We draw inspiration from the sparse parity problem (Barak et al., 2023; Abbe et al., 2023) and the modular addition problem (Power et al., 2022; Nanda et al., 2023) to formulate the sparse modular addition problem.
@@
 Formally, we aim to learn a mapping:
-$\begin{array}[]{cccc}f^{*}:&\mathbb{F}_{p}^{L}&\to&\mathbb{F}_{p}\\ &x=(x_{1},x_{2},\ldots,x_{L})&\mapsto&y=\sum_{t\in[k]}x_{t}.\end{array}$
+$\begin{array}[]{cccc}f^{*}:&(\mathbb{F}_{p})^{L}&\to&\mathbb{F}_{p}\\ &x=(x_{1},\ldots,x_{L})&\mapsto&y=\sum_{t\in[k]}x_{t}.\end{array}$
 This mapping defines deterministic conditional distributions linking input and output data through the formula $p(Y=y|X=x)=\mathbf{1}_{\{f^{*}(x)=y\}}$.
-The sparse modular addition task offers the benefits of simplicity while allowing for the observation of a variety of training behaviors.
-2.2 One-layer Transformer Architecture
-This paper focuses on a specific architecture designed to address our problem.
+2.2 Transformer Architecture
+Our study is made with a transformer model that adapts the original implementation by Vaswani et al. (2023) to our task.
 Sequence embeddings.
@@
 We first embed each token with both semantic and positional information.
-Given a learnable token embedding $E:{\mathbb{F}}_{p}\to{\mathbb{R}}^{d}$, and a learnable position embedding $P:[L]\to{\mathbb{R}}^{d}$, a sentence in token space is lifted to a sentence in embedding space through the following embedding operation
-$\forall\,t\in[L],\qquad z_{t}:=Z(x_{t},t):=\frac{E(x_{t})+P(t)}{\left\|E(x_{t})+P(t)\right\|}.$ (embedding)
-This type of embedding, known as absolute position embedding, incorporates both the semantic meaning of the token and its position within the sequence.
+Given a learnable token embedding $E:{\mathbb{F}}_{p}\to{\mathbb{R}}^{d}$, and a learnable position embedding $P:[L]\to{\mathbb{R}}^{d}$, a sentence in token space is lifted to a sentence in embedding space through the following embedding operation that applies on each $x_{t}$ for $t\in[L]$
+$z_{t}:=Z(x_{t},t):=\frac{E(x_{t})+P(t)}{\left\|E(x_{t})+P(t)\right\|}.$
+This type of embedding, known as absolute position embedding, incorporates both the token’s semantic meaning and its position within the sequence.
 The normalization, which projects each element of the sentence onto the sphere, is known as RMS-norm (Zhang & Sennrich, 2019).
-Next, an attention mechanism is applied to the sequence of normalized embeddings $(z_{t})$ to aggregate them into a single sentence embedding $\xi\in{\mathbb{R}}^{d}$.
+Next, a cross-attention mechanism is applied to the sequence of normalized embeddings $(z_{t})$ to aggregate them into a single sentence embedding $\xi\in{\mathbb{R}}^{d}$.
 It utilizes a query vector $q\in{\mathbb{R}}^{d}$, and a value matrix $V\in{\mathbb{R}}^{d\times d}$.
@@
 Compared to the attention mechanism in Vaswani et al.
-(2023), we omit both the key and output matrices, which would act as extra parameters that do not increase the expressivity of our model.
+(2023), we omit both the key and output matrices, which would act as extra parameters that do not increase the expressivity of our model since $z$ can be set to anything thanks to $E$ and $P$.
 Feedforward neural network.
@@
 It implements the following transformation:
-$\zeta:=\xi+\sum_{i\in[h]}\sigma\left(\frac{w_{i}^{\top}\xi}{\|\xi\|}+b_{i}\right)\cdot u_{i}\in\mathbb{R}^{d},$
+$\psi:=\xi+\sum_{i\in[h]}u_{i}\cdot\sigma\left(\frac{w_{i}^{\top}\xi}{\|\xi\|}\right)\in\mathbb{R}^{d},$
 where $\sigma:\mathbb{R}\to\mathbb{R}$ is a non-linear function, chosen to be $\sigma(x)=x\varphi(x)$ with $\varphi$ being the Gaussian cumulative function.
-This function $\sigma$ is known as the GELU activation.
-The final embedding vector $\zeta$ is decoded back to token space based on how it aligns with the respective token embeddings.
-A softmax layer converts these alignments into a probability vector over the different output classes:
-$p_{\zeta}(y)=\operatorname{softmax}((E(y)^{\top}\zeta)_{y\in\mathbb{F}_{p}}).$
-Abstracting all the learnable weights into a single parameter $\theta$, our architecture takes as input a sentence $x\in\mathbb{F}_{p}^{L}$ and outputs a probability vector over the classes $p_{\theta}(y|x)$.
+The function $\sigma$ is known as the GELU activation.
+The final embedding vector $\zeta$ is decoded back to token space based on how it aligns with the respective token embeddings which amounts to using weight-tying in the implementation (Press & Wolf, 2017).
+More specifically, the model’s logits write
+$\zeta=(E(v)^{\top}\psi)_{v\in\mathbb{F}_{p}}\in\mathbb{R}^{p}.$
+A softmax layer converts these alignments into a probability vector $\mu=\operatorname{softmax}(\zeta)\in\Delta_{p}$ over the different output classes.
+Abstracting all the learnable weights into a single parameter $\theta$, our architecture takes as input a sentence $x\in(\mathbb{F}_{p})^{L}$ and outputs the probability $\hat{p}_{\theta}(y|x)=\mu_{y}$.
 The parameters of our model are optimized by minimizing the cross-entropy loss defined as:
-$\mathcal{L}(\theta):=\mathbb{E}_{(X,Y)\sim p}\left[-\log(p_{\theta}(Y|X))\right],$
-where $p$ is a training distribution, typically chosen as the counting measure over some training data. This loss is a proxy for the measure we aim to optimize, which is accuracy, defined as:
-$\mathcal{L}(\theta):=\mathbb{E}_{(X,Y)\sim p}\left[\mathbf{1}_{\{Y=\operatorname*{arg\,max}_{y}p_{\theta}(y|X)\}}\right],$
-with $p$ denoting a data distribution, typically chosen as the counting measure over some testing data.
-Default Configuration.
-In our experiments, we trained our networks using $n=2048=2^{11}$ data points, which were sampled uniformly with replacement from the $p^{L}$ possible sentences.
-We utilized the Adam optimizer with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$ (Kingma & Ba, 2017), and we initialized the network weights using the default schemes provided by PyTorch (Paszke et al., 2019).
-2.3 Visualization Tools
-The primary motivation of this paper is to embed all computations in dimension $d=2$, enabling us to visualize everything occurring within our model in a relatively straightforward manner.
+$\mathcal{L}(\theta):=\mathbb{E}_{(x,y)\sim p}\left[-\log(\hat{p}_{\theta}(y|x))\right],$
+where $p$ is a training distribution, typically chosen as the counting measure over some training data.
+This loss is a proxy for the measure we aim to optimize, i.e., the accuracy [^1] In practice, the cross-entropy is a differentiable proxy of the error, i.e., the $0-1$ loss which is the counterpart of the accuracy., defined as
+$\mathcal{A}(\theta):=\mathbb{E}_{(x,y)\sim q}\left[\mathbbm{1}_{\{y=\operatorname*{arg\,max}_{v}\hat{p}_{\theta}(v|x)\}}\right],$
+with $q$ denoting a data distribution, typically chosen as the counting measure over some testing data. The implementation details are given in Appendix A.
+2.3 Visual Sandbox
+To visualize everything occurring within our model along the training, we embed all computations in the plan by taking $d=2$.
+Note that we also validate our insights in larger models with embeddings size $d>2$ in Appendix B.
 Our codebase is designed to generate videos of the training dynamics, tracking several key aspects.
@@
 Value transform. We visualize $VZ(x,t)$ as a point cloud. This allows us to understand how sequence embeddings are built and how the value matrix may overcome faulty attention patterns.
+Figure 1: Clustering Head. Implementation of an idealized circuit that captures the invariants of the problem.
 Sequence embeddings/transforms.
@@
 Transform level lines.
-We visualize the mapping from sentence embedding $\xi$ to their associated learned probabilities $p_{\zeta}(y=1)$.
+We visualize the mapping from sentence embedding $\xi$ to their associated learned probabilities $\hat{p}_{\theta}(y|x)$.
 We also plot the sentence embedding on the same plot to better understand the level line changes.
@@
 It is interesting to put them in relation to the other visualizations to better understand the loss spikes, loss plateaus, and phase transitions.
-3 Family of Circuits
-In this section, we describe idealized circuits designed to solve the sparse modular addition problem and discuss concrete solutions implemented by transformers.
-Figure 1: Idealized circuit capturing the invariants of the problem
-Figure 2: Faulty attention corrected by value collapse.
-Figure 3: Embeddings having only learned some invariants.
-Figure 4: Fuzzy construction.
-3.1 Idealized Circuits
+3 Clustering Head: A Circuit Perspective
+In this section, we adopt a circuit perspective and describe an efficient pathway for solving the sparse modular addition problem.
+In particular, we describe an idealized “clustering head” circuit, as well as variants observed in practice.
+3.1 Idealized Circuit
 In the sparse modular addition problem, the output $y$ is invariant to two sets of transformations of the inputs $x=(x_{1},\ldots,x_{L})$.
-  • “Permutation invariance:” $y$ does not depend on permutation of non-spurious tokens, i.e., if $\pi\in\mathfrak{S}_{k}$ is a permutation, then $y=f^{*}((x_{i}))=f^{*}((x_{\pi(1)},\ldots,x_{\pi(k)},x_{k+1},\ldots,x_{L}))$.
+  • “Permutation invariance:” $y$ does not depend on permutation of non-spurious tokens, i.e. for $\pi\in\mathfrak{S}_{k}$, $f^{*}((x_{i}))=f^{*}((x_{\pi(1)},\ldots,x_{\pi(k)},x_{k+1},\ldots,x_{L}))$.
   • “Suffix invariance:” $y$ does not depend on the suffix $(x_{k+1},\ldots,x_{L})$.
-[…]
-Figure 4 illustrates this ideal model with our visualization.
+These two sets of invariants can be easily enforced by the embedding layer.
+Any architecture where the position embeddings satisfy $p_{t}=p_{1}$ for $t\in[k]$, will be permutation invariant, meaning that its output will be invariant to permutations of the non-spurious tokens.
+Similarly, suffix invariance can be enforced by ensuring that the query vector primarily aligns with $Z(x,t)$ for $t\in[k]$, allowing the sequence embedding $\xi$ to be invariant to the sequence suffix.
+Such a construction would yield (${k+p-1\choose k}$ clusters of sequence embeddings, [^2] This number corresponds to the number of ways to split $k$ into $p$ buckets, which is also the number of stars and bars configurations with $k$ stars and $p-1$ bars.
+which the feedforward layer could scatter into as many decision regions to map each sequence embedding $\xi$ to the correct output class $y\in[p]$.
+The pathway that realizes such operations is called a Clustering Head.
+Figure 1 illustrates this ideal model using our visual sandbox.
 From left to right, we first plot the two token embeddings $E(x)$ for $x\in[p]$ with $p=2$.
@@
 This arrow points in the direction of the embedding $Z(x,t)$ for $t\in[k]$, allowing the attention mechanism to focus exclusively on non-spurious tokens.
-Finally, we plot some sequence embeddings, annotated with $(x_{t})$, and the output of the feedforward transform.
-The feedforward layer is then able to map each cluster to the appropriate output class.
+Finally, we plot the sequence embeddings $\xi$ of some $(x_{t})$, and the output of the feedforward transform.
+The embeddings are clustered and respect the invariance, and the feedforward layer can map each cluster to the appropriate output class.
 Its output is a probability vector over the classes, which we map to a color according to the RGB color wheel.
-3.2 Concrete Realizations
-In practice, the model weights learned through gradient descent deviate from the idealized model. We observe several variations of the idealized model, which we categorize as follows:
-Faulty Attention, Corrected by Value.
-In many instances, we find that the attention scores are not fully concentrated on the first five tokens.
-However, this faulty attention pattern is compensated by the value matrix, which effectively collapses the embeddings of the spurious tokens that are attended to.
-This adjustment allows the sequence embeddings to remain invariant to the suffix of the sequence.
-An example of such a configuration is depicted in Figure 4.
-We also observe examples where the attention focuses solely on non-spurious tokens that are not, for example, zeros, which still allows the sequence embedding to encode for the prefix sums.
-Partially Learned Invariants.
-Frequently, we observe that the sequence embeddings have not fully learned all the suffix and prefix invariants, resulting in more than six clusters of sequence embeddings.
-Specifically, Figure 4 illustrates a scenario where the sequence embeddings are not invariant to the value $x_{6}$, as evidenced by the positions of the blue squares $(0,6)$ and $(1,6)$ on the plot.
-They also lack invariance to permutations of the token in the first or fifth positions with another of the non-spurious tokens.
-This results in a sequence embedding that presents more clusters than the idealized model, leading to a greater number of connected decision regions in the feedforward layer.
-Fuzzy Constructions.
-Occasionally, we encounter fuzzy constructions where the sequence embeddings are clustered according to unconventional patterns that nonetheless generalize to unseen data.
-Such a construction is presented at the bottom of Figure 4.
-To conclude, we observe numerous variations from the idealized construction.
-The networks discover a variety of weight configurations to effectively solve our task.
-While these configurations do not exhibit particularly striking geometric structures, they often capture the invariants of our problem.
-4 Training Insights
-This section discusses several training insights we obtained from our sandbox. It focuses on training dynamics, the transferability of circuits, and loss spikes.
-4.1 Training dynamics
-Our toolbox enables us to precisely track the training dynamics of the model.
-We notably observe that the loss curves typically present drops, corresponding to the learning of different parts of the network, hierarchically from the first to the last layers.
-They are illustrated in Figure 7.
 Figure 2: Two-phase learning.
@@
 (4) A second loss drop occurs as the MLP gets fitted (last snapshot).
-Figure 6: Ablation studies regarding test accuracy as a function of batch size, hidden neurons, learning rates, and MLP learning rates discount factor for a single run with 1000 epochs.
-We ensured consistency in initial weights and batch designs when changing hyperparameters.
-Figure 7: This study, akin to Figure 7, yet averaged over $100$ runs, highlights some regularity in the effect of hyperparameters on the resulting test accuracy.
-The solid color indicates standard deviations.
+3.2 Concrete Realizations
+We expect clustering heads to naturally appear during training given that they are an efficient and simple way to solve the task.
+In practice, the sequence embeddings $\xi$ learned through gradient descent can be grouped in similar but different ways, which leads to several variations of the idealized head.
+These variations are categorized below.
+Faulty attention corrected by value.
+In many instances, we observe that the attention scores are not fully concentrated on the first five tokens but are compensated by the value matrix, which effectively collapses the embeddings of the spurious tokens that are attended to.
+This allows the sequence embeddings to remain invariant to the suffix of the sequence as can be seen in Fig.
+13.
+Partially learned invariants.
+Frequently, the sequence embeddings have not fully learned all the suffix and prefix invariants, resulting in more than six clusters of sequence embeddings [^3] This corresponds to (${k+p-1\choose k}$ for $p=2,k=5$..
+Specifically, Figure 13 illustrates a scenario where the sequence embeddings are not invariant to the value $x_{6}$, as evidenced by the positions of the blue squares $(0,6)$ and $(1,6)$ on the plot.
+They also lack invariance to permutations of the token in the first or fifth positions with another of the non-spurious tokens.
+This results in a sequence embedding that presents more clusters than the idealized model, leading to a greater number of connected decision regions in the feedforward layer.
+Fuzzy constructions.
+Occasionally, we encounter fuzzy constructions where the sequence embeddings are clustered according to unconventional patterns that nonetheless generalize to unseen data.
+Such a construction is presented in Figure 13.
+4 Learning Clustering Heads
+This section investigates the training dynamics of solving the sparse modular addition problem.
+We begin by making empirical observations: we observe that transformers typically learn clustering heads through a two-stage learning process and that loss spikes can occur during this process.
+We then discuss theoretical insights that can be derived without any model simplification assumptions.
+4.1 Two-Phase, One Loss: Escaping Saddle Points
+A first experimental observation is the typical two-stage learning process that transformers undergo in our setting: first clustering sequence embeddings, before fitting a classifier on top.
+In particular, the loss curves typically present two drops, corresponding to the learning of different parts of the network, hierarchically from the first to the last parts of the network.
+This is illustrated in Figure 2.
+Our findings resonate with theoretical arguments in the deep learning theory literature on saddle point (Dauphin et al., 2014; Du et al., 2017; Chi et al., 2019a) and hierarchical learning (Abbe et al., 2021).
 First loss drop: learning of the sequence embeddings.
 The first loss drop coincides with the learning of the sequence embeddings.
-It really corresponds to a phase change in the dynamics of the weights.
+It corresponds to a phase change in the dynamics of the weights.
 Before the first phase change, the weights seem to wander as if trapped in a saddle point, waiting for a clear signal to escape it.
 At one point, they all move quite rapidly to create a relatively definitive structure for the sequence embeddings.
-Interestingly, we also notice that when changing the training hyperparameters, the time for this phase change to occur can vary quite a lot, reflecting the highly unpredictable time needed to escape from the saddle point.
-Figure 7 shows the high variability of the test accuracy after 1000 epochs when slightly varying one hyperparameter (Figure 7 shows more regularity when averaging over the runs).
-The exit from a saddle point can also be understood via the gradient norms as illustrated in Figures 16 and 16 of Appendix A.2.
-Our findings resonate with some theoretical arguments found in the deep learning theory literature (Dauphin et al., 2014; Du et al., 2017; Chi et al., 2019), results that we hope our toolbox could help theorists strengthen.
+Interestingly, we also notice that the time it takes for this phase change to occur when changing the training hyperparameters can vary quite a lot, reflecting the highly unpredictable time needed to escape from the saddle point.
+This is illustrated in Section A.4 where Figure 16(a) shows the high variability of the test accuracy after 1000 epochs when slightly varying one hyperparameter (Figure 16(b) shows the same phenomenon with more regularity when averaging over the runs).
+The exit from a saddle point can also be understood via the gradient norms, as discussed at the end of this section.
+Figure 3: Connection to saddle points.
+From left to right: Evolution of train and test losses, the corresponding accuracies, the evolution of gradient norms for each layer in log-scale, and the similar evolution in linear scale in full-batch.
+We see that the learning phases of Section 4.1 appear in tandem with high gradient norms.
+This can be seen in the last subfigure where the three pics correspond to the loss drops and their corresponding plateaus.
+Figure 4: High-curvature feed-forward.
+Loss spikes (both left) resulting from a small change from one iteration to another in sequence embeddings that are close to the decision boundaries of the subsequent feedforward layer (both right).
 Second loss drop: fitting of the MLP.
@@
 We notice that this second loss drop appears soon after the first one, if not simultaneously.
-Once again, we hope that theorists could use our visual toolbox to shed more light on this hierarchical learning of the weights and strengthen previous results on the matter (see, e.g., Abbe et al., 2021).
-The influence of initialization.
-As we have seen in the previous subsection, the final configuration can vary quite a lot from one run to another.
-From our manual inspection, the final configuration seems to be highly correlated with the initial weight configurations.
-We notably see a strong similarity between the attention patterns at the start and at the end of the training, as illustrated by Figure 8.
-We also notice that modifying the training hyperparameters (batch size, learning rates, etc.) without changing the initial weights does not significantly alter the final weight configurations.
-(a) Pretraining with $p=2$.
-(b) Finetuning with $p=3$.
-Figure 8: Influence of the initialization.
-Visualization of attention maps: for each sequence on the left, we plot the corresponding attention pattern at the start and the end of training.
-The final attention maps are correlated with the original one, illustrating that final variations of the original circuit depend on the original weight configuration.
-4.2 Transfer Experiments
-Among the lessons learned from training very large models is the importance of careful data engineering (Touvron et al., 2023b; Team, 2024).
-Cabannes et al.
-(2024a) has highlighted that certain sources of data may facilitate the learning of invariances, while Abbe et al.
-(2023) discusses how data curation enables models to escape saddle points more quickly.
-These insights are consistent with the observations we made regarding our problem.
-In the sparse modular addition problem, the number of unique sequences is equal to $p^{T}$, which increases rapidly with both $p$ and $T$ and makes the problem quite hard to learn.
-In particular, when setting $T=12$ and limiting the training set to $n=2048$ data points, training for 1000 epochs does not result in any learning for $p>3$.
-(a) Mean
-(b) Maximum
-Figure 9: Accuracies obtained from pretraining only with $p\in[2,10]$ and finetuning with $p\in[3,10]$ starting from $p=2$ for various vocabulary sizes.
-Left: Averaged accuracy; Right: Maximum accuracy.
-Finetuned models display better performances than pretrained-only models.
-As previously seen in Figure 4, when training with $p=2$, we often find circuits that capture both permutation and suffix invariants.
-These invariants generalize for any $p\in{\mathbb{N}}$ when $k$ and $T$ are fixed.
-Consequently, initializing models with these invariants makes learning the sparse modular addition problem much easier.
-This observation was made after conducting the following experiments: we first trained a model with sequences in ${\mathbb{F}}_{p}$ for $p=2$ over 1000 epochs, before switching the dataset to sequences in ${\mathbb{F}}_{p}$ for $p=3$ for another 1000 epochs.
-We found that this procedure significantly facilitates the learning of the sparse modular addition problem for $p=3$, which we summarize in Figure 9 where each bar plot is obtained by over $250$ runs.
-Remarkably, we found that the only models that achieved 100% test accuracy were those that captured both the token and permutation invariances after the first 1000 epochs.
-Specifically, these were the models that created six sequence embedding clusters, as shown in Figures 4 and 4, rather than those depicted in Figures 4 or 4.
-Figure 11 shows the final circuit found in one of our finetuning experiments.
-The training was initialized with the circuit in Figure 4, after adding a token embedding to encode for $x=2$, resulting in Figure 11.
-The final embeddings are not that far from the initial one, with the transformer having learned to mainly pay attention to the non-spurious tokens that are not equal to 2.
-It also pays some attention to 0 and 1 in positions $t=7$ and $t=10$.
-However, this faulty attention is corrected by the value matrix.
-Once again, the final configuration seems somewhat close to the initial one, as shown by the attention pattern reported in Figure 8.
-This is consistent with the observations made in the previous subsections.
-Figure 10: Same circuit as in Figure 4 with an additional token embedding for $p=3$.
-Figure 11: Circuit learned after $1000$ epochs of finetuning with $p=3$.
-Overall, our transfer experiment highlights the transferability of circuits and the usefulness of a curriculum in facilitating the learning of challenging tasks by inducing effective circuits through tasks that are easier to solve.
-Although our experiments are performed in two stages, we hypothesize that the same type of mechanism can explain the importance of data curation.
-4.3 Loss Spikes
+When the model’s training stagnates, the loss plateaus for several epochs before decreasing again.
+This behavior may be understood by the optimizer reaching a saddle point in the loss function w.r.t.
+the model’s parameter, that is the gradient of the loss is (almost) zero, but the optimization has not yet reached a local minimum.
+The connection between gradient norms and learning phases is salient in Fig.
+3 where we can see that loss drops occur in tandem with high gradient norms for each layer.
+4.2 Occurence of Loss Spikes
+Better understanding why loss spikes occur is of great interest as they can be detrimental to the training stability, particularly with very large-scale transformers such as large language models (Chowdhery et al., 2023).
+While some know-how exists for mitigating the impact of peak losses, it is often experimentally motivated and requires a lot of computational resources to achieve operational success (Grattafiori et al., 2024).
 One interesting aspect of our sandbox is that it generates loss spikes that we can study quite precisely.
-Our visual inspection showcases two aspects beyond loss spikes.
-The high curvature of multi-layer perceptrons with heavy weights, or with numerous small correlated weights, as well as the high curvature of the RMS normalization layer near the origin.
-Figure 12: Loss spikes are linked to the high curvature of internal network functions.
-A small update to an element $x$ can result in a substantial change to its normalized version $x/\left\|x\right\|$, significantly altering the network’s subsequent behavior.
-Figure 13: Loss spikes (both left) resulting from a small change from one iteration to another in sequence embeddings that are close to the decision boundaries of the subsequent feedforward layer (both right).
-Figure 14: A small change in sequence embeddings (left) can lead to a big change in MLP response (right). This is due to heavy, or small but heavily correlated assemblers (middle).
+Our visual inspection showcases two aspects leading to loss spikes: the high curvature of the RMS normalization layer near the origin as well as the high curvature of multi-layer perceptrons with heavy weights, or with numerous small correlated weights.
+High-curvature normalization.
 Figure 5 illustrates that a small modification to an element can result in a disproportionately large change in its normalized version.
@@
 (2022), further theoretical insights being provided by Cabannes et al.
-(2024b).
+(2024c).
 Interestingly, this analysis suggests removing some loss spikes by smoothing out the normalization layer.
 For example, consider using $f(x)=\sigma(\left\|x\right\|)x/\left\|x\right\|$ where $\sigma$ is a smooth function with $\sigma(0)=0$ and $\sigma([1,\infty))=\{1\}$.
-This demonstrates the usefulness of our visual sandbox in gaining insights and building intuition, which can then be validated on a larger scale in separate projects.
+This demonstrates the usefulness of our visual sandbox in gaining insights and building intuition, which can then be validated on a larger scale in subsequent works.
+Figure 5: High-curvature normalization.
+Loss spikes are linked to the high curvature of internal network functions.
+A small update to an element $x$ can result in a substantial change to its normalized version $x/\left\|x\right\|$, significantly altering the network’s subsequent behavior.
+High-curvature feed-forward.
 Another source of loss spikes is illustrated in Figure 4.
@@
 Once again, one can imagine different ways to regularize these types of loss spikes, with various regularization measures, or by ensuring that the capacity of the MLP is large enough for the MLP to avoid creating these heavy or highly correlated weights.
-To conclude, our visual sandbox enabled us to peek into some reasons behind loss spikes in large neural networks and to build intuition on how to circumvent them, giving us ideas for large-scale experiments, which, if successful, could help increase the amount of intelligence learned for a given amount of training compute.
-Connection to gradient norms and sparsity is discussed in Appendix A.2 and A.3.
-Conclusion
-This paper aims to advance our understanding of neural network training dynamics through a detailed study of a small-scale transformer model, facilitated by a novel visual sandbox.
-This tool allows us to observe and analyze the model’s internal mechanisms vividly, providing both theoretical and practical insights that could lead to more efficient training strategies.
-Our findings highlight a hierarchical learning process where low-level features are developed first, followed by a more predictable refinement of the model.
-We also emphasize the impact of initial configurations on the final model outcomes and the seemingly unpredictable nature of some feature learning phase transitions during training.
-Additionally, we explored the transferability of learned behaviors, which relates to the importance of data curation and curriculum learning in enhancing model performance.
-We also addressed loss spikes caused by high curvature in the model’s internal functions and proposed potential solutions to mitigate these issues.
-Looking ahead, future work will aim to apply these insights to larger and more complex models, assess the scalability of observed phenomena, and enhance our visualization tools for higher-dimensional data.
-Societal Impact.
-This paper aims to understand the complex dynamics at play in neural networks.
-The insights gained from our research stream could help in designing more efficient and robust neural networks, ultimately aiding in reducing the computational cost and environmental impact of training large models.
-Appendix A Appendix
-A.1 Reproducibility Statement
-We open-source the code at https://github.com/facebookresearch/pal and provide the implementation details in Section 2.2 to ensure reproducibility.
-A.2 Understanding Saddle points in Loss Optimization via Gradient Norms
-When the model’s training stagnates, the loss plateau for several epochs before it decreases again.
-This behavior may be understood by the optimizer reaching a saddle point in the loss function w.r.t.
-the model’s parameter, that is the gradient of the loss is (almost) zero, but the optimization has not reached yet a local minimum.
-We illustrate this phenomenon in the full-batch and mini-batch setup, respectively in Figures 16 and 16.
-The connection between gradient norms and learning phases is salient in Figure 16.
-We can see that loss drops occur in tandem with high gradient norms for each layer.
-As studied in Section 4.1, these drops correspond to successive learning phases.
-This is even more salient in the last subfigure of Figure 16 where the pics precisely match the drops.
-This shows the connection between the exit of a saddle point and high gradient norms.
-Similarly, in Figure 16, we can see that the first drop in the loss (around epoch $250$ in the first plot) appears when the first increase in gradient norm occurs and the second point of inflection (around epoch $700$) also match an infection point on the gradient norms.
-It should be noted that this is less salient compared to the full-batch setup.
-Figure 15: Connection between loss spikes and gradient norm in full-batch setup.
-From left to right: Evolution of train and test losses, the corresponding accuracies, the evolution of gradient norms for each layer in log-scale, and the similar evolution in linear scale.
-We see that the learning phases of Section 4.1 appear in tandem with high gradient norms.
-This can be seen in the last subfigure where the three pics correspond to the loss drops and their corresponding plateaus
-Figure 16: Mini-batch setup.
-This study, akin to Figure 16, makes the connection between gradient norm and loss profile even more salient for our problem.
-The first loss drop appears simultaneously with the first increase in gradient norms.
-Loss spikes occur in tandem with spikes in gradient norms.
-However, this setup enables us to study another training behavior mentioned in Sectin 4.3.
-Indeed, we observe in Figure 16 that the loss spikes appear in tandem with high gradient norms, indicating that when a too-large step size deviates the model from its current small loss region, it is taken back to where it was with large updates.
-It has been shown in the literature (Ilbert et al., 2024; Foret et al., 2021; Zhang et al., 2024) that studying the gradient and the hessian of the loss could provide valuable insights on the neural network optimization both from a theoretical and experimental point of view.
+Figure 6: A small change in sequence embeddings (left) can lead to a big change in MLP response (right). This is due to heavy, or small but heavily correlated assemblers (middle).
+Connection to gradient norms.
+Finally, we observe in Fig.
+7 that the loss spikes appear in tandem with high gradient norms, indicating that when a too-large step size deviates the model from its current small loss region, it is taken back to where it was with large updates.
+We note that we observe similar patterns in larger models with embeddings size $d>2$ (see Appendix B).
+It has been shown in the literature that studying the gradient and the hessian of the loss could provide valuable insights on the neural network optimization both from a theoretical and experimental point of view (Ilbert et al., 2024; Foret et al., 2021; Zhang et al., 2024b).
 We believe using our visual sandbox could help in future work on such an investigation.
-A.3 Impact of Activation Sparsity on Models Performance
+Figure 7: Loss spikes and gradient norms.
+This study, akin to Fig.
+3 in mini-batch, makes the connection between gradient norm and loss spikes salient as loss spikes occur in tandem with spikes in gradient norms.
+4.3 Further Observations
+Our sandbox allows us to make other interesting observations, such as the impact of initialization on the type of final circuits, the importance of active neurons, and the usefulness of curriculum learning to induce the correct invariance.
+Impact of the initialization.
+We observe in Fig.
+8 the strong similarity between the attention patterns at the start and at the end of the training, indicating that initialization plays a crucial role in the concrete realizations of clustering heads.
+We note that the transformer learns to focus on the $1$ among the non-spurious tokens, before counting them and deducing the number of 00 to make its final prediction.
+Figure 8: Impact of the initialization.
+For each sequence (left), we plot the corresponding attention pattern at the start (middle) and the end of training (right).
+The darker, the higher the attention weight.
+The final attention maps are correlated with the original one, illustrating that variations of the idealized circuit depend on the original configuration.
+We see that the transformer learns to focus solely on the first $k=5$ input tokens, which are the ones defining the output $y$.
+Figure 9: This study is akin to Fig.
+8 when the model is first pre-trained with $p=2$ and then finetuned with $p=3$.
+This illustrates that the final variations of the original circuit depend on the original weight configuration.
+We notice that the model learns better to focus on the non-spurious tokens.
+Figure 10: Connection between performance and sparsity.
+Since the activation is the GELU that does not map all the negative entries to 00 (contrary to the ReLU), we define the activation sparsity as the percentage of entries with values lower than $\varepsilon>0$.
+We display the evolution of the activation sparsity of $20$ trained models with $\varepsilon\in[10^{-5},10^{2}]$.
+Left: Successful models (i.e., with test accuracy above $0.9$) in blue have less sparse activation than failed models in red.
+Right: The color indicates the models’ test accuracy (the lighter, the better).
+The performance increases as the activation sparsity decreases.
+Transferability of circuits.
+Among the lessons of training very large models is that some sources of data may facilitate the learning of certain skills (Touvron et al., 2023b; Llama Team, 2024; Cabannes et al., 2024a).
+In Section A.3, we illustrate this phenomenon in the learning of clustering heads.
+In particular, we show that training first a model on a simpler task (e.g., $p=2$) helps it better solve a more challenging one (e.g., $p=3$) as can be seen in Fig.
+9 along with the impact of initialization.
+Although in two stages, our experiments might indicate that the transferability of circuits can explain the importance of data curation in more practical settings.
+Connectivity of neurons.
+As explained in Section 3.1, clustering heads capture the invariance of the sparse modular addition problem.
+Motivated by recent works studying the sparsity of transformers’ activations (Li et al., 2023a; Mirzadeh et al., 2024), we show in Fig.
+10 that the models that successfully solve the task have denser activation.
+This might indicate that many neurons are necessary to correctly understand the invariance, hence approximating the idealized circuit well (more details in Section A.2).
+Generalization to bigger embeddings.
+While the previous observations were enabled by our toolbox, which precisely tracks the training dynamics when embeddings reside in ${\mathbb{R}}^{2}$, our findings are confirmed when considering larger embedding dimensions in Appendix B.
+In particular, we discuss the two-phase learning in Section B.1, the loss spikes in Section B.2, and the activation sparsity in Section B.3.
+We also note the coherence of our results with existing literature in more practical settings (Li et al., 2023b; Mirzadeh et al., 2024; Zhang et al., 2024b; Ilbert et al., 2024; Cabannes et al., 2024c).
+4.4 Theoretical Insights
+In this section, we provide theoretical insights on the learning of clustering heads.
+In practice, our transformer is trained via gradient descent which converges to the stationary points of $\mathcal{L}$, i.e., solutions of $\nabla\mathcal{L}=0$.
+The closed-form gradients are given in Eq.
+5 and these stationary points are characterized exactly in C.2.
+They notably include the clustering heads construction, and all their variants which allow them to reach the loss minimum ${\mathcal{L}}^{*}=H(Y\mid X)$ which is the conditional entropy of $Y$ given $X$.
+To better understand the two-staged dynamic observed in Section 4.1, we investigate in Section 4.4 how the gradient magnitude is controlled along the training.
+The formal statement is deferred to Section C.4.
+{boxprop}[Gradient upper-bound, informal] Assuming bounded token embedding, the gradient norm is upper-bounded by
+$\lVert\nabla\mathcal{L}(\theta)\rVert\leq\mathcal{O}\left(\tilde{B}\cdot\sqrt{\mathscr{E}}\right),$
+where $\mathscr{E}=\mathbb{E}_{x}\left[d_{\mathrm{TV}}\left(\hat{p}_{\theta}(\cdot|x),p(\cdot|x)\right)\right]$ captures the ability of the transformer to solve the task and $\tilde{B}$ depends solely on the spectral norm of the model’s parameters.
+Since our model does not include any theoretical simplification to the transformer architecture, the proofs are somewhat involved.
+Section 4.4 is useful to understand the loss profiles identified in Section 4.1.
+Notably, we see that the term $\mathscr{E}$ captures the ability of the model to correctly solve the task.
+In the early iterations, the model behaves poorly which imposes a weak control on the gradient norm.
+This allows for large gradient norms and hence for bigger updates to the weights.
+The loss can vary rapidly as can be seen by the steep descent of Fig.
+2.
+During training, the model improves as can be seen by the increase in accuracy (both train and test).
+This leads to a smaller $\mathscr{E}$ and strengthens the control on the norm.
+In parallel, the gradient updates are of smaller magnitude as well as the loss variations.
+This implies a less steep descent in the loss as observed in Fig.
+2.
+5 Discussion
+This paper aims to advance our understanding of the transformer’s training dynamics through a detailed study of a small-scale model.
+We introduced a visual sandbox that provides comprehensive visualizations of each layer throughout the training process.
+It allowed us to reveal a type of circuit, called "clustering heads," which learns the invariants of the problem.
+We analyzed the learning dynamics of these circuits, showcasing two-stage learning, and loss spikes caused by high curvature in the model’s internal functions.
+In addition, we studied the effect of initialization and curriculum learning.
+Our theoretical insights highlighted that clustering heads can indeed be learned by gradient descent and provided an understanding of the two-stage learning observed in practice.
+Future work.
+Future work will aim to apply these insights to larger and more complex models, assess the scalability of observed phenomena, and enhance our visualization tools for higher-dimensional models.
+Impact Statement
+This paper aims to understand the complex dynamics at play in transformers.
+The insights gained from our research stream could help design more efficient and robust models and optimization schemes, ultimately aiding in reducing the computational cost and environmental impact of training large models.
+Acknowledgements
+We thank Elvis Dohmatob, Eshaan Nichani, Giuseppe Paolo, Faniriana Rakoto Endor, and Ievgen Redko for the fruitful discussions that led to this work.
+Author Contributions
+Ambroise worked on the codebase, the experiments, the proofs, and the writing.
+Wassim worked on the codebase, the experiments, and the writing.
+Vivien proposed the initial idea, worked on the codebase, and supervised the project.
+Each author was essential to this work.
+Appendix
+Roadmap.
+We provide additional experiments in Appendix A.
+We go beyond our controlled setting and present results with higher-dimensional embeddings in Appendix B.
+In Appendix C, we discuss in detail the theoretical insights on the learning dynamics.
+Finally, the proofs of our theoretical results are given in Appendix D.
+The corresponding table of contents is presented below.
+Appendix A Experimental Study of Clustering Heads
+In this section, we present additional results related to the study of the circuits implemented by the transformer.
+Notably, we investigate the concrete realizations to solve the task, the sparsity of the model’s activation along the training, and the transferability of circuits.
+Implementation details.
+We open-source the code at https://github.com/facebookresearch/pal.
+The implementation details are given in Section 2 along with the default configuration below.
+In our experiments, we trained our networks using $n=2048=2^{11}$ data points, which were sampled uniformly with replacement from the $p^{L}$ possible sentences.
+We utilized the Adam optimizer with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$ (Kingma & Ba, 2017), and we initialized the network weights using the default schemes provided by PyTorch (Paszke et al., 2019).
+A.1 Concrete Realizations
+In practice, the model weights learned through gradient descent deviate from the idealized model.
+In this section, we visualize the variations of the idealized circuit described in Section 3.2 and categorized as follows:
+  1. 1.
+    Faulty attention corrected by value: the attention is compensated by the value matrix to capture well the invariance,
+  2. 2.
+    Partially learned invariants: the sentence embeddings present more clusters than the idealized model,
+  3. 3.
+    Fuzzy constructions: the clustering is unconventional but the model still generalizes well.
+Figure 11: Faulty attention corrected by value collapse.
+Figure 12: Embeddings having only learned some invariants.
+Figure 13: Fuzzy construction.
+A.2 Connectivity of Neurons
 Sparsity is a phenomenon of interest in many fields such as signal processing, neuroscience, and machine learning (Mairal et al., 2009; Chen et al., 1998; Barth & Poulet, 2012).
@@
 In particular, Li et al.
-(2023) showed that trained transformers have sparse activations and concluded that it was caused by the training dynamics rather than by a compact representation of the training data as commonly thought in computer vision and NLP.
+(2023a) showed that trained transformers have sparse activations and concluded that it was caused by the training dynamics rather than by a compact representation of the training data as commonly thought in computer vision and NLP.
 Mirzadeh et al.
@@
 It should be noted that those works study deep transformers and identified that the sparsity increases with the depth while we only consider a one-layer transformer.
+In our setting, the activation sparsity can help us better understand how neurons connect to implement clustering heads.
+Successful models have dense activations.
 Following the framework from Li et al.
-(2023), we recall that the activation sparsity corresponds to the percentage of non-zero entries of the feed-forward activation map.
+(2023a), we recall that the activation sparsity corresponds to the percentage of non-zero entries of the feed-forward activation map.
 Without loss of generality, the feed-forward block is an MLP with weights $W_{1},W_{2}$ and a non-linear activation $\sigma$ that outputs for any input $x$ a vector $z=W_{1}\sigma(W_{2}x)$.
@@
 This is the case of the SiLU (Elfwing et al., 2018), used in Llama models (Touvron et al., 2023a), and of the GeLU Hendrycks & Gimpel (2023) used in Falcon (Almazrouei et al., 2023), PaLM (Chowdhery et al., 2024), and in our transformer implementation.
-Instead of replacing such activations by a ReLU (Mirzadeh et al., 2024; Li et al., 2023), we compute a smoothed sparsity [^1] This is akin to using the $\ell_{1}$-norm, respectively the nuclear norm, instead of the $\ell_{0}$ quasi-norm, respectively the rank (Ilbert et al., 2024; Gribonval & Nielsen, 2003).
+Instead of replacing such activations with a ReLU (Mirzadeh et al., 2024; Li et al., 2023a), we compute a smoothed sparsity.
+This is akin to using the $\ell_{1}$-norm, respectively the nuclear norm, instead of the $\ell_{0}$ quasi-norm, respectively the rank (Ilbert et al., 2024; Gribonval & Nielsen, 2003).
 as the percentage of entries with an absolute value lower than $\varepsilon>0$.
 A sparsity of $1$ means that all entries are $\varepsilon$-close to 00 (i.e., sparse activations) and a sparsity of 00 means that all entries are at least $\varepsilon$-away from 00 (i.e., dense activations).
-Figure 17: Connection between performance and sparsity.
-We display the evolution of the activation sparsity of $20$ trained models with $\varepsilon\in[10^{-5},10^{2}]$.
-Left: Successful models (i.e., with test accuracy above $0.9$) in blue have less sparse activation than failed models in red.
-Right: The color indicates the models’ test accuracy (the lighter, the better).
-The performance increases as the activation sparsity decreases.
-To better understand the impact of sparsity, we train $20$ independent models and display in Figure 17 their sparsity after training for $\varepsilon\in[10^{-5},10^{2}]$ (the range is chosen such that the sparsity reaches its extremal values 00 and $1$).
+Evolution of the sparsity with performance.
+To better understand the impact of sparsity, we train $20$ independent models and display in Fig.
+10 their sparsity after training for $\varepsilon\in[10^{-5},10^{2}]$ (the range is chosen such that the sparsity reaches its extremal values 00 and $1$).
 Given the task’s difficulty, achieving an accuracy above $0.9$ is a success; otherwise, it is a failure.
@@
 We observe a striking separation between successful and failed training.
-In the admissible range $[10^{-3},10]$, successful models tend to have less sparse activation than failed ones.
-To further study this phenomenon, we plot in the right subplot of Figure 17 the evolution of the sparsity with $\varepsilon$, and here, the color indicates the models’ test accuracy (the lighter the color, the better the model).
+In the permissible range $[10^{-3},10]$, successful models tend to have less sparse activation than failed ones.
+To further study this phenomenon, we plot in the right subplot of Fig.
+10 the evolution of the sparsity with $\varepsilon$, and here, the color indicates the models’ test accuracy (the lighter the color, the better the model).
 We can see that the sparsity decreases as the performance increases.
 This explains the sharp transition between failure and success in the left subplot.
-This experiment seems to indicate that, contrary to images and textual data (Li et al., 2023; Mirzadeh et al., 2024), the sparse modular addition problem needs the involvement of many neurons during inference, and hence requires non-sparse activations.
+These experiments seem to indicate that, contrary to images and textual data (Li et al., 2023a; Mirzadeh et al., 2024), the sparse modular addition problem needs the involvement of many neurons during inference, and hence requires non-sparse activations.
+A.3 Transferability of Circuits
+Among the lessons learned from training very large models is the importance of careful data engineering (Touvron et al., 2023b; Llama Team, 2024).
+Cabannes et al.
+(2024a) has highlighted that certain sources of data may facilitate the learning of invariances, while Abbe et al.
+(2023) discusses how data curation enables models to escape saddle points more quickly.
+These insights are consistent with the observations we made regarding our problem.
+In the sparse modular addition problem, the number of unique sequences is equal to $p^{T}$, which increases rapidly with both $p$ and $T$ and makes the problem quite hard to learn.
+In particular, when setting $T=12$ and limiting the training set to $n=2048$ data points, training for 1000 epochs does not result in any learning for $p>4$.
+(a) Mean
+(b) Maximum
+Figure 14: Accuracies obtained from pretraining only with $p\in[2,10]$ and finetuning with $p\in[3,10]$ starting from $p=2$ for various vocabulary sizes.
+Left: Averaged accuracy; Right: Maximum accuracy.
+Finetuned models display better performances than pretrained-only models.
+Impact on the performance.
+As previously seen in Figure 13, when training with $p=2$, we often find circuits that capture both permutation and suffix invariants.
+These invariants generalize for any $p\in{\mathbb{N}}$ when $k$ and $T$ are fixed.
+Consequently, initializing models with these invariants makes learning the sparse modular addition problem much easier.
+This observation was made after conducting the following experiments: we first trained a model with sequences in ${\mathbb{F}}_{p}$ for $p=2$ over 1000 epochs, before switching the dataset to sequences in ${\mathbb{F}}_{p}$ for $p=3$ for another 1000 epochs.
+We found that this procedure significantly facilitates the learning of the sparse modular addition problem for $p=3$, which we summarize in Figure 14 where each bar plot is obtained by over $250$ runs.
+Remarkably, we found that the only models that achieved 100% test accuracy were those that captured both the token and permutation invariances after the first 1000 epochs.
+Specifically, these were the models that created six sequence embedding clusters, as shown in Figures 1 and 13, rather than those depicted in Figures 13 or 13.
+Impact on the learned circuits.
+Figure 15(b) shows the final circuit found in one of our finetuning experiments.
+The training was initialized with the circuit in Figure 13, after adding a token embedding to encode for $x=2$, resulting in Figure 15(a).
+The final embeddings are not that far from the initial one, with the transformer having learned to mainly pay attention to the non-spurious tokens that are not equal to 2.
+It also pays some attention to 0 and 1 in positions $t=7$ and $t=10$.
+However, this faulty attention is corrected by the value matrix.
+Once again, the final configuration seems somewhat close to the initial one, as shown by the attention pattern reported in Figure 8.
+This is consistent with the observations made in the previous subsections.
+(a) Initialization with an additional token embedding for $p=3$.
+(b) Circuit learned after $1000$ epochs of finetuning with $p=3$.
+Figure 15: Transfer of clustering head. The obtained circuit is akin to Fig. 13 where the value corrects the faulty attention.
+Benefits of transfer.
+Overall, our transfer experiment highlights the transferability of circuits and the usefulness of a curriculum in facilitating the learning of challenging tasks by inducing effective circuits through tasks that are easier to solve.
+Although our experiments are performed in two stages, we hypothesize that the same type of mechanism can explain the importance of data curation.
+A.4 Impact of the Hyperparameters
+In this section, we conduct ablation studies on the training hyperparameters.
+In Fig.
+16(a), we observe the high variability of the test accuracy for a fixed seed after $1000$ epochs when varying the batch size, hidden dimension, learning rates, and MLP learning rates.
+Fig.
+16(b) is similar with more regularity as it is averaged over $100$ runs with different seeds.
+This study shows that the exit of saddle points is highly dependent on the hyperparameters.
+(a) Single run with 1000 epochs.
+(b) Averaged over $100$ runs with standard deviations.
+Figure 16: Ablation studies regarding test accuracy as a function of batch size, hidden neurons, learning rates, and MLP learning rates discount factor for a single run with 1000 epochs.
+We ensured consistency in initial weights and batch designs when changing hyperparameters.
+We observe (mostly in Fig.
+16(b)) some regularity in the effect of hyperparameters on the resulting test accuracy.
+Appendix B Beyond Low-dimensional Embeddings
+Relying on our visual sandbox, we studied several phenomena on the training dynamics of neural networks from the learning of the different parts of the network in Section 4.1 to the loss spikes occurring during the optimization in Section 4.2 through the efficiency of transfer learning in Section A.3.
+We also analyzed the connection between saddle points, loss spikes, and gradient norms in Sections 4.1 and 4.2 and the impact of the activation sparsity on the models’ performance in Section A.2.
+The low-dimensional embeddings allowed us to pair loss profiles with the visualization of each layer to decipher the training behaviors.
+However, we note that the studied behaviors and the insights we obtain can be analyzed independently of the embedding dimension $d$.
+Figure 17: Connection to saddle points.
+This study is akin to Fig.
+3 with $d=3$ in the top plot and $d=64$ in the bottom one.
+From left to right: Evolution of train and test losses, the corresponding accuracies, the evolution of gradient norms for each layer in log-scale, and the similar evolution in linear scale in full-batch.
+Akin to Fig.
+3, we see that the learning phases studied in Section 4.1 appear in tandem with high gradient norms.
+This can be seen in the last subfigure where the pics correspond to the loss drops and spikes.
+Motivation.
+Taking an embedding size $d=2$ was mainly motivated by visualization purposes and we saw that such a setting was already challenging enough to observe training phenomena of interest in more practical settings.
+For the sake of self-consistency, we extend our experiments with higher-dimensional embeddings and show that our insights remain valid in more practical settings with $d>2$.
+B.1 Two-Phase Learning and Saddle Points
+In this section, we extend the experiments of Section 4.1 for $d>2$.
+In Fig.
+17, we consider $d=3$ (top) and $d=64$ (bottom).
+We obtain similar conclusions than in the case $d=2$ in Fig.
+3 where the loss drops occur in tandem with high gradient norms for each layer, hinting at the exit of saddle points.
+B.2 Loss Spikes and Gradient Norms
+In this section, we extend the experiments of Section 4.2 for $d>2$.
+In Fig.
+18, we consider $d=3$ (top) and $d=64$ (bottom).
+We observe a similar behavior as in the case $d=2$ in Fig.
+7 where the loss spikes and gradient norm spikes occur at the same time.
+This again highlights the connection between gradient norms and loss spikes.
+Figure 18: Loss spikes and gradient norms.
+This study is akin to Fig.
+7 in mini-batch with $d=3$ in the top plot and $d=64$ in the bottom one.
+Again, we observe that the loss spikes and gradient norms spikes appear in tandem.
+B.3 Activation Sparsity and Performance
+Finally, we extend the experiments of Apppendix A.2 with an embedding size $d\in\{2,3,4,8,16,32\}$ in Fig.
+19.
+We first note that the higher the embedding size, the more the model succeeds at the task.
+Especially, as of $d=8$, all the models are successful, i.e., they all achieve an accuracy higher than $0.9$ (as defined in Section A.2).
+This is expected given that imposing low-dimensional embeddings limits the expressiveness and the generalization power of our model.
+It should be noted that this was one of the many challenges of our study: obtaining a generalizable neural network with embeddings in $R^{2}$ for a mathematical reasoning task such as the sparse modular addition problem.
+We obtain similar conclusions than in Section A.2 with successful models having more dense activations.
+Figure 19: Connection between performance and sparsity.
+We display the evolution of the activation sparsity of $20$ trained models with $\varepsilon\in[10^{-5},10^{2}]$ for $d\in\{2,3,4,8,16,32\}$.
+Successful models (i.e., with test accuracy above $0.9$) in blue have less sparse activation than failed models in red.
+Appendix C Theoretical Insights on the Learning Dynamics
+In this section, we detail our theoretical insights on the learning dynamics.
+It should be noted that all our derivations are conducted with arbitrary embedding dimension $d$ and we do not consider theoretical simplifications of the transformer we study (introduced in Section 2.2).
+C.1 Notations
+In this section, we recall the notations used in our theoretical results and proofs.
+We denote $\{1,\cdots,N\}$ as $[N]$.
+The $i$-th row of a matrix $A\in{\mathbb{R}}^{n\times m}$ is denoted by $A_{i}$, its $j$-th column is denoted by $A_{\cdot,j}$ and its transpose is denoted by by $A^{\top}$.
+By abuse of notation, we denote by $\lVert A\rVert$ the entry-wise $\ell_{2}$-norm of the matrix $A$ which amounts to computing the Euclidean norm of $A$ flattened.
+Note that we have $\lVert A\rVert=\lVert A\rVert_{\mathrm{F}}=\sqrt{\sum_{i=1}^{\min\{n,m\}}\sigma_{i}(A)^{2}}$, where $\lVert\cdot\rVert_{\mathrm{F}}$ is the Frobenius norm of $A$ and the $\sigma_{i}(A)$ are its singular values.The operator norm (also called spectral norm) of $A$ writes $\lVert A\rVert_{\mathrm{op}}=\sigma_{\mathrm{max}}(A)$, with $\sigma_{\mathrm{max}}(A)$ the largest singular value of $A$.
+The max norm of $A$ is denoted by $\lVert A\rVert_{\infty}=\max_{ij}\lvert A_{ij}\rvert$.
+The identity matrix of size $n$ is denoted by $I_{n}\in{\mathbb{R}}^{n\times n}$.
+The indicator function is denoted by $\mathbbm{1}$ and verifies $\mathbbm{1}\{a=b\}$ with value $1$ if $a=b$ and 00 otherwise.
+The transpose of a vector $x$ is denoted by $x^{\top}$ and its Euclidean norm is denoted by $\lVert x\rVert=\sqrt{x^{\top}x}$.
+The total variation distance between two probability distributions $p,q$ is denoted by $d_{\mathrm{TV}}\left(p,q\right)$.
+For ease of notation, we define the following quantities for $\xi\in{\mathbb{R}}^{d},W\in{\mathbb{R}}^{h\times d},z\in{\mathbb{R}}^{d\times N},q\in{\mathbb{R}}^{d}$ and the GELU function $\sigma\colon{\mathbb{R}}\to R$:
+$\bar{\xi}=\frac{\xi}{\lVert\xi\rVert}\in{\mathbb{R}}^{d},$ (1)
+$A_{z}=\mathrm{diag}\left(\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\right)\in{\mathbb{R}}^{N\times N},$ (2)
+$\Sigma_{\bar{\xi}}=\mathrm{diag}(\sigma^{\prime}(W\bar{\xi}))\in{\mathbb{R}}^{h\times h}.$ (3)
+$\Delta_{z}=\frac{z}{\sqrt{d}}\left[I_{N}-\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\mathbbm{1}_{N}^{\top}\right].$ (4)
+C.2 Gradients Derivations
+In this section, we derive the gradient of the loss with respect to the neural network’s weights.
+Transformer model.
+We first recall below the quantities involved in the gradient computations:
+  1. 1.
+    Input: $x=[x_{1},\ldots,x_{L}]\in(\mathbb{F}_{p})^{L}$,
+  2. 2.
+    Embedding: $z_{t}:=Z(x_{t},t):=\frac{E(x_{t})+P(t)}{\left\|E(x_{t})+P(t)\right\|}\in{\mathbb{R}}^{d}$ with $P,E$ learnable,
+  3. 3.
+    Attention: $\xi:=(Vz)\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\in{\mathbb{R}}^{d}$ with $V\in{\mathbb{R}}^{d\times d},q\in{\mathbb{R}}^{d}$ learnable,
+  4. 4.
+    Feed-forward: $\psi:=\xi+\sum_{i\in[h]}u_{i}\cdot\sigma\left(\frac{w_{i}^{\top}\xi}{\|\xi\|}\right)=\xi+U\sigma\left(\frac{W\xi}{\|\xi\|}\right)\in\mathbb{R}^{d}$ where $U\in\mathbb{R}^{d\times h}$ has columns $(u_{i})_{i=1}^{h}$, $W\in\mathbb{R}^{h\times d}$ has rows $(w_{i}^{\top})_{i=1}^{h}$ and $U,W$ are learnable,
+  5. 5.
+    Unembedding: $\zeta=(E(v)^{\top}\psi)_{v\in\mathbb{F}_{p}}=W_{U}\psi\in\mathbb{R}^{p}$ where $W_{U}\in{\mathbb{R}}^{p\times d}$ has rows $E(v)^{T}$ and is learnable.
+To avoid cumbersome derivations that do not fundamentally change the analysis and since we are mostly interested in the dynamics at work in the attention and the feed-forward layers, we assume that that the embedding matrices $E,P$ are randomly initialized and do not participate in the optimization.
+It also concerns the unembedding matrix $W_{U}$ thanks to weight tying.
+This is a common assumption in prior works studying the training dynamics of transformers (Cabannes et al., 2024b; Bietti et al., 2023; Cabannes et al., 2024c; Nichani et al., 2024).
+To derive the loss gradients, we first reformulate the loss.
+Gradient of the training loss.
+We consider the vector $\theta$ that encompasses the model’s parameters.
+More specifically, $\theta$ can be seen as the (flattened) concatenation of $q\in{\mathbb{R}}^{d},V\in{\mathbb{R}}^{d\times d},W\in{\mathbb{R}}^{h\times d},U\in{\mathbb{R}}^{d\times h}$, i.e., as a vector of $R^{m}$ with $m=d(1+d+2h)$.
+We recall that the cross-entropy loss can be reformulated as
+$\mathcal{L}(\theta):=\mathbb{E}_{(x,y)\sim q}\left[\ell(\theta;x,y\right],$
+where we have
+$\ell(\theta;x,y)=-\sum_{i=1}^{p}\mathbbm{1}\{y=i\}\log(\frac{\exp(\zeta_{i})}{\sum_{n=1}^{p}\exp(\zeta_{n})})=\log(\sum_{n=1}^{p}\exp(\zeta_{n}))-\sum_{i=1}^{p}\mathbbm{1}\{y=i\}\cdot\zeta_{i}.$ (5)
+We can now proceed with computing the gradient of the loss function.
+The following lemma, whose proof is deferred to Section D.1, gives the closed-form expression of the gradient.
+{boxlem}[Closed-form gradients] Let $\Delta_{z}=\frac{z}{\sqrt{d}}\left[I_{N}-\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\mathbbm{1}_{N}^{\top}\right]\in{\mathbb{R}}^{d\times N}$, $A_{z}=\mathrm{diag}\left(\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\right)\in{\mathbb{R}}^{N\times N}$, $\Sigma_{\bar{\xi}}=\mathrm{diag}(\sigma^{\prime}(W\bar{\xi}))\in{\mathbb{R}}^{h\times h}$, $M=\frac{1}{\lVert\xi\rVert}\left(I_{d}-\bar{\xi}\bar{\xi}^{\top}\right)\in{\mathbb{R}}^{d\times d}$ and $Q=U\Sigma_{\bar{\xi}}WM\in{\mathbb{R}}^{d\times d}$.
+Assume that the embedding weights $E,P$ are not trainable and let []delimited-[][$\cdot$][ ⋅ ] indicate vector concatenation where matrices are flattened into vectors whenever needed.
+The gradient of the loss function is a vector $\nabla_{\theta}\mathcal{L}(\theta)\in{\mathbb{R}}^{m}$ with $m=d(1+d+2h)$ that writes
+$\nabla\mathcal{L}(\theta)=\sum_{j=1}^{p}\mathbb{E}_{x}\left[(\hat{p}_{\theta}(y=j|x)-p(y=j|x))\cdot\nabla_{\theta}\zeta_{j}\right],$
+where
+$\begin{cases}\nabla_{\theta}\zeta_{j}&=\quad[\nabla_{q}\zeta_{j},\nabla_{V}\zeta_{j},\nabla_{W}\zeta_{j},\nabla_{U}\zeta_{j}]\in{\mathbb{R}}^{m},\\ &\\ \nabla_{q}\zeta_{j}&=\quad\Delta_{z}A_{z}(Vz)^{\top}\left(I_{d}+Q^{\top}\right)E(j)\in{\mathbb{R}}^{d},\\ &\\ \nabla_{V}\zeta_{j}&=\quad\left(I_{d}+Q^{\top}\right)E(j)\left[z\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\right]^{\top}\in{\mathbb{R}}^{d\times d},\\ &\\ \nabla_{W}\zeta_{j}&=\quad\Sigma_{\bar{\xi}}U^{\top}E(j)\bar{\xi}^{\top}\in{\mathbb{R}}^{h\times d},\\ &\\ \nabla_{U}\zeta_{j}&=\quad E(j)\sigma(W\bar{\xi})^{\top}\in{\mathbb{R}}^{d\times h}.\end{cases}$
+We first notice the importance of the token embeddings since each gradient depends (linearly) on the $E(j)$.
+We can then see that the attention gradients are both controlled by $\left(I_{d}+Q^{\top}\right)E(j)$.
+We note that $M=\frac{1}{\lVert\xi\rVert}\left(I_{d}-\bar{\xi}\bar{\xi}^{\top}\right)$ projects any vector into $\bar{\xi}$’s orthogonal complementary space, which is included in the unit-sphere of $R^{d}$ denoted by $\mathbb{S}^{d}=\{x\in{\mathbb{R}}^{d}\text{ s.t.
+}\lVert x\rVert=1\}$.
+Similar orthogonal projections appear in the derivations of Tian et al.
+(2023) and are used to mimic normalization layers in Geshkovski et al.
+(2023).
+With $M$ symmetric, $Q^{\top}=M(U\Sigma_{\bar{\xi}}W)^{\top}$ simply perturbs the vector $E(j)$ before projecting it in the sphere and $\left(I_{d}+Q^{\top}\right)E(j)$ can be seen as a perturbed residual connection.
+Finally, we observe that the gradients of the feed-forward weights have a linear dependency on $U,W$ which seems to allow implementing the decision boundaries once the sequence embeddings are clustered (see Section 4.1).
+Given that the gradient formulae in Eq.
+5 are not straightforward to analyze due to the high dimension and the cross-dependency of the terms, we do not attempt to study the gradient updates, noting that a precise analysis is an important theoretical question for future work.
+Remark C.1 (Faithful to practice).
+The derivations to obtain the gradients in closed form are somewhat involved due to the complexity of the network since we do not theoretically simplify the transformer architecture or the data contrary to most prior works that either model associative memories (Cabannes et al., 2024c, b; Bietti et al., 2023; Nichani et al., 2024), consider orthogonal token embeddings (Bietti et al., 2023; Tian et al., 2023), linearize the attention (Nichani et al., 2024; Zhang et al., 2024a; Mahankali et al., 2024; Ahn et al., 2023), discard residual connections (Tian et al., 2023; Geshkovski et al., 2023; Jelassi et al., 2022) or the feed-forward block (Geshkovski et al., 2023; Edelman et al., 2024; Mahankali et al., 2024; Ahn et al., 2023; Tian et al., 2023; Jelassi et al., 2022).
+Moreover, we preserve the classification setting with the cross-entropy loss used in practice instead of considering the regression case with a square loss (Chistikov et al., 2023; Pesme & Flammarion, 2023).
+While this ensures our analysis stays close to the practice, it also comes with technical challenges.
+C.3 Stationary Points Characterization
+The neural networks considered in this work are learned via gradient descent which converges to one of its fixed points (Nesterov, 2014).
+Such points are stationary points (or critical points) of the loss function (Chi et al., 2019b), formally defined below.
+{boxdef}[Stationary points] The stationary points of $\mathcal{L}$ are the solutions of the Euler equation
+$\nabla\mathcal{L}(\theta)=0.$ (6)
+A stationary point $\bar{\theta}$ can be either a local minimum, a local maximum, or a saddle point of $\mathcal{L}$.
+In deep learning, since most objectives are non-convex and in high-dimension, minima [^1] We omit the mention of maxima since we are mostly interested in minimization problems.
+are rarely global and the gradient descent likely converges to a local minimum or a saddle point.
+In particular, saddle points are prevalent in deep neural networks loss landscape (Dauphin et al., 2014), which can hinder the optimization.
+Remark C.2 (Beyond stationary points).
+We note that distinguishing the critical points between local minimum, local maximum, or saddle points involves computing the hessian of the training loss (Chi et al., 2019b) which is beyond the scope of this work.
+As a first step towards better understanding the training dynamics, we show that in the following theorem that clustering heads can be learned by gradient descent.
+The proof is deferred to Section D.2.
+{boxprop}[Learning clustering heads by gradient descent] Let $\Delta_{z}=\frac{z}{\sqrt{d}}\left[I_{N}-\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\mathbbm{1}_{N}^{\top}\right]\in{\mathbb{R}}^{d\times N}$, $A_{z}=\mathrm{diag}\left(\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\right)\in{\mathbb{R}}^{N\times N}$, $\Sigma_{\bar{\xi}}=\mathrm{diag}(\sigma^{\prime}(W\bar{\xi}))\in{\mathbb{R}}^{h\times h}$, $M=\frac{1}{\lVert\xi\rVert}\left(I_{d}-\bar{\xi}\bar{\xi}^{\top}\right)\in{\mathbb{R}}^{d\times d}$, $Q=U\Sigma_{\bar{\xi}}WM\in{\mathbb{R}}^{d\times d}$, $\bar{Q}=\bar{U}\Sigma_{\bar{\xi}}\bar{W}M\in{\mathbb{R}}^{d\times d}$ and $\mathscr{C}_{x}=\sum_{j=1}^{p}(\hat{p}_{\theta}(y=j|x)-p(y=j|x))E(j)$.
+The stationary points $\theta$ of $\mathcal{L}$ verify
+$\begin{split}&\mathbb{E}_{x}\left[\Delta_{z}A_{z}(Vz)^{\top}\left(I_{d}+Q^{\top}\right)\mathscr{C}_{x}\right]=0,\\ &\mathbb{E}_{x}\left[\left(I_{d}+Q^{\top}\right)\mathscr{C}_{x}\left[z\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\right]^{\top}\right]=0,\\ &\mathbb{E}_{x}\left[\Sigma_{\bar{\xi}}U^{\top}\mathscr{C}_{x}\bar{\xi}^{\top}\right]=0,\\ &\mathbb{E}_{x}\left[\mathscr{C}_{x}\sigma(W\bar{\xi})^{\top}\right]=0.\end{split}$ (7)
+Moreover, Eq.
+7 holds for transformers that implement clustering heads.
+This implies that such circuits can be learned by gradient descent.
+Proposition C.2 allows us to access the final configuration of the model in the form of a system of equations.
+It involves the main quantities of the problem: the trainable weights $q,V,W,U$, the embedding tokens $E(j)$, and the training data $(x,y)$.
+A particularly interesting observation is the linear dependency of each equation on the term $\mathscr{C}_{x}$.
+This term captures the model’s prediction discrepancy with the ground truth through the $\hat{p}_{\theta}(y=j|x)-p(y=j|x)$ and plays an important role in the proof.
+Once the stationary points are characterized, showing that clustering heads can be learned by gradient descent is rather straightforward using the $\mathscr{C}_{x}$ but bears a very important meaning.
+Notably, it shows that the natural way we proposed observed to solve the sparse modular addition problem also qualifies as a fixed-point of gradient descent.
+Remark C.3 (Other pathways).
+While C.2 shows that clustering heads are viable fixed-points of the gradient descent, it does not provide conditions to ensure it is reached, nor qualify them as local minima, global minima, or saddle points.
+This interest research direction is left for future work.
+We note that our experimental investigation suggests that clustering heads solve the sparse modular addition problem but there may be other pathways to do so too.
+In addition, as discussed in Section 3.2, the pathways learned in practice can differ from the idealized circuit although the variations still capture most of the invariants of the problem.
+C.4 Bounding the Gradients
+In Section 4, we connect some of our findings with the evolution of gradient norms along the training.
+Leveraging the closed-form expressions given in Eq.
+5, we can study upper-bound gradient norms to elucidate their dependency on the parameters and data.
+The following proposition, whose proof is deferred to Section D.3 makes those upper bounds explicit.
+{boxprop}[Formal statement of Section 4.4] Assume the token embedding is bounded, i.e., there exists $B>0$ such that $\lVert E\rVert_{\infty}\leq B$.
+We introduce $\mathscr{E}=\mathbb{E}_{x}\left[d_{\mathrm{TV}}\left(\hat{p}_{\theta}(\cdot|x),p(\cdot|x)\right)\right]$ that captures the ability of the model to solve the sparse modular addition problem task.
+We have
+$\lVert\nabla\mathcal{L}(\theta)\rVert\leq\tilde{B}\cdot\sqrt{\mathscr{E}},$
+where
+$\tilde{B}=\mathcal{O}\left(B\sqrt{\lVert V\rVert_{\mathrm{op}}\left[1+\lVert U\rVert_{\mathrm{op}}\lVert W\rVert_{\mathrm{op}}^{2}\right]^{2}+\left[1+\lVert U\rVert_{\mathrm{op}}\lVert W\rVert_{\mathrm{op}}\right]^{2}+\lVert U\rVert_{\mathrm{op}}^{2}+\lVert W\rVert_{\mathrm{op}}^{2}}\right).$
+This result is of interest for two main reasons.
+The first one is that it can be used to understand the loss profiles identified in Section 4.1.
+Notably, we see that the term $\mathscr{E}$ captures the ability of the model to correctly solve the task.
+In the early iterations, the model behaves poorly which imposes a weak control on the gradient norm.
+This allows for large gradient norms and hence for bigger updates to the weights.
+The loss can vary rapidly as can be seen by the steep descent of Fig.
+2.
+Along with the training, the model improves as can be seen by the increase in accuracy (both train and test).
+This leads to a smaller $\mathscr{E}$ and strengthens the control on the norm.
+In parallel, the gradient updates are of smaller magnitude as well as the loss variations.
+This goes along a less steep descent in the loss as observed in Fig.
+2.
+The second benefit of Section C.4 is that one could derive sufficient conditions on the term related to the weight matrices and on the term related to classification to achieve an approximate critical point (Chi et al., 2019b), that is parameters $\theta$ such that the gradient verifies $\lVert\nabla\mathcal{L}(\theta)\rVert\leq\varepsilon$ for some $\varepsilon>0$.
+Since finding local approximate minima can be faster (and more beneficial) than finding critical points (Agarwal et al., 2017), practitioners could derive optimization algorithms that impose an appropriate bound on the two terms of Section C.4 instead of using the gradient descent.
+These interesting considerations are left for future work.
+Appendix D Proofs
+In this section, we provide detailed proof of our theoretical results.
+D.1 Proof of Eq. 5
+We detail below the proof of Eq. 5.
+Proof.
+The proof consists of three main steps:
+  1. 1.
+    Expressing the gradient of the loss with respect to the gradients of the attention embeddings,
+  2. 2.
+    Deriving the gradients for each trainable weight,
+  3. 3.
+    Putting everything together.
+We first formulate the gradient of the loss with respect to the gradient of the attention embeddings.
+The next lemma provides the formula.
+{boxlem}[Gradient compact formulation] Denoting the model’s trainable parameters by $\theta\in{\mathbb{R}}^{m}$ and recalling that the attention embedding $\xi\in{\mathbb{R}}^{d}$ can be seen as a function of $\theta$, we have
+$\nabla_{\theta}\mathcal{L}(\theta)=\sum_{j=1}^{p}\mathbb{E}_{x}\left[(\hat{p}_{\theta}(y=j|x)-p(y=j|x))\cdot\nabla_{\theta}\zeta_{j}\right]$ (8)
+Proof.
+Noting that the logits can be written as $\zeta=g(\theta;x,y)$ for some $g$, the loss [^2] We keep denoting the loss by $\ell$ by abuse of notation.
+can then be written as $\ell(\theta;x,y)=\ell(\zeta)=\ell(g(\theta;x,y))$.
+We drop the $(x,y)$ letter when there is no possible ambiguity to avoid cumbersome notations.
+Using the chain rule, with $D$ denoting the Jacobian operator, we have
+$\begin{split}\nabla_{\theta}\mathcal{L}(\theta)&=\mathbb{E}_{(x,y)\sim q}\left[\nabla_{\theta}\ell(\theta)\right]\\ &=\mathbb{E}_{(x,y)\sim q}\left[\nabla_{\theta}\ell(g(\theta))\right]\\ &=\mathbb{E}_{(x,y)\sim q}\left[(D_{\theta}g(\theta))^{\top}(D_{g(\theta)}\ell(g(\theta)))^{\top}\right]\\ &=\mathbb{E}_{(x,y)\sim q}\left[\nabla_{\theta}g(\theta)\nabla_{\zeta}\ell(\zeta)\right].\end{split}$
+By definition, we have $\nabla_{\theta}g(\theta)=(\nabla_{\theta}\zeta_{1},\ldots,\nabla_{\theta}\zeta_{p})$ and $\nabla_{\zeta}\ell(\zeta)=(\frac{\partial\ell}{\partial\zeta_{j}}(\zeta))\in\mathbb{R}^{p}$.
+Hence, the matrix-vector product writes
+$\begin{split}(\nabla_{\theta}g(\theta))\nabla_{\zeta}\ell(\zeta)&=\sum_{j=1}^{p}\frac{\partial\ell}{\partial\zeta_{j}}(\zeta)\cdot[\nabla_{\theta}g(\theta)]_{\cdot,j}\\ &=\sum_{j=1}^{p}\frac{\partial\ell}{\partial\zeta_{j}}(\zeta)\cdot\nabla_{\theta}\zeta_{j}.\end{split}$
+Using Eq. (5), we have
+$\begin{split}\frac{\partial\ell}{\partial\zeta_{j}}(\zeta)&=\frac{\exp(\zeta_{j})}{\sum_{n=1}^{p}\exp(\zeta_{n})}-\mathbbm{1}\{y=j\}\\ &=\operatorname{softmax}(\zeta)_{j}-\mathbbm{1}\{y=j\}\\ &=\hat{p}_{\theta}(y=j|x)-\mathbbm{1}\{y=j\},\end{split}$
+Putting everything together, we obtain
+$\begin{split}\nabla_{\theta}\mathcal{L}(\theta)&=\mathbb{E}_{(x,y)\sim p}\left[\sum_{j=1}^{p}(\hat{p}_{\theta}(y=j|x)-\mathbbm{1}\{y=j\})\cdot\nabla_{\theta}\zeta_{j}\right]\\ &=\sum_{j=1}^{p}\mathbb{E}_{(x,y)\sim p}\left[(\hat{p}_{\theta}(y=j|x)-\mathbbm{1}\{y=j\})\cdot\nabla_{\theta}\zeta_{j}\right]\\ &=\sum_{j=1}^{p}\mathbb{E}_{x}\left[\mathbb{E}_{y}\left[(\hat{p}_{\theta}(y=j|x)-\mathbbm{1}\{y=j\})\cdot\nabla_{\theta}\zeta_{j}|x\right]\right],\end{split}$
+where the first equality comes from the linearity of the expectation and the second stems from the definition of the conditional expectation.
+As $\hat{p}_{\theta}(y=j|x)$ and $\nabla_{\theta}\zeta_{j}$ do not depend on the random variable $y$, we have
+$\begin{split}&\mathbb{E}_{y}\left[(\hat{p}_{\theta}(y=j|x)-\mathbbm{1}\{y=j\})\cdot\nabla_{\theta}\zeta_{j}|x\right]\\ &\qquad=(\hat{p}_{\theta}(y=j|x)-\mathbb{E}_{y}\left[\mathbbm{1}\{y=j\}|x\right])\cdot\nabla_{\theta}\zeta_{j}\\ &\qquad=(\hat{p}_{\theta}(y=j|x)-p(y=j|x))\cdot\nabla_{\theta}\zeta_{j},\end{split}$
+since we have $x,y\sim p$. In summary, the gradient of the loss writes
+$\nabla_{\theta}\mathcal{L}(\theta)=\sum_{j=1}^{p}\mathbb{E}_{x}\left[(\hat{p}_{\theta}(y=j|x)-p(y=j|x))\cdot\nabla_{\theta}\zeta_{j}\right].$
+Section D.1 implies that deriving the $\nabla_{\theta}\zeta_{j}=\left(\frac{\partial\zeta_{j}}{\partial\theta_{n}}\right)_{n=1}^{m}$ allows us to obtain the final gradient formula.
+The following lemmas provide for a fixed $j$ the gradients of $\xi_{j}$ with respect to the different parts of the neural networks, namely the feed-forward weights $U$ and $W$ and the attention weights $q$ and $V$.
+D.1.1 Attention Derivations
+In this section, we derive the gradients with respect to the attention weights.
+Attention Query.
+The following lemma provides the formula of $\nabla_{q}\xi_{j}$.
+{boxlem}[Query] Let $\Delta_{z}=\frac{z}{\sqrt{d}}\left[I_{N}-\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\mathbbm{1}_{N}^{\top}\right]\in{\mathbb{R}}^{d\times N}$ following Eq.
+(4), $A_{z}=\mathrm{diag}\left(\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\right)\in{\mathbb{R}}^{N\times N}$ following Eq.
+(2), $\Sigma_{\bar{\xi}}=\mathrm{diag}(\sigma^{\prime}(W\bar{\xi}))\in{\mathbb{R}}^{h\times h}$ following Eq.
+(3), $M=\frac{1}{\lVert\xi\rVert}\left(I_{d}-\bar{\xi}\bar{\xi}^{\top}\right)\in{\mathbb{R}}^{d\times d}$ following Eq.
+(15) and $Q=U\Sigma_{\bar{\xi}}WM\in{\mathbb{R}}^{d\times d}$.
+The entries of $\nabla_{\theta}\zeta_{j}$ associated to the value $V$ are the entries of $\nabla_{q}\zeta_{j}\in{\mathbb{R}}^{d}$ where we have
+$\nabla_{q}\zeta_{j}=\Delta_{z}A_{z}\left[\left(I_{d}+Q\right)Vz\right]^{\top}E(j).$
+Proof.
+We want to compute the $\frac{\partial\zeta_{j}}{\partial\theta_{n}}$ for the $\theta_{n}$ that are entries of $q$. We first note that
+$\frac{\partial\zeta_{j}}{\partial\theta_{n}}=\sum_{l=1}^{p}(W_{U})_{jl}\frac{\partial\psi_{l}}{\partial\theta_{n}},$ (9)
+where we have thanks to weight tying, $(W_{U})_{jl}=E(j)_{l}$ is fixed. Moreover, we have
+$\psi_{l}=\xi_{l}+\sum_{i=1}^{h}(u_{i})_{l}\sigma(w_{i}^{\top}\bar{\xi}),$ (10)
+where we recall that $\bar{\xi}=\frac{\xi}{\lvert\xi\rVert}$ following Eq. (1). Considering $n$ such that $\theta_{n}=q_{a}$ for some $a\in[d]$, we have
+$\frac{\partial\psi_{l}}{\partial\theta_{n}}=\frac{\partial\psi_{l}}{\partial q_{a}}=\frac{\partial\xi_{l}}{\partial q_{a}}+\frac{\partial}{\partial q_{a}}(\sum_{i=1}^{h}(u_{i})_{l}\sigma(w_{i}^{\top}\bar{\xi}))=\underbrace{\frac{\partial f_{1}(q_{a})}{\partial q_{a}}}_{\mathrm{LHS}}+\underbrace{\sum_{i=1}^{h}(u_{i})_{l}\frac{\partial f_{2}^{i}(q_{a})}{\partial q_{a}}}_{\mathrm{RHS}},$ (11)
+with
+$\begin{cases}f_{1}\colon&{\mathbb{R}}\to{\mathbb{R}},\quad q_{a}\mapsto\xi_{l}\\ f_{2}^{i}\colon&{\mathbb{R}}\to{\mathbb{R}},\quad q_{a}\mapsto(\sigma\circ f_{3}^{i}\circ f_{4}^{i}\circ f_{5}^{i})(q_{a})=\sigma(w_{i}^{\top}\bar{\xi})\\ f_{3}^{i}\colon&{\mathbb{R}}^{d}\to{\mathbb{R}},\quad\bar{\xi}\mapsto w_{i}^{\top}\bar{\xi}\\ f_{4}^{i}\colon&{\mathbb{R}}^{d}\to{\mathbb{R}}^{d},\quad\xi\mapsto\bar{\xi}=\frac{\xi}{\lVert\xi\rVert}\\ f_{5}^{i}\colon&{\mathbb{R}}\to{\mathbb{R}}^{d},\quad q_{a}\mapsto\xi.\\ \end{cases}$
+We proceed in two steps and first compute the LHS of Eq. (11).
+LHS of Eq. (11): we first note that
+$\xi_{l}=\sum_{t=1}^{N}(Vz)_{lt}\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})_{t}.$
+This leads to
+$\frac{\partial f_{1}(q_{a})}{\partial q_{a}}=\frac{\partial\xi_{l}}{\partial q_{a}}=\sum_{t=1}^{N}(Vz)_{lt}\frac{\partial s_{t}}{\partial q_{a}},$ (12)
+where
+$s_{t}=\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})_{t}=\frac{\exp{(\frac{1}{\sqrt{d}}\sum_{o=1}^{d}z_{ot}q_{o})}}{\sum_{t^{\prime}=1}^{N}\exp{(\frac{1}{\sqrt{d}}\sum_{o=1}^{d}z_{ot^{\prime}}q_{o})}}.$ (13)
+This can be rewritten as
+$s_{t}=\frac{\exp{(\frac{1}{\sqrt{d}}z_{at}q_{a}+\frac{1}{\sqrt{d}}\sum_{o\neq a}z_{ot}q_{o})}}{\sum_{t^{\prime}=1}^{N}\exp{(\frac{1}{\sqrt{d}}z_{at^{\prime}}q_{a}+\frac{1}{\sqrt{d}}\sum_{o\neq a}z_{ot^{\prime}}q_{o})}}=\frac{g(q_{a})}{\tilde{g}(q_{a})},$
+where
+$\begin{cases}&g\colon x\to\exp{(\frac{1}{\sqrt{d}}z_{at}x+\frac{1}{\sqrt{d}}\sum_{o\neq a}z_{ot}q_{o})}\\ &\tilde{g}\colon x\to\sum_{t^{\prime}=1}^{N}\exp{(\frac{1}{\sqrt{d}}z_{at^{\prime}}x+\frac{1}{\sqrt{d}}\sum_{o\neq a}z_{ot^{\prime}}q_{o})}\end{cases}.$
+This leads to
+$\frac{\partial s_{t}}{\partial q_{a}}=\frac{g^{\prime}(q_{a})\tilde{g}(q_{a})-\tilde{g}^{\prime}(q_{a})g(q_{a})}{[\tilde{g}(q_{a})]^{2}}.$
+By simple application of the chain rule, we have
+$g^{\prime}(q_{a})$ $=\frac{1}{\sqrt{d}}z_{at}\exp{(\frac{1}{\sqrt{d}}\sum_{o=1}^{d}z_{ot}q_{o})}$
+$=\frac{1}{\sqrt{d}}z_{at}\exp{(\frac{1}{\sqrt{d}}z_{\cdot,t}^{\top}q)}.$
+Similarly, we have
+$\tilde{g}^{\prime}(q_{a})=\sum_{t^{\prime}=1}^{N}\frac{1}{\sqrt{d}}z_{at^{\prime}}\exp{(\frac{1}{\sqrt{d}}z_{\cdot,t^{\prime}}^{\top}q)}.$
+It follows that
+$g^{\prime}(q_{a})\tilde{g}(q_{a})$ $=\frac{1}{\sqrt{d}}z_{at}\exp{(\frac{1}{\sqrt{d}}z_{\cdot,t}^{\top}q)}\sum_{t^{\prime}=1}^{N}\exp{(\frac{1}{\sqrt{d}}z_{\cdot,t^{\prime}}^{\top}q)}$
+$=\frac{1}{\sqrt{d}}z_{at}\sum_{t^{\prime}=1}^{N}\exp{(\frac{1}{\sqrt{d}}\sum_{o=1}^{d}(z_{\cdot,t}+z_{\cdot,t^{\prime}})^{\top}q)}$
+and similarly
+$g(q_{a})\tilde{g}^{\prime}(q_{a})$ $=\exp{(\frac{1}{\sqrt{d}}z_{\cdot,t}^{\top}q)}\sum_{t^{\prime}=1}^{N}\frac{1}{\sqrt{d}}z_{at^{\prime}}\exp{(\frac{1}{\sqrt{d}}z_{\cdot,t^{\prime}}^{\top}q)}$
+$=\frac{1}{\sqrt{d}}\sum_{t^{\prime}=1}^{N}z_{at^{\prime}}\exp{(\frac{1}{\sqrt{d}}\sum_{o=1}^{d}(z_{\cdot,t}+z_{\cdot,t^{\prime}})^{\top}q)}.$
+Putting everything together, we obtain
+$\frac{\partial s_{t}}{\partial q_{a}}$ $=\frac{1}{\sqrt{d}}\frac{\sum_{t^{\prime}=1}^{N}(z_{at}-z_{at^{\prime}})\exp{(\frac{1}{\sqrt{d}}(z_{\cdot,t}+z_{\cdot,t^{\prime}})^{\top}q)}}{[\sum_{r=1}^{N}\exp{(\frac{1}{\sqrt{d}}\sum_{o=1}^{d}z_{or}q_{o})}]^{2}}$
+$=\frac{1}{\sqrt{d}}\sum_{t^{\prime}=1}^{N}(z_{at}-z_{at^{\prime}})\frac{\exp{(\frac{1}{\sqrt{d}}z_{\cdot,t}^{\top}q)}}{\sum_{r=1}^{N}\exp{(\frac{1}{\sqrt{d}}\sum_{o=1}^{d}z_{or}q_{o})}}\cdot\frac{\exp{(\frac{1}{\sqrt{d}}z_{\cdot,t^{\prime}}^{\top}q)}}{\sum_{r=1}^{N}\exp{(\frac{1}{\sqrt{d}}\sum_{o=1}^{d}z_{or}q_{o})}}$
+$=\frac{1}{\sqrt{d}}\sum_{t^{\prime}=1}^{N}(z_{at}-z_{at^{\prime}})\frac{\exp{(\frac{1}{\sqrt{d}}\sum_{o=1}^{d}z_{ot}q_{o})}}{\sum_{r=1}^{N}\exp{(\frac{1}{\sqrt{d}}\sum_{o=1}^{d}z_{or}q_{o})}}\cdot\frac{\exp{(\frac{1}{\sqrt{d}}\sum_{o=1}^{d}z_{ot^{\prime}}q_{o})}}{\sum_{r=1}^{N}\exp{(\frac{1}{\sqrt{d}}\sum_{o=1}^{d}z_{or}q_{o})}}$
+$=\frac{1}{\sqrt{d}}\sum_{t^{\prime}=1}^{N}(z_{at}-z_{at^{\prime}})\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})_{t}\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})_{t^{\prime}}$
+$=\frac{1}{\sqrt{d}}\sum_{t^{\prime}=1}^{N}(z_{at}-z_{at^{\prime}})s_{t}s_{t^{\prime}}.$
+Using Eq. (12) leads to
+$\begin{split}\frac{\partial f_{1}(q_{a})}{\partial q_{a}}=\frac{\partial\xi_{l}}{\partial q_{a}}&=\sum_{t=1}^{N}(Vz)_{lt}\frac{1}{\sqrt{d}}\sum_{t^{\prime}=1}^{N}(z_{at}-z_{at^{\prime}})s_{t}s_{t^{\prime}}\\ &=\frac{1}{\sqrt{d}}\sum_{tt^{\prime}}(Vz)_{lt}z_{at}s_{t}s_{t^{\prime}}-\frac{1}{\sqrt{d}}\sum_{tt^{\prime}}(Vz)_{lt}z_{at^{\prime}}s_{t}s_{t^{\prime}}\\ &=\frac{1}{\sqrt{d}}\sum_{t=1}^{N}(Vz)_{lt}z_{at}s_{t}\underbrace{\sum_{t^{\prime}=1}^{N}s_{t^{\prime}}}_{=1}-\frac{1}{\sqrt{d}}\sum_{t=1}^{N}(Vz)_{lt}s_{t}\underbrace{\sum_{t=1}^{N}z_{at^{\prime}}s_{t^{\prime}}}_{=\mu(z_{a})}\\ &=\frac{1}{\sqrt{d}}\sum_{t=1}^{N}(z_{at}-\mu(z_{a})(Vz)_{lt}s_{t}\\ &=\sum_{t=1}^{N}\Delta_{at}(Vz)_{lt}s_{t},\end{split}$ (14)
+where $\Delta_{at}=\frac{z_{at}-\mu(z_{a})}{\sqrt{d}}$ corresponds to the deviation from the mean embedding at dimension $a$, assuming the distribution probability follows the softmax $(s_{t})_{t=1}^{N}$.
+We note that the $\Delta_{at}$ are the entries of
+$\Delta_{z}=\frac{z}{\sqrt{d}}\left[I_{N}-\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\mathbbm{1}_{N}^{\top}\right]\in{\mathbb{R}}^{d\times N}$
+defined in Eq.
+(4).
+Hence, $\frac{\partial f_{1}(q_{a})}{\partial q_{a}}$ can be understood as the $l$-element of a sentence embedding where the deviation from the mean embedding has rescaled the value matrix.
+We now proceed to the RHS of Eq.
+(11).
+RHS of Eq. (11): a simple application of the chain rule gives
+$\frac{\partial f_{2}^{i}(q_{a})}{\partial q_{a}}=\underbrace{\nabla f_{5}^{i}(q_{a})}_{\in{\mathbb{R}}^{1\times d}}\underbrace{\nabla f_{4}^{i}(\xi)}_{\in{\mathbb{R}}^{d\times d}}\underbrace{\nabla f_{3}^{i}(\bar{\xi})}_{\in{\mathbb{R}}^{d}}\underbrace{\sigma^{\prime}(w_{i}^{\top}\bar{\xi})}_{\in{\mathbb{R}}},$
+where we recall that $\bar{\xi}=\xi/\lVert\xi\rVert$. We have
+$\nabla f_{5}^{i}(q_{a})=\frac{\partial\xi}{\partial q_{a}}=\left[\frac{\partial\xi_{1}}{\partial q_{a}},\ldots,\frac{\partial\xi_{d}}{\partial q_{a}}\right],$
+and
+$\nabla f_{3}^{i}(\bar{\xi})=w_{i}.$
+For $f_{4}^{i}$, we note that $f_{4}^{i}(\xi)=(h_{1}\circ h_{2})(\xi)\cdot\xi$, with $h_{1}\colon{\mathbb{R}}\to{\mathbb{R}},t\mapsto\frac{1}{t}$ and $h_{2}\colon{\mathbb{R}}^{d}\to{\mathbb{R}},x\mapsto\lVert x\rVert$.
+It leads to
+$\nabla f_{4}^{i}(\xi)$ $=\frac{\partial(h_{1}\circ h2)(\xi)}{\partial\xi}\cdot\xi^{\top}+(h_{1}\circ h_{2})(\xi)\cdot\frac{\partial\xi}{\partial\xi}$
+$=\frac{\partial(h_{1}\circ h2)(\xi)}{\partial\xi}\cdot\xi^{\top}+\frac{1}{\lVert\xi\rVert}I_{d}.$
+Since we have $h_{2}(\xi)=\lVert\xi\rVert=\sqrt{\xi^{\top}\xi}$, we know that
+$\frac{\partial h2(\xi)}{\partial\xi}=\frac{1}{2\sqrt{\xi^{\top}\xi}}\cdot 2\xi=\frac{\xi}{\lVert\xi\rVert}.$
+It leads to
+$\frac{\partial(h_{1}\circ h2)(\xi)}{\partial\xi}=-\frac{1}{\lVert\xi\rVert^{2}}\cdot\frac{\xi}{\lVert\xi\rVert}=-\frac{\xi}{\lVert\xi\rVert^{3}}.$
+Putting everything together, we obtain
+$\nabla f_{4}^{i}(\xi)=-\frac{\xi}{\lVert\xi\rVert^{3}}\xi^{\top}+\frac{1}{\lVert\xi\rVert}I_{d}=\frac{1}{\lVert\xi\rVert}\left(I_{d}-\bar{\xi}\bar{\xi}^{\top}\right).$
+Going back to $\frac{\partial f_{2}^{i}(q_{a})}{\partial q_{a}}$, we have
+$\frac{\partial f_{2}^{i}(q_{a})}{\partial q_{a}}=\left[\frac{\partial\xi_{1}}{\partial q_{a}},\ldots,\frac{\partial\xi_{d}}{\partial q_{a}}\right]\frac{1}{\lVert\xi\rVert}\left(I_{d}-\bar{\xi}\bar{\xi}^{\top}\right)w_{i}\sigma^{\prime}(w_{i}^{\top}\bar{\xi})=\lambda Mw_{i}\alpha_{i},$
+where we introduce
+$\lambda=\left[\frac{\partial\xi_{1}}{\partial q_{a}},\ldots,\frac{\partial\xi_{d}}{\partial q_{a}}\right]\in{\mathbb{R}}^{1\times d},\quad M=\frac{1}{\lVert\xi\rVert}\left(I_{d}-\bar{\xi}\bar{\xi}^{\top}\right)\in{\mathbb{R}}^{d\times d},\quad\alpha_{i}=\sigma^{\prime}(w_{i}^{\top}\bar{\xi}),$ (15)
+to ease notations. Developing the matrix multiplication, we obtain
+$\lambda Mw_{i}\alpha_{i}$ $=\sum_{r=1}^{d}\frac{\partial\xi_{r}}{\partial q_{a}}\left[Mw_{i}\right]_{r}\alpha_{i}$
+$=\alpha_{i}\sum_{r=1}^{d}\sum_{t=1}^{N}\Delta_{at}(Vz)_{rt}s_{t}\left[Mw_{i}\right]_{r}$ (using Eq. (14))
+$=\alpha_{i}\sum_{t=1}^{N}\Delta_{at}s_{t}\sum_{r=1}^{d}(Vz)_{rt}\sum_{o=1}^{d}M_{ro}(w_{i})_{o}$
+$=\alpha_{i}\sum_{t=1}^{N}\Delta_{at}s_{t}\sum_{o=1}^{d}(w_{i})_{o}\sum_{r=1}^{d}(Vz)_{rt}M_{ro}$
+$=\alpha_{i}\sum_{t=1}^{N}\Delta_{at}s_{t}\sum_{o=1}^{d}(w_{i})_{o}\sum_{r=1}^{d}M_{ro}(Vz)_{rt}$
+$=\alpha_{i}\sum_{t=1}^{N}\Delta_{at}s_{t}\sum_{o=1}^{d}(w_{i})_{o}\sum_{r=1}^{d}M^{\top}_{or}(Vz)_{rt}$
+$=\alpha_{i}\sum_{t=1}^{N}\Delta_{at}s_{t}\sum_{o=1}^{d}(w_{i})_{o}\left[M^{\top}Vz\right]_{ot}$
+$=\alpha_{i}\sum_{t=1}^{N}\Delta_{at}s_{t}\sum_{o=1}^{d}(w_{i})_{o}\tilde{M}_{ot},$
+where we introduce
+$\tilde{M}=M^{\top}Vz\in{\mathbb{R}}^{d\times N}.$ (16)
+Now multiplying by $(u_{i})_{l}$ and summing over $i\in[h]$, we obtain
+$\sum_{i=1}^{h}(u_{i})_{l}\frac{\partial f_{2}^{i}(q_{a})}{\partial q_{a}}$ $=\sum_{i=1}^{h}(u_{i})_{l}\alpha_{i}\sum_{t=1}^{N}\Delta_{at}s_{t}\sum_{o=1}^{d}(w_{i})_{o}\tilde{M}_{ot}$
+$=\sum_{t=1}^{N}\Delta_{at}s_{t}\sum_{o=1}^{d}\tilde{M}_{ot}\sum_{i=1}^{h}\alpha_{i}(u_{i})_{l}(w_{i})_{o}.$
+We can now put everything together to rewrite Eq. (11).
+LHS + RHS of Eq. (11): We have
+$\frac{\partial\psi_{l}}{\partial q_{a}}=\frac{\partial\xi_{l}}{\partial q_{a}}+\sum_{t=1}^{N}\Delta_{at}s_{t}\sum_{o=1}^{d}\tilde{M}_{ot}\sum_{i=1}^{h}\alpha_{i}(u_{i})_{l}(w_{i})_{o}.$
+and Eq. (9) leads to
+$\frac{\partial\zeta_{j}}{\partial q_{a}}=\underbrace{\sum_{l=1}^{p}(W_{U})_{jl}\frac{\partial\xi_{l}}{\partial q_{a}}}_{=\mathrm{LHS}}+\underbrace{\sum_{l=1}^{p}(W_{U})_{jl}\sum_{t=1}^{N}\Delta_{at}s_{t}\sum_{o=1}^{d}\tilde{M}_{ot}\sum_{i=1}^{h}\alpha_{i}(u_{i})_{l}(w_{i})_{o}}_{=\mathrm{RHS}}.$ (17)
+The LHS can be rewitten as
+$\mathrm{LHS}$ $=\sum_{l=1}^{p}(W_{U})_{jl}\sum_{t=1}^{N}\Delta_{at}(Vz)_{lt}s_{t}$
+$=\sum_{t=1}^{N}\Delta_{at}s_{t}\sum_{l=1}^{p}(W_{U})_{jl}(Vz)_{lt}$
+$=\sum_{t=1}^{N}\Delta_{at}s_{t}\sum_{l=1}^{p}E(j)_{l}(Vz)_{lt}$
+$=\sum_{t=1}^{N}\Delta_{at}s_{t}E(j)^{\top}(Vz)_{\cdot,t}$
+$=\left[\sum_{t=1}^{N}\Delta_{at}s_{t}[(Vz)_{\cdot,t}]^{\top}\right]E(j).$
+The RHS can be rewritten as
+$\mathrm{RHS}$ $=\sum_{l=1}^{p}(W_{U})_{jl}\sum_{t=1}^{N}\Delta_{at}s_{t}\sum_{o=1}^{d}\tilde{M}_{ot}\sum_{i=1}^{h}\alpha_{i}(u_{i})_{l}(w_{i})_{o}$
+$=\sum_{t=1}^{N}\Delta_{at}s_{t}\sum_{l=1}^{p}(W_{U})_{jl}\sum_{i=1}^{h}\alpha_{i}(u_{i})_{l}\sum_{o=1}^{d}\tilde{M}_{ot}(w_{i})_{o}$
+$=\sum_{t=1}^{N}\Delta_{at}s_{t}\sum_{l=1}^{p}(W_{U})_{jl}\sum_{i=1}^{h}\alpha_{i}(u_{i})_{l}\sum_{o=1}^{d}(w_{i})_{o}\tilde{M}_{ot}$
+$=\sum_{t=1}^{N}\Delta_{at}s_{t}\sum_{l=1}^{p}(W_{U})_{jl}\sum_{i=1}^{h}\alpha_{i}(u_{i})_{l}\sum_{o=1}^{d}W_{io}\tilde{M}_{ot}$
+$=\sum_{t=1}^{N}\Delta_{at}s_{t}\sum_{l=1}^{p}(W_{U})_{jl}\sum_{i=1}^{h}\alpha_{i}(u_{i})_{l}(W\tilde{M})_{it}.$
+We now introduce the matrix
+$\tilde{P}=U\begin{pmatrix}\alpha_{1}&0&\cdots\\ \cdots&\cdots&\cdots\\ \cdots&0&\alpha_{h}\end{pmatrix}W\tilde{M}\in{\mathbb{R}}^{d\times N}.$ (18)
+Its entry in row $l$ and column $t$ writes
+$\tilde{P}_{lt}$ $=\left[U\begin{pmatrix}\alpha_{1}&0&\cdots\\ \cdots&\cdots&\cdots\\ \cdots&0&\alpha_{h}\end{pmatrix}W\tilde{M}\right]_{lt}$
+$=\sum_{i=1}^{h}\left[U\begin{pmatrix}\alpha_{1}&0&\cdots\\ \cdots&\cdots&\cdots\\ \cdots&0&\alpha_{h}\end{pmatrix}\right]_{li}(W\tilde{M})_{it}$
+$=\sum_{i=1}^{h}\alpha_{i}U_{li}(W\tilde{M})_{it}$
+$=\sum_{i=1}^{h}\alpha_{i}(u_{i})_{l}(W\tilde{M})_{it}.$ (as $U$ has columns $u_{i}$)
+Plugging this term in the computation of the RHS leads to
+$\mathrm{RHS}$ $=\sum_{t=1}^{N}\Delta_{at}s_{t}\sum_{l=1}^{p}(W_{U})_{jl}\tilde{P}_{lt}$
+$=\sum_{t=1}^{N}\Delta_{at}s_{t}\sum_{l=1}^{p}\tilde{P}_{lt}(W_{U})_{jl}$
+$=\sum_{t=1}^{N}\Delta_{at}s_{t}\sum_{l=1}^{p}\tilde{P}_{lt}E(j)_{l}$ (as $W_{U}$ has rows $E(j)^{\top}$)
+$=\sum_{t=1}^{N}\Delta_{at}s_{t}\sum_{l=1}^{p}\tilde{P}^{\top}_{tl}E(j)_{l}$
+$=\sum_{t=1}^{N}\Delta_{at}s_{t}\tilde{P}^{\top}_{t}E(j).$
+Going back to Eq. (17), we have
+$\frac{\partial\zeta_{j}}{\partial q_{a}}$ $=\left[\sum_{t=1}^{N}\Delta_{at}s_{t}[(Vz)_{\cdot,t}]^{\top}\right]E(j)+\sum_{t=1}^{N}\Delta_{at}s_{t}\tilde{P}^{\top}_{t}E(j)$
+$=\sum_{t=1}^{N}\Delta_{at}s_{t}\left[(Vz)_{\cdot,t}+\tilde{P}^{\top}_{t}\right]E(j)$
+$=\sum_{t=1}^{N}\Delta_{at}s_{t}\left[(Vz)_{\cdot,t}+\tilde{P}_{\cdot,t}\right]E(j)$
+$=\sum_{t=1}^{N}\Delta_{at}s_{t}\left[Vz+\tilde{P}\right]_{\cdot,t}E(j)$
+$=\sum_{t=1}^{N}\Delta_{at}s_{t}\left[Vz+\tilde{P}\right]^{\top}_{t}E(j)$
+$=\sum_{t=1}^{N}\Delta_{at}\left[\begin{pmatrix}s_{1}&0&\cdots\\ \cdots&\cdots&\cdots\\ \cdots&0&s_{N}\end{pmatrix}\left[Vz+\tilde{P}\right]^{\top}\right]_{t}E(j)$
+$=\Delta_{a}\begin{pmatrix}s_{1}&0&\cdots\\ \cdots&\cdots&\cdots\\ \cdots&0&s_{N}\end{pmatrix}\left[Vz+\tilde{P}\right]^{\top}E(j),$
+where $\Delta_{a}$ is the $a$-th column of $\Delta$ defined in Eq. (4). Using Eq. (2) and Eq. (13), we notice that
+$A_{z}=\mathrm{diag}(\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}}))=\mathrm{diag}((s_{t})_{t=1}^{N})=\begin{pmatrix}s_{1}&0&\cdots\\ \cdots&\cdots&\cdots\\ \cdots&0&s_{N}\end{pmatrix}\in{\mathbb{R}}^{N\times N}.$
+Using Eq. (3) and Eq. (15), we notice that
+$\Sigma_{\bar{\xi}}=\mathrm{diag}(\sigma^{\prime}(W\bar{\xi})=\mathrm{diag}((\sigma^{\prime}(w_{i}^{\top}\bar{\xi})_{i=1}^{h})=\mathrm{diag}((\alpha_{i})_{i=1}^{h})=\begin{pmatrix}\alpha_{1}&0&\cdots\\ \cdots&\cdots&\cdots\\ \cdots&0&\alpha_{h}\end{pmatrix}\in{\mathbb{R}}^{N\times N}.$
+Using Eq. (3) and Eq. (18), we thus have
+$\tilde{P}=U\begin{pmatrix}\alpha_{1}&0&\cdots\\ \cdots&\cdots&\cdots\\ \cdots&0&\alpha_{h}\end{pmatrix}W\tilde{M}=U\Sigma_{\bar{\xi}}M\tilde{W}$
+and we obtain
+$Vz+\tilde{P}$ $=Vz+U\Sigma_{\bar{\xi}}M\tilde{W}$
+$=Vz+U\Sigma_{\bar{\xi}}WM^{\top}Vz$ (using Eq. (16))
+$=Vz+U\Sigma_{\bar{\xi}}WMVz$ ($M$ symmetric from Eq. (15))
+$=\left(I_{d}+U\Sigma_{\bar{\xi}}WM\right)Vz.$
+It follows that
+$\frac{\partial\zeta_{j}}{\partial q_{a}}=\Delta_{a}A_{z}\left[\left(I_{d}+U\Sigma_{\bar{\xi}}WM\right)Vz\right]^{\top}E(j).$
+It leads to $\nabla_{q}\zeta_{j}=\left(\frac{\partial\zeta_{j}}{\partial q_{a}}\right)_{a}\in{\mathbb{R}}^{d}$ that writes
+$\nabla_{q}\zeta_{j}=\Delta_{z}A_{z}\left[\left(I_{d}+U\Sigma_{\bar{\xi}}WM\right)Vz\right]^{\top}E(j).$
+The corresponding entries of $\nabla_{\theta}\zeta_{j}$ are the entries of $\nabla_{q}\zeta_{j}$. ∎
+Attention Value.
+The following lemma provides the formula of $\nabla_{V}\xi_{j}$.
+{boxlem}[Value] Let $\Sigma_{\bar{\xi}}=\mathrm{diag}(\sigma^{\prime}(W\bar{\xi}))\in{\mathbb{R}}^{h\times h}$ following Eq.
+(3), $M=\frac{1}{\lVert\xi\rVert}\left(I_{d}-\bar{\xi}\bar{\xi}^{\top}\right)\in{\mathbb{R}}^{d\times d}$ following Eq.
+(15) and $Q=U\Sigma_{\bar{\xi}}WM\in{\mathbb{R}}^{d\times d}$.
+The entries of $\nabla_{\theta}\zeta_{j}$ associated to the value $V$ are the entries of $\nabla_{U}\zeta_{j}\in{\mathbb{R}}^{d\times d}$ flattened in ${\mathbb{R}}^{d^{2}}$ where we have
+$\nabla_{V}\zeta_{j}=\left(I_{d}+Q^{\top}\right)E(j)z\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}}).$
+Proof.
+We want to compute the $\frac{\partial\zeta_{j}}{\partial\theta_{n}}$ for the $\theta_{n}$ that are entries of $V$.
+Using Eq.
+(9) we first need to derive $\frac{\partial\psi_{l}}{\partial\theta_{n}}$.
+As $\theta$ encompasses all the learnable weights, we will derive it depending on the value of $n$.
+Considering $n$ such that $\theta_{n}=V_{bc}$ for some $b,c\in[d]$ and using Eq.
+(10), we have
+$\frac{\partial\psi_{l}}{\partial\theta_{n}}=\frac{\partial\psi_{l}}{\partial V_{bc}}=\frac{\partial\xi_{l}}{\partial V_{bc}}+\frac{\partial}{\partial V_{bc}}(\sum_{i=1}^{h}(u_{i})_{l}\sigma(w_{i}^{\top}\bar{\xi}))=\underbrace{\frac{\partial\beta_{1}(V_{bc})}{\partial V_{bc}}}_{\mathrm{LHS}}+\underbrace{\sum_{i=1}^{h}(u_{i})_{l}\frac{\partial\beta_{2}^{i}(V_{bc})}{\partial V_{bc}}}_{\mathrm{RHS}},$ (19)
+with
+$\begin{cases}\beta_{1}\colon&{\mathbb{R}}\to{\mathbb{R}},\quad V_{bc}\mapsto\xi_{l}\\ \beta_{2}^{i}\colon&{\mathbb{R}}\to{\mathbb{R}},\quad V_{bc}\mapsto(\sigma\circ\beta_{3}^{i}\circ\beta_{4}^{i}\circ\beta_{5}^{i})(V_{bc})=\sigma(w_{i}^{\top}\bar{\xi})\\ \beta_{3}^{i}\colon&{\mathbb{R}}^{d}\to{\mathbb{R}},\quad\bar{\xi}\mapsto w_{i}^{\top}\bar{\xi}\\ \beta_{4}^{i}\colon&{\mathbb{R}}^{d}\to{\mathbb{R}}^{d},\quad\xi\mapsto\bar{\xi}=\frac{\xi}{\lVert\xi\rVert}\\ \beta_{5}^{i}\colon&{\mathbb{R}}\to{\mathbb{R}}^{d},\quad V_{bc}\mapsto\xi.\\ \end{cases}$
+We proceed in two steps and first compute the LHS of Eq. (19).
+LHS of Eq. (19): we first note that
+$\xi_{l}=\sum_{t=1}^{N}(Vz)_{lt}s_{t}=\sum_{t=1}^{N}\sum_{c=1}^{d}V_{lc}z_{ct}s_{t}=\sum_{c=1}^{d}V_{lc}\sum_{t=1}^{N}z_{ct}s_{t},$
+where $s_{t}=\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})_{t}$.
+We notice that $\xi_{l}$ depends only on the $(V_{lc})$ hence the gradients with respect to $V_{bc}$ for some $c\in[d]$ and $b\neq l$ equals zero.
+It follows that
+$\frac{\partial\beta_{1}(V_{bc})}{\partial V_{bc}}=\frac{\partial\xi_{l}}{\partial V_{bc}}=\begin{cases}\sum_{t=1}^{N}z_{ct}s_{t}&\text{ if }b=l\\ 0&\text{ otherwise }\end{cases}.$ (20)
+RHS of Eq. (19): a simple application of the chain rule gives
+$\frac{\partial\beta_{2}^{i}(V_{bc})}{\partial V_{bc}}=\underbrace{\nabla\beta_{5}^{i}(V_{bc})}_{\in{\mathbb{R}}^{1\times d}}\underbrace{\nabla\beta_{4}^{i}(\xi)}_{\in{\mathbb{R}}^{d\times d}}\underbrace{\nabla\beta_{3}^{i}(\bar{\xi})}_{\in{\mathbb{R}}^{d}}\underbrace{\sigma^{\prime}(w_{i}^{\top}\bar{\xi})}_{\in{\mathbb{R}}},$
+where we recall that $\bar{\xi}=\xi/\lVert\xi\rVert$. We have, using similar arguments as the computation of $\frac{\partial\xi_{l}}{\partial V_{bc}}$ in Eq. (20) that
+$\nabla\beta_{5}^{i}(V_{bc})=\frac{\partial\xi}{\partial V_{bc}}=\left[\frac{\partial\xi_{1}}{\partial V_{bc}},\ldots,\frac{\partial\xi_{d}}{\partial V_{bc}}\right]=\left[0,\ldots,0,\sum_{t=1}^{N}z_{ct}s_{t},0,\ldots,0\right],$
+where the only non-zero entry is in position $b$ of the row-vector $\nabla\beta_{5}^{i}(V_{bc})\in{\mathbb{R}}^{1\times d}$.
+We notice that the functions $\beta_{3}^{i}$ and $\beta_{4}^{i}$ of Eq.
+(19) corresponds to the functions $f_{3}^{i}$ and $f_{4}^{i}$ of Eq.
+(11).
+Hence, the gradients are the same and we have
+$\nabla\beta_{3}^{i}(\bar{\xi})=\nabla f_{3}^{i}(\bar{\xi})=w_{i}$
+and
+$\nabla\beta_{4}^{i}(\bar{\xi})=\nabla f_{4}^{i}(\bar{\xi})=\frac{1}{\lVert\xi\rVert}\left(I_{d}-\bar{\xi}\bar{\xi}^{\top}\right).$
+Going back to $\frac{\partial\beta_{2}^{i}(V_{bc})}{\partial V_{bc}}$, we have
+$\frac{\partial\beta_{2}^{i}(V_{bc})}{\partial V_{bc}}=\left[0,\ldots,0,\sum_{t=1}^{N}z_{ct}s_{t},0,\ldots,0\right]\frac{1}{\lVert\xi\rVert}\left(I_{d}-\bar{\xi}\bar{\xi}^{\top}\right)w_{i}\sigma^{\prime}(w_{i}^{\top}\bar{\xi})=\mu Mw_{i}\alpha_{i},$
+where we introduce
+$\mu=\left[0,\ldots,0,\sum_{t=1}^{N}z_{ct}s_{t},0,\ldots,0\right]\in{\mathbb{R}}^{1\times d},\quad M=\frac{1}{\lVert\xi\rVert}\left(I_{d}-\bar{\xi}\bar{\xi}^{\top}\right)\in{\mathbb{R}}^{d\times d},\quad\alpha_{i}=\sigma^{\prime}(w_{i}^{\top}\bar{\xi}),$ (21)
+to ease notations, with $M,\alpha_{i}$ first defined in Eq. (15). Developing the matrix multiplication, we obtain
+$\mu Mw_{i}\alpha_{i}=\alpha_{i}\sum_{t=1}^{N}z_{ct}s_{t}\left[Mw_{i}\right]_{b}=\alpha_{i}\sum_{t=1}^{N}z_{ct}s_{t}\sum_{o=1}^{d}M_{bo}(w_{i})_{o}.$
+It follows that
+$\sum_{i=1}^{h}(u_{i})_{l}\frac{\partial\beta_{2}^{i}(V_{bc})}{\partial V_{bc}}=\sum_{i=1}^{h}(u_{i})_{l}\alpha_{i}\sum_{t=1}^{N}z_{ct}s_{t}\sum_{o=1}^{d}M_{bo}(w_{i})_{o}=\sum_{t=1}^{N}z_{ct}s_{t}\sum_{i=1}^{h}\alpha_{i}(u_{i})_{l}\sum_{o=1}^{d}M_{bo}(w_{i})_{o}.$
+We can now put everything together to rewrite Eq. (19).
+LHS + RHS of Eq. (19): We have
+$\frac{\partial\psi_{l}}{\partial V_{bc}}=\frac{\partial\xi_{l}}{\partial V_{bc}}+\sum_{t=1}^{N}z_{ct}s_{t}\sum_{i=1}^{h}\alpha_{i}(u_{i})_{l}\sum_{o=1}^{d}M_{bo}(w_{i})_{o}$
+and Eq. (9) leads to
+$\frac{\partial\zeta_{j}}{\partial V_{bc}}=\underbrace{\sum_{l=1}^{p}(W_{U})_{jl}\frac{\partial\xi_{l}}{\partial V_{bc}}}_{=\mathrm{LHS}}+\underbrace{\sum_{l=1}^{p}(W_{U})_{jl}\sum_{t=1}^{N}z_{ct}s_{t}\sum_{i=1}^{h}\alpha_{i}(u_{i})_{l}\sum_{o=1}^{d}M_{bo}(w_{i})_{o}}_{=\mathrm{RHS}}.$ (22)
+Using Eq. (20), the LHS can be rewritten as
+$\mathrm{LHS}=(W_{U})_{jb}\sum_{t=1}^{N}z_{ct}s_{t}=E(j)_{b}\sum_{t=1}^{N}z_{ct}s_{t}.$
+The RHS can be rewritten as
+$\mathrm{RHS}$ $=\sum_{l=1}^{p}(W_{U})_{jl}\sum_{t=1}^{N}z_{ct}s_{t}\sum_{i=1}^{h}\alpha_{i}(u_{i})_{l}\sum_{o=1}^{d}M_{bo}(w_{i})_{o}$
+$=\sum_{t=1}^{N}z_{ct}s_{t}\sum_{l=1}^{p}(W_{U})_{jl}\left[\sum_{i=1}^{h}\alpha_{i}(u_{i})_{l}\sum_{o=1}^{d}M_{bo}(w_{i})_{o}\right]$
+$=\sum_{t=1}^{N}z_{ct}s_{t}\sum_{l=1}^{p}(W_{U})_{jl}\left[\sum_{i=1}^{h}\alpha_{i}(u_{i})_{l}\sum_{o=1}^{d}W_{io}M_{bo}\right]$ (as $W$ has rows $w_{i}^{\top}$)
+$=\sum_{t=1}^{N}z_{ct}s_{t}\sum_{l=1}^{p}(W_{U})_{jl}\left[\sum_{i=1}^{h}\alpha_{i}(u_{i})_{l}(WM)_{ib}\right].$
+We now introduce the matrix
+$Q=U\begin{pmatrix}\alpha_{1}&0&\cdots\\ \cdots&\cdots&\cdots\\ \cdots&0&\alpha_{h}\end{pmatrix}WM\in{\mathbb{R}}^{d\times d}.$ (23)
+Its entry in row $l$ and column $b$ writes
+$Q_{lt}$ $=\left[U\begin{pmatrix}\alpha_{1}&0&\cdots\\ \cdots&\cdots&\cdots\\ \cdots&0&\alpha_{h}\end{pmatrix}WM\right]_{lb}$
+$=\sum_{i=1}^{h}\left[U\begin{pmatrix}\alpha_{1}&0&\cdots\\ \cdots&\cdots&\cdots\\ \cdots&0&\alpha_{h}\end{pmatrix}\right]_{li}(WM)_{ib}$
+$=\sum_{i=1}^{h}\alpha_{i}U_{li}(WM)_{ib}$
+$=\sum_{i=1}^{h}\alpha_{i}(u_{i})_{l}(WM)_{ib}^{\top}.$ (as $W$ has rows $w_{i}$)
+Plugging this term in the computation of the RHS leads to
+$\mathrm{RHS}$ $=\sum_{t=1}^{N}z_{ct}s_{t}\sum_{l=1}^{p}(W_{U})_{jl}Q_{lb}$
+$=\sum_{t=1}^{N}z_{ct}s_{t}\sum_{l=1}^{p}E(j)_{l}Q_{lb}$
+$=\sum_{t=1}^{N}z_{ct}s_{t}\sum_{l=1}^{p}Q^{\top}_{bl}E(j)_{l}$
+$=\sum_{t=1}^{N}z_{ct}s_{t}\left[Q^{\top}E(j)\right]_{b}$
+Going back to Eq. (22), we have
+$\frac{\partial\zeta_{j}}{\partial V_{bc}}$ $=E(j)_{b}\sum_{t=1}^{N}z_{ct}s_{t}+\sum_{t=1}^{N}z_{ct}s_{t}\left[Q^{\top}E(j)\right]_{b}$
+$=\left[E(j)+Q^{\top}E(j)\right]_{b}\sum_{t=1}^{N}z_{ct}s_{t}$
+$=\left[\left(I_{d}+Q^{\top}\right)E(j)\right]_{b}\sum_{t=1}^{N}z_{ct}s_{t}.$
+We introduce the vector $z\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\in{\mathbb{R}}^{d}$ whose entries writes
+$\left[z\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\right]_{c}=\sum_{t=1}^{N}z_{ct}\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})_{t}=\sum_{t=1}^{N}z_{ct}s_{t},$
+where we used the fact that $s_{t}=\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})_{t}$ following Eq. (13). It follows that
+$\frac{\partial\zeta_{j}}{\partial V_{bc}}=\left[\left(I_{d}+Q^{\top}\right)E(j)\right]_{b}\left[z\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\right]_{c}.$
+Using Eq. (3) and Eq (23), we have
+$Q=U\begin{pmatrix}\alpha_{1}&0&\cdots\\ \cdots&\cdots&\cdots\\ \cdots&0&\alpha_{h}\end{pmatrix}WM=U\Sigma_{\bar{\xi}}WM.$
+Hence, $\nabla_{V}\zeta_{j}=\left(\frac{\partial\zeta_{j}}{\partial V_{bc}}\right)_{b,c}\in{\mathbb{R}}^{d\times d}$ is a matrix equal to
+$\nabla_{V}\zeta_{j}=\left(I_{d}+\left[U\Sigma_{\bar{\xi}}WM\right]^{\top}\right)E(j)\left[z\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\right]^{\top}.$
+The corresponding entries of $\nabla_{\theta}\zeta_{j}$ amounts to the entries of $\nabla_{V}\zeta_{j}$ flattened in ${\mathbb{R}}^{d^{2}}$. This concludes the proof. ∎
+Feed-forward receptors.
+The following lemma provides the formula of $\nabla_{W}\xi_{j}$.
+{boxlem}[Receptors] Let $\Sigma_{\bar{\xi}}=\mathrm{diag}(\sigma^{\prime}(W\bar{\xi}))\in{\mathbb{R}}^{h\times h}$ following Eq.
+(3).
+The entries of $\nabla_{\theta}\zeta_{j}$ associated to the assemblers $(w_{i})_{i=1}^{h}$ are the entries of $\nabla_{W}\zeta_{j}\in{\mathbb{R}}^{htimesd}$ flattened in ${\mathbb{R}}^{dh}$ where we have
+$\nabla_{W}\zeta_{j}=\Sigma_{\bar{\xi}}U^{\top}E(j)\bar{\xi}^{\top}.$
+Proof.
+We want to compute the $\frac{\partial\zeta_{j}}{\partial\theta_{n}}$ for the $\theta_{n}$ that are entries of $W$.
+Using Eq.
+(9) we first need to derive $\frac{\partial\psi_{l}}{\partial\theta_{n}}$.
+As $\theta$ encompasses all the learnable weights, we will derive it depending on the value of $n$.
+Here, $\psi_{l}$ depends on all the $(w_{i})_{e}$.
+Considering $n$ such that $\theta_{n}=(w_{i})_{e}$ for some $i\in[h]$ and some $e\in[d]$ and using Eq.
+(10), we have
+$\frac{\partial\psi_{l}}{\partial\theta_{n}}=\frac{\partial\psi_{l}}{\partial(w_{i})_{e}}=(u_{i})_{l}\frac{\partial\sigma(w_{i}^{\top}\bar{\xi})}{\partial(w_{i})_{e}}.$
+Noting that $w_{i}^{\top}\xi=\sum_{f=1}^{d}(w_{i})_{f}\xi_{f}$ and using the chain rule, we have
+$\frac{\partial\sigma(w_{i}^{\top}\xi/\lVert\xi\rVert)}{\partial(w_{i})_{e}}$ $=\frac{\partial(w_{i}^{\top}\bar{\xi})}{\partial(w_{i})_{e}}\sigma^{\prime}(w_{i}^{\top}\bar{\xi})$
+$=\frac{\partial(\sum_{f=1}^{d}(w_{i})_{f}\bar{\xi}_{f})}{\partial(w_{i})_{e}}\sigma^{\prime}(w_{i}^{\top}\bar{\xi})$
+$=\frac{\partial((w_{i})_{e}\bar{\xi}_{e})}{\partial(w_{i})_{e}}\sigma^{\prime}(w_{i}^{\top}\bar{\xi})$
+$=\bar{\xi}_{e}\sigma^{\prime}(w_{i}^{\top}\bar{\xi}).$
+We recall that $\sigma$ is the GELU activation defined as $\sigma\colon x\to x\varphi(x)$ with $\varphi$ the cumulative distribution function of the standard normal distribution.
+By the fundamental theorem of analysis, it leads that
+$\sigma^{\prime}(x)=\varphi(x)+x\varphi^{\prime}(x)=\varphi(x)+xf(x),$
+with $f\colon u\to\frac{1}{\sqrt{2\pi}}\exp{(-\frac{u^{2}}{2})}$ the probability density function of the standard normal distribution. Putting everything together, we have
+$\frac{\partial\psi_{l}}{\partial(w_{i})_{e}}=(u_{i})_{l}\bar{\xi}_{e}\sigma^{\prime}(w_{i}^{\top}\bar{\xi}).$
+Using Eq. (9), we have
+$\frac{\partial\zeta_{j}}{\partial(w_{i})_{e}}$ $=\sum_{l=1}^{p}(W_{U})_{jl}(u_{i})_{l}\bar{\xi}_{e}\sigma^{\prime}(w_{i}^{\top}\bar{\xi})$
+$=\sum_{l=1}^{p}E(j)_{l}(u_{i})_{l}\bar{\xi}_{e}\sigma^{\prime}(w_{i}^{\top}\bar{\xi})$
+$=E(j)^{\top}u_{i}\bar{\xi}_{e}\sigma^{\prime}(w_{i}^{\top}\xi/\lVert\xi\rVert)$
+$=\sigma^{\prime}(w_{i}^{\top}\bar{\xi})u_{i}^{\top}E(j)\bar{\xi}_{e}.$
+To ease the notations, we denote $\alpha_{i}=\sigma^{\prime}(w_{i}^{\top}\bar{\xi})$.
+Recalling that $\sigma$ and thus $\sigma^{\prime}$ is entry-wise, it leads to $\nabla_{W}\zeta_{j}=(\frac{\partial\zeta_{j}}{\partial(w_{i})_{e}})_{i,e}\in{\mathbb{R}}^{h\times d}$ that writes
+$\nabla_{W}\zeta_{j}$ $=\begin{pmatrix}\alpha_{1}u_{1}^{\top}E(j)\\ \cdots\\ \alpha_{h}u_{h}^{\top}E(j)\\ \end{pmatrix}\bar{\xi}^{\top}$
+$=\begin{pmatrix}\alpha_{1}u_{1}^{\top}\\ \cdots\\ \alpha_{h}u_{h}^{\top}\end{pmatrix}E(j)\bar{\xi}^{\top}$
+$=\begin{pmatrix}\alpha_{1}&0&\cdots\\ \cdots&\cdots&\cdots\\ \cdots&0&\alpha_{h}\end{pmatrix}U^{\top}E(j)\bar{\xi}^{\top}$
+$=\begin{pmatrix}\sigma^{\prime}(w_{1}^{\top}\bar{\xi})&0&\cdots\\ \cdots&\cdots&\cdots\\ \cdots&0&\sigma^{\prime}(w_{h}^{\top}\bar{\xi})\end{pmatrix}U^{\top}E(j)\bar{\xi}^{\top}$
+$=\mathrm{diag}(\sigma^{\prime}(W\bar{\xi}))U^{\top}E(j)\bar{\xi}^{\top},$
+where $\mathrm{diag}(\sigma^{\prime}(W\bar{\xi}))\in{\mathbb{R}}^{h\times h}$ is diagonal with entries $(\sigma^{\prime}(w_{i}^{\top}\bar{\xi}))_{i=1}^{h}$. Using Eq. (3), it finally leads to
+$\nabla_{W}\zeta_{j}=\Sigma_{\bar{\xi}}U^{\top}E(j)\bar{\xi}^{\top},$
+and the corresponding entries of $\nabla_{\theta}\zeta_{j}$ amounts to the entries of $\nabla_{W}\zeta_{j}$ flattened in ${\mathbb{R}}^{hd}$. This concludes the proof. ∎
+Feed-forward assemblers.
+The following lemma provides the formula of $\nabla_{U}\xi_{j}$.
+{boxlem}[Assemblers] The entries of $\nabla_{\theta}\zeta_{j}$ associated to the assemblers $(u_{i})_{i=1}^{h}$ are the entries of $\nabla_{U}\zeta_{j}\in{\mathbb{R}}^{d\times h}$ flattened in ${\mathbb{R}}^{dh}$ where we have
+$\nabla_{U}\zeta_{j}=E(j)\sigma(W\bar{\xi})^{\top}\in R^{d\times h}.$
+Proof.
+We want to compute the $\frac{\partial\zeta_{j}}{\partial\theta_{n}}$ for the $\theta_{n}$ that are entries of $U$.
+Using Eq.
+(9) we first need to derive $\frac{\partial\psi_{l}}{\partial\theta_{n}}$.
+As $\theta$ encompasses all the learnable weights, we will derive it depending on the value of $n$.
+As $\psi_{l}$ only depends on the $l$-elements of the $(u_{i})$, we only consider indices $n$ such that $\theta_{n}=(u_{i})_{l}$ for some $i\in[h]$ as the other values of $n$ will lead to a null gradient.
+Using Eq.
+(10), we have
+$\frac{\partial\psi_{l}}{\partial\theta_{n}}=\frac{\partial\psi_{l}}{\partial(u_{i})_{l}}=\frac{\partial(u_{i})_{l}}{\partial(u_{i})_{l}}\sigma(w_{i}^{\top}\bar{\xi})=\sigma(w_{i}^{\top}\bar{\xi}).$
+Using Eq. (9), we have
+$\frac{\partial\zeta_{j}}{\partial(u_{i})_{l}}=(W_{U})_{jl}\sigma(w_{i}^{\top}\bar{\xi})=E(j)_{l}\sigma(w_{i}^{\top}\bar{\xi}).$
+Hence, recalling that $\sigma$ is applied element-wise, $\nabla_{U}\zeta_{j}=(\frac{\partial\zeta_{j}}{\partial(u_{i})_{l}})_{l,i}\in{\mathbb{R}}^{d\times h}$ is a matrix equal to
+$\nabla_{U}\zeta_{j}=E(j)\sigma(W\bar{\xi})^{\top},$
+and the corresponding entries of $\nabla_{\theta}\zeta_{j}$ amounts to the entries of $\nabla_{U}\zeta_{j}$ flattened in ${\mathbb{R}}^{dh}$. This concludes the proof. ∎
+We can now put everything together. Using Section D.1 in combination with the gradientsderived in Section D.1.1, Section D.1.1, Section D.1.1 and Section D.1.1 concludes the proof of Eq. 5. ∎
+D.2 Proof of C.2
+We detail below the proof of C.2.
+Proof.
+We assume the gradient descent continues until convergence at a critical point and discard other practical considerations such as early stopping.
+Following Chi et al.
+(2019b) and using Section C.3, we know that the parameters $\theta$ of the transformer verify
+$\nabla_{\theta}\mathcal{L}(\theta)=0$ (24)
+From Eq. 5, we know that $\nabla_{\theta}\mathcal{L}(\theta)\in{\mathbb{R}}^{m}$ writes
+$\nabla_{\theta}\mathcal{L}(\theta)=[\nabla_{q}\mathcal{L}(\theta),\nabla_{V}\mathcal{L}(\theta),\nabla_{W}\mathcal{L}(\theta),\nabla_{U}\mathcal{L}(\theta)],$ (25)
+using []delimited-[][$\cdot$][ ⋅ ] as the vector concatenation and flattening matrices whenever needed. We now reformulate the gradients of each weight group. We have
+$\nabla_{q}\mathcal{L}(\theta)$ $=\sum_{j=1}^{p}\mathbb{E}_{x}\left[(\hat{p}_{\theta}(y=j|x)-p(y=j|x))\cdot\nabla_{q}\zeta_{j}\right]$
+$=\sum_{j=1}^{p}\mathbb{E}_{x}\left[(\hat{p}_{\theta}(y=j|x)-p(y=j|x))\cdot\left(\Delta_{z}A_{z}(Vz)^{\top}\left(I_{d}+Q^{\top}\right)E(j)\right)\right]$
+$=\mathbb{E}_{x}\left[\sum_{j=1}^{p}(\hat{p}_{\theta}(y=j|x)-p(y=j|x))\cdot\left(\Delta_{z}A_{z}(Vz)^{\top}\left(I_{d}+Q^{\top}\right)E(j)\right)\right]$ (by linearity of the expectation)
+$=\mathbb{E}_{x}\left[\Delta_{z}A_{z}(Vz)^{\top}\left(I_{d}+Q^{\top}\right)\cdot\sum_{j=1}^{p}(\hat{p}_{\theta}(y=j|x)-p(y=j|x))E(j)\right]$ ($\hat{p}_{\theta}(y=j|x)-p(y=j|x)\in{\mathbb{R}}$)
+$=\mathbb{E}_{x}\left[\Delta_{z}A_{z}(Vz)^{\top}\left(I_{d}+Q^{\top}\right)\mathscr{C}_{x}\right].$ (by definition of $\mathscr{C}_{x}$)
+The same arguments lead to
+$\nabla_{V}\mathcal{L}(\theta)$ $=\sum_{j=1}^{p}\mathbb{E}_{x}\left[(\hat{p}_{\theta}(y=j|x)-p(y=j|x))\cdot\nabla_{V}\zeta_{j}\right]$
+$=\sum_{j=1}^{p}\mathbb{E}_{x}\left[(\hat{p}_{\theta}(y=j|x)-p(y=j|x))\cdot\left(\left(I_{d}+Q^{\top}\right)E(j)\left[z\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\right]^{\top}\right)\right]$
+$=\mathbb{E}_{x}\left[\sum_{j=1}^{p}(\hat{p}_{\theta}(y=j|x)-p(y=j|x))\cdot\left(\left(I_{d}+Q^{\top}\right)E(j)\left[z\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\right]^{\top}\right)\right]$
+$=\mathbb{E}_{x}\left[\left(I_{d}+Q^{\top}\right)\cdot\left(\sum_{j=1}^{p}(\hat{p}_{\theta}(y=j|x)-p(y=j|x))E(j)\right)\cdot\left[z\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\right]^{\top}\right]$
+$=\mathbb{E}_{x}\left[\left(I_{d}+Q^{\top}\right)\mathscr{C}_{x}\left[z\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\right]^{\top}\right],$
+then to
+$\nabla_{W}\mathcal{L}(\theta)$ $=\sum_{j=1}^{p}\mathbb{E}_{x}\left[(\hat{p}_{\theta}(y=j|x)-p(y=j|x))\cdot\nabla_{W}\zeta_{j}\right]$
+$=\sum_{j=1}^{p}\mathbb{E}_{x}\left[(\hat{p}_{\theta}(y=j|x)-p(y=j|x))\cdot\left(\Sigma_{\bar{\xi}}U^{\top}E(j)\bar{\xi}^{\top}\right)\right]$
+$=\mathbb{E}_{x}\left[\sum_{j=1}^{p}(\hat{p}_{\theta}(y=j|x)-p(y=j|x))\cdot\left(\Sigma_{\bar{\xi}}U^{\top}E(j)\bar{\xi}^{\top}\right)\right]$
+$=\mathbb{E}_{x}\left[\Sigma_{\bar{\xi}}U^{\top}\cdot\left(\sum_{j=1}^{p}(\hat{p}_{\theta}(y=j|x)-p(y=j|x))E(j)\right)\cdot\bar{\xi}^{\top}\right]$
+$=\mathbb{E}_{x}\left[\Sigma_{\bar{\xi}}U^{\top}\mathscr{C}_{x}\bar{\xi}^{\top}\right],$
+and finally to
+$\nabla_{U}\mathcal{L}(\theta)$ $=\sum_{j=1}^{p}\mathbb{E}_{x}\left[(\hat{p}_{\theta}(y=j|x)-p(y=j|x))\cdot\nabla_{U}\zeta_{j}\right]$
+$=\sum_{j=1}^{p}\mathbb{E}_{x}\left[(\hat{p}_{\theta}(y=j|x)-p(y=j|x))\cdot\left(E(j)\sigma(W\bar{\xi})^{\top}\right)\right]$
+$=\mathbb{E}_{x}\left[\sum_{j=1}^{p}(\hat{p}_{\theta}(y=j|x)-p(y=j|x))\cdot\left(E(j)\sigma(W\bar{\xi})^{\top}\right)\right]$
+$=\mathbb{E}_{x}\left[\left(\sum_{j=1}^{p}(\hat{p}_{\theta}(y=j|x)-p(y=j|x))E(j)\right)\cdot\sigma(W\bar{\xi})^{\top}\right]$
+$=\mathbb{E}_{x}\left[\mathscr{C}_{x}\sigma(W\bar{\xi})^{\top}\right].$
+The equality to 00 of the vector in ${\mathbb{R}}^{m}$ is equivalent to each of its components being equal to 00, and in particular, this is the case of the gradients with respect to $q,V,W$, and $U$.
+Combining Eq.
+24 with Eq.
+5 leads to the following stationarity conditions
+$\begin{cases}&\mathbb{E}_{x}\left[\Delta_{z}A_{z}(Vz)^{\top}\left(I_{d}+Q^{\top}\right)\mathscr{C}_{x}\right]=0\\ &\mathbb{E}_{x}\left[\left(I_{d}+Q^{\top}\right)\mathscr{C}_{x}\left[z\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\right]^{\top}\right]=0\\ &\mathbb{E}_{x}\left[\Sigma_{\bar{\xi}}U^{\top}\mathscr{C}_{x}\bar{\xi}^{\top}\right]=0\\ &\mathbb{E}_{x}\left[\mathscr{C}_{x}\sigma(W\bar{\xi})^{\top}\right]=0.\end{cases}$
+Once those conditions are established, the rest of the proof is straightforward.
+Indeed, it suffices to recall that the clustering heads described in Section 3.1 capture all the invariants of the problem.
+This leads to a perfect separation of the clusters that can be seen in Fig.
+1.
+In particular, the transformer correctly classifies all the sequences, which amounts to the $\hat{p}_{\theta}(y=j|x)$ and the $p(y=j|x)$ being aligned.
+By definition, this implies $\mathscr{C}_{x}=0$ for all $x$, and noticing the linear dependency of the gradients with $\mathscr{C}_{x}$ ensures that the stationary conditions previously stated are verified.
+Hence, a transformer that implements a circuit head has parameters $\theta$ for which $\nabla\mathcal{L}(\theta)=0$ holds and that can thus be reached via gradient descent.
+This implies circuit heads can emerge during the training of the transformer.
+This concludes the proof.
+∎
+D.3 Proof of Section 4.4
+We detail below the proof of Section 4.4 whose formal statement is given in Section C.4.
+Proof.
+We first decompose the gradient of the loss using Eq. 25. Gathering the terms by group leads to
+$\lVert\nabla_{\theta}\mathcal{L}(\theta)\rVert^{2}=\sum_{n=1}^{m}\left|\frac{\partial\mathcal{L}(\theta)}{\partial\theta_{n}}\right|^{2}=\lVert\nabla_{q}\mathcal{L}(\theta)\rVert^{2}+\lVert\nabla_{V}\mathcal{L}(\theta)\rVert^{2}+\lVert\nabla_{W}\mathcal{L}(\theta)\rVert^{2}+\lVert\nabla_{U}\mathcal{L}(\theta)\rVert^{2},$ (26)
+where the norm is the Euclidean one with matrices flattened as vectors. Similarly to the proof of C.2, we have
+$\begin{cases}&\nabla_{q}\mathcal{L}(\theta)=\mathbb{E}_{x}\left[\Delta_{z}A_{z}(Vz)^{\top}\left(I_{d}+Q^{\top}\right)\mathscr{C}_{x}\right]\\ &\nabla_{V}\mathcal{L}(\theta)=\mathbb{E}_{x}\left[\left(I_{d}+Q^{\top}\right)\mathscr{C}_{x}\left[z\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\right]^{\top}\right]\\ &\nabla_{W}\mathcal{L}(\theta)=\mathbb{E}_{x}\left[\Sigma_{\bar{\xi}}U^{\top}\mathscr{C}_{x}\bar{\xi}^{\top}\right]\\ &\nabla_{U}\mathcal{L}(\theta)=\mathbb{E}_{x}\left[\mathscr{C}_{x}\sigma(W\bar{\xi})^{\top}\right].\end{cases}$
+We now proceed with upper-bounding each term of Eq.
+26.
+For clarity of the proofs, we first recall the following basic relations.
+{boxlem}[Properties] Let $a\in{\mathbb{R}}^{m},b\in{\mathbb{R}}^{n},A\in{\mathbb{R}}^{n\times m}$ and $B\in{\mathbb{R}}^{m\times l}$.
+We have
+  1. 1.
+    $\lVert ab^{\top}\rVert=\lVert a\rVert\cdot\lVert b\rVert$.
+  2. 2.
+    $\lVert Ab\rVert\leq\lVert A\rVert_{\mathrm{op}}\lVert b\rVert$.
+  3. 3.
+    The operator norm (resp. the Frobenius norm) is sub-multiplicative i.e., $\lVert AB\rVert_{\mathrm{op}}\leq\lVert A\rVert_{\mathrm{op}}\lVert B\rVert_{\mathrm{op}}$.
+  4. 4.
+    If $A$ has elements bounded by $C>0$, then $\lVert A\rVert_{\mathrm{op}}\leq C\sqrt{nm}$.
+  5. 5.
+    If $P$ be an orthogonal projection, then, we have $\lVert P\rVert_{\mathrm{op}}=1$.
+  6. 6.
+    The singular values of $A$ and $A^{\top}$ coincides. In particular, $\lVert A\rVert_{\mathrm{op}}=\lVert A^{\top}\rVert_{\mathrm{op}}$.
+Proof.
+The first point follows from
+$\lVert ab^{\top}\rVert^{2}=\sum_{i,j}(ab^{\top})_{ij}^{2}=\sum_{i,j}(a_{i}b_{j})^{2}=\left(\sum_{i}a_{i}^{2}\right)\cdot\left(\sum_{j}b_{j}^{2}\right)=\lVert a\rVert^{2}\cdot\lVert b\rVert^{2},$
+where taking the square root gives the desired equality.
+The second point comes from the definition of the operator norm
+$\lVert A\rVert_{\mathrm{op}}=\sup\{\frac{\lVert Ax\rVert}{\lVert x\rVert}:x\in{\mathbb{R}}^{m},x\neq 0\}=\sup\{\lVert Ax\rVert:x\in{\mathbb{R}}^{m},\lVert x\rVert=1\}.$
+It implies that for any $x\in{\mathbb{R}}^{m}$ non-zero, by definition of the sup, we have
+$\lVert Ax\rVert=\frac{\lVert Ax\rVert}{\lVert x\rVert}\cdot\lVert x\rVert\leq\lVert A\rVert_{\mathrm{op}}\lVert x\rVert.$
+The inequality holds (with equality) for $x=0$.
+The third point is simply a property of matrix norms.
+The fourth point comes from the definition of operator norm. Let $x\in{\mathbb{R}}^{m}$ with unit norm. We have
+$\lVert Ax\rVert^{2}=\sum_{i=1}^{n}\left(\sum_{j=1}^{m}A_{ij}x_{j}\right)^{2}\leq\sum_{i=1}^{n}\left(\sum_{j=1}^{m}A_{ij}^{2}\right)\left(\sum_{j=1}^{m}x_{j}^{2}\right)\leq nm\cdot C^{2},$ (27)
+using the fact that $\lVert x\rVert^{2}=\sum_{j=1}^{m}x_{j}^{2}=1$ and the bound on the entries of $A$ for the last inequality.
+Taking the root square and the sup on the left-hand term gives the desired formula.
+For the fifth point, we recall that an orthogonal projection is symmetric.
+Hence, we have $P=P^{2}=p^{\top}P$ which also ensures $P$ to be semi-definite positive, i.e., with nonnegative eigenvalues.
+Let $\lambda$ be such an eigenvalue and $v$ an eigenvector.
+We must have $Pv=\lambda v$ but applying $P$ on this relation also leads to
+$Pv=P^{2}v=P(Pv)=P(\lambda v)=\lambda^{2}v.$
+Taking the norm forces $\lambda=\lambda^{2}$ which implies $\lambda\in\{0,1\}$.
+As the singular values of $P$ are its eigenvalues (by the symmetry of $P$), it implies that $\lVert P\rVert_{\mathrm{op}}\leq 1$.
+Moreover, writing the singular value decomposition of $P$ leads to $P=\sum_{i}\sigma_{i}u_{i}v_{i}^{\top}$.
+Let $u_{k}$ be in the range of $P$ (it suffices to show that $Pv_{k}=\sigma_{k}u_{k}$ and since $P\neq 0$, there is at least one $\sigma_{k}$ non-zero which provides at least one $u_{k}$ in the range of $P$).
+Using $P^{2}=P$ implies that
+$Pu_{k}=u_{k}\iff\sum_{i}\sigma_{i}u_{i}v_{i}^{\top}u_{k}=u_{k}\iff\sum_{i}(\sigma_{i}v_{i}^{\top}u_{k})u_{i}=u_{k}.$
+Since the $(u_{i})$ forms an orthogonal basis, it forces to have $(\sigma_{k}v_{k}^{\top}u_{k})=1$.
+A simple application of Cauchy-Schwartz (recalling that the $u_{i},v_{i}$ have unit norm) leads to $v_{k}^{\top}u_{k}\leq 1$, and finally it implies that $\sigma_{k}\geq 1$.
+This forces by definition of the operator norm $\lVert P\rVert_{\mathrm{op}}\geq 1$.
+Putting everything together gives $\lVert P\rVert_{\mathrm{op}}=1$.
+For the last point, assume $A$ has a singular value decomposition that writes
+$A=\sum_{i=1}^{\min\{n,m\}}\sigma_{i}u_{i}v_{i}^{\top},$
+where the $u_{i}\in{\mathbb{R}}^{n},v_{i}\in{\mathbb{R}}^{m}$ are orthogonal basis. Hence, we have
+$A^{\top}=\sum_{i=1}^{\min\{n,m\}}\sigma_{i}v_{i}u_{i}^{\top},$
+which shows that the singular values are the same. Since the operator norm computes the largest singular value, we have the last equality. ∎
+In the rest, we upper-bound all the terms by extensively using the monotonicity of the expectation.
+Query term.
+We have
+$\lVert\nabla_{q}\mathcal{L}(\theta)\rVert^{2}$ $=\lVert\mathbb{E}_{x}\left[\Delta_{z}A_{z}(Vz)^{\top}\left(I_{d}+Q^{\top}\right)\mathscr{C}_{x}\right](\theta)\rVert^{2}$
+$\leq\mathbb{E}_{x}\left[\lVert\Delta_{z}A_{z}(Vz)^{\top}\left(I_{d}+Q^{\top}\right)\mathscr{C}_{x}\rVert^{2}\right]$ (Jensen inequality since $\lVert\cdot\rVert^{2}$ is convex)
+$\leq\mathbb{E}_{x}\left[\lVert\Delta_{z}A_{z}(Vz)^{\top}\left(I_{d}+Q^{\top}\right)\rVert_{\mathrm{op}}^{2}\lVert\mathscr{C}_{x}\rVert^{2}\right]$ (Point $(2)$ of Section D.3)
+Using the fact that operator norms are sub-multiplicative, we have
+$\lVert\Delta_{z}A_{z}(Vz)^{\top}\left(I_{d}+Q^{\top}\right)\rVert_{\mathrm{op}}\leq\lVert\Delta_{z}\rVert_{\mathrm{op}}\lVert A_{z}\rVert_{\mathrm{op}}\lVert(Vz)^{\top}\rVert_{\mathrm{op}}\lVert\left(I_{d}+Q^{\top}\right)\rVert_{\mathrm{op}}.$
+Recalling from Eq. 4 that
+$\Delta_{z}=\frac{z}{\sqrt{d}}\left[I_{N}-\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\mathbbm{1}_{N}^{\top}\right],$
+and using the fact that the embeddings $z\in{\mathbb{R}}^{d\times M}$ have columns in $\mathbb{S}^{d}$ (thanks to the RMS-norm), we have that $I_{N}-\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\mathbbm{1}_{N}^{\top}$ has entries bounded by $2$.
+Indeed, given that
+$\forall t\in[N],\lVert z_{\cdot,t}\rVert^{2}=\sum_{i=1}^{d}z_{dt}^{2}=1,$
+we know that the $z_{dt}$ are bounded by one.
+As the softmax outputs a probability vector, the $z_{dt}$ are reweighted by elements in $[0,1]$.
+This implies that the entries of $\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\mathbbm{1}_{N}^{\top}$ (a matrix with rows equal to $\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})$) are bounded by 1.
+Taking the minus sign into account leads to $I_{N}-\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\mathbbm{1}_{N}^{\top}$ with entries bounded by $2$ and using the point $(3)$ and $(4)$ of Section D.3, we have
+$\lVert\Delta_{z}\rVert_{\mathrm{op}}$ $\leq\frac{1}{\sqrt{d}}\lVert z\rVert_{\mathrm{op}}\lVert I_{N}-\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\mathbbm{1}_{N}^{\top}\rVert_{\mathrm{op}}$
+$\leq\frac{1}{\sqrt{d}}\lVert z^{\top}\rVert_{\mathrm{op}}\cdot(2N)$ (last point of Section D.3)
+Moreover, using a proof similar to the point $(4)$ of Section D.3 (in particular involving Eq. 27), one can show that for any $A\in{\mathbb{R}}^{n\times m}$,
+$\lVert A\rVert_{\mathrm{op}}\leq\sqrt{m}\max_{i\in[m]}\lVert A_{i}\rVert.$
+Applying the previous inequality to $z^{\top}\in{\mathbb{R}}^{N\times d}$ that has rows $z_{\cdot,t}$ with unit norm leads to
+$\lVert z^{\top}\rVert_{\mathrm{op}}\leq\sqrt{d}.$ (28)
+We thus obtain $\lVert\Delta_{z}\rVert_{\mathrm{op}}\leq 2N$. Recalling from (2) that
+$A_{z}=\mathrm{diag}\left(\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\right)$
+is diagonal with entries that form a probability vector, we know that
+$\lVert A_{z}\rVert_{\mathrm{op}}=\sigma_{\mathrm{max}}(A_{z})=\lambda_{\mathrm{max}}(A_{z})=\max_{t\in[N]}\{\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})_{t}\}\leq 1.$
+Then, using Eq. 28, we notice that
+$\lVert(Vz)^{\top}\rVert_{\mathrm{op}}=\lVert z^{\top}V^{\top}\rVert_{\mathrm{op}}\leq\lVert z^{\top}\rVert_{\mathrm{op}}\lVert V^{\top}\rVert_{\mathrm{op}}\leq\sqrt{d}\lVert V^{\top}\rVert_{\mathrm{op}}=\sqrt{d}\lVert V\rVert_{\mathrm{op}},$
+where the last equality comes from the last point of Section D.3.
+We now proceed to $I+Q^{\top}$.
+Using the triangular inequality leads to $\lVert I+Q^{\top}\rVert_{\mathrm{op}}\leq\lVert I\rVert_{\mathrm{op}}+\lVert Q^{\top}\rVert_{\mathrm{op}}$.
+We know that $\lVert I\rVert_{\mathrm{op}}=1$.
+Moreover, we have
+$\lVert Q^{\top}\rVert_{\mathrm{op}}=\lVert M(U\Sigma_{\bar{\xi}}W)^{\top}\rVert_{\mathrm{op}}\leq\lVert M\rVert_{\mathrm{op}}\cdot\lVert(U\Sigma_{\bar{\xi}}W)^{\top}\rVert_{\mathrm{op}}.$
+Using the fact that $M$ is an orthogonal projection ensures from the point $(6)$ of Section D.3 that $\lVert M\rVert_{\mathrm{op}}=1$.
+Moreover, using the points $(3)$ and $(6)$ of Section D.3 leads to
+$\lVert(U\Sigma_{\bar{\xi}}W)^{\top}\rVert_{\mathrm{op}}=\lVert U\Sigma_{\bar{\xi}}W\rVert_{\mathrm{op}}\leq\lVert U\rVert_{\mathrm{op}}\lVert\Sigma_{\bar{\xi}}\rVert_{\mathrm{op}}\lVert W\rVert_{\mathrm{op}}\lVert.$
+Noting that $\Sigma_{\bar{\xi}}$ is diagonal with entries $\sigma^{\prime}(w_{i}^{\top}\bar{\xi})$ (Eq.
+3) and recalling that $\sigma^{\prime}(x)=\psi(x)+xf(x)$ with $\varphi$ the cumulative distribution function of the normal distribution and $f$ its probability density function (in particular, $\varphi$ has values in $[0,1]$ and $f$ in $[0,1/\sqrt{2\pi}]$), we know that $\lvert\sigma^{\prime}(x)\rvert\leq 1+(1/\sqrt{2\pi})\lvert x\rvert$.
+Since $\sigma^{\prime}$ is applied entry-wise, using again the point (2) of Section D.3 leads to
+$\lVert\Sigma_{\bar{\xi}}\rVert_{\mathrm{op}}\leq 1+\frac{1}{\sqrt{2\pi}}\lVert W\bar{\xi}\rVert\leq 1+\frac{1}{\sqrt{2\pi}}\lVert W\rVert_{\mathrm{op}}.$
+Putting everything together leads to
+$\lVert\Delta_{z}A_{z}(Vz)^{\top}\left(I_{d}+Q^{\top}\right)\rVert_{\mathrm{op}}\leq 2n\sqrt{d}\lVert V\rVert_{\mathrm{op}}(1+\sqrt{\frac{2}{\pi}}\lVert W\rVert_{\mathrm{op}}^{2}\lVert U\rVert_{\mathrm{op}}).$
+Since the weight matrices are randomly initialized and learned by gradient descent whose updates have the form $\mathbb{E}_{x}\left[\varphi(\mathrm{parameters})\right]$, they do not depend on the expectation with respect to the training data distribution.
+We can thus take those terms outside of the expectation which leads to
+$\lVert\nabla_{q}\mathcal{L}(\theta)\rVert^{2}\leq\left[2n\sqrt{d}\lVert V\rVert_{\mathrm{op}}(1+\sqrt{\frac{2}{\pi}}\lVert W\rVert_{\mathrm{op}}^{2}\lVert U\rVert_{\mathrm{op}})\right]^{2}\cdot\mathbb{E}_{x}\left[\lVert\mathscr{C}_{x}\rVert^{2}\right].$ (29)
+Value term.
+We have
+$\lVert\nabla_{V}\mathcal{L}(\theta)\rVert^{2}$ $\leq\mathbb{E}_{x}\left[\lVert\underbrace{\left(I_{d}+Q^{\top}\right)\mathscr{C}_{x}}_{\in{\mathbb{R}}^{d}}\underbrace{\left[z\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\right]^{\top}}_{\in{\mathbb{R}}^{1\times d}}\rVert^{2}\right]$ (Jensen inequality)
+$\leq\mathbb{E}_{x}\left[\lVert\left(I_{d}+Q^{\top}\right)\mathscr{C}_{x}\rVert^{2}\left\lVert\left[z\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\right]^{\top}\right\rVert^{2}\right]$ (Point (1) of Section D.3)
+Again using the point (2) of Section D.3 and the derivations done for the query term, we have
+$\lVert\left(I_{d}+Q^{\top}\right)\mathscr{C}_{x}\rVert\leq\lVert\left(I_{d}+Q^{\top}\right)\rVert_{\mathrm{op}}\cdot\lVert\mathscr{C}_{x}\rVert\leq(1+2\lVert W\rVert_{\mathrm{op}}\lVert U\rVert_{\mathrm{op}})\cdot\lVert\mathscr{C}_{x}\rVert.$
+Similarly, we have
+$\left\lVert\left[z\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\right]^{\top}\right\rVert=\left\lVert z\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\right\rVert\leq\lVert z\rVert_{\mathrm{op}}\lVert\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\rVert.$
+As shown in the proof of the query term, we have $\lVert z\rVert_{\mathrm{op}}=\lVert z^{\top}\rVert_{\mathrm{op}}\leq\sqrt{d}$ and $p=\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})$ is a probability vector which implies
+$\lVert p\rVert^{2}=\sum_{i=1}^{N}p_{i}^{2}\leq\sum_{i=1}^{N}p_{i}=1,$
+where the first inequality comes from $x\mapsto x^{2}\leq x\mapsto x$ on $[0,1]$. It follows that
+$\left\lVert\left[z\operatorname{softmax}(\frac{z^{\top}q}{\sqrt{d}})\right]^{\top}\right\rVert\leq\sqrt{d}.$
+In summary, with the same arguments as before, we can take the weight terms outside of the expectation and we obtain
+$\lVert\nabla_{V}\mathcal{L}(\theta)\rVert^{2}\leq d(1+2\lVert W\rVert_{\mathrm{op}}\lVert U\rVert_{\mathrm{op}})^{2}\cdot\mathbb{E}_{x}\left[\lVert\mathscr{C}_{x}\rVert^{2}\right].$ (30)
+Receptors terms.
+With similar arguments, we have
+$\lVert\nabla_{W}\mathcal{L}(\theta)\rVert^{2}$ $=\lVert\mathbb{E}_{x}\left[\Sigma_{\bar{\xi}}U^{\top}\mathscr{C}_{x}\bar{\xi}^{\top}\right]\rVert^{2}$
+$\leq\mathbb{E}_{x}\left[\lVert\underbrace{\Sigma_{\bar{\xi}}U^{\top}\mathscr{C}_{x}}_{\in{\mathbb{R}}^{d}}\underbrace{\bar{\xi}^{\top}}_{\in{\mathbb{R}}^{1\times d}}\rVert^{2}\right]$ (Jensen)
+$\leq\mathbb{E}_{x}\left[\lVert\Sigma_{\bar{\xi}}U^{\top}\mathscr{C}_{x}\rVert^{2}\lVert\bar{\xi}^{\top}\rVert^{2}\right]$ (Point (1) of Section D.3)
+By definition, $\bar{\xi}=\xi/\lVert\xi\rVert$ has unit norm. Moreover, using the point (2) of Section D.3, we know that
+$\lVert\Sigma_{\bar{\xi}}U^{\top}\mathscr{C}_{x}\rVert\leq\lVert\Sigma_{\bar{\xi}}U^{\top}\rVert_{\mathrm{op}}\cdot\lVert\mathscr{C}_{x}\rVert\leq\lVert\Sigma_{\bar{\xi}}\rVert_{\mathrm{op}}\cdot\lVert U^{\top}\rVert_{\mathrm{op}}\cdot\lVert\mathscr{C}_{x}\rVert.$
+where the last inequality comes from the sub-multiplicity of the operator norm.
+As we showed in the deviation of the query term, we have $\lVert\Sigma_{\bar{\xi}}\rVert_{\mathrm{op}}\leq 2$ and we know that $\lVert U^{\top}\rVert_{\mathrm{op}}=\lVert U\rVert_{\mathrm{op}}$.
+In summary, taking the weight terms outside of the expectation, we have
+$\lVert\nabla_{W}\mathcal{L}(\theta)\rVert^{2}\leq 4\lVert U\rVert_{\mathrm{op}}^{2}\mathbb{E}_{x}\left[\lVert\mathscr{C}_{x}\rVert^{2}\right].$ (31)
+Assemblers terms.
+The same arguments as before lead to
+$\lVert\nabla_{U}\mathcal{L}(\theta)\rVert^{2}=\lVert\mathbb{E}_{x}\left[\mathscr{C}_{x}\sigma(W\bar{\xi})^{\top}\right]\rVert^{2}\leq\mathbb{E}_{x}\left[\lVert\mathscr{C}_{x}\sigma(W\bar{\xi})^{\top}\rVert^{2}\right]\leq\mathbb{E}_{x}\left[\lVert\mathscr{C}_{x}\rVert^{2}\cdot\lVert\sigma(W\bar{\xi})^{\top}\rVert^{2}\right]$
+where the last inequality comes from the first point of Section D.3.
+Recalling that the GELU function is defined as $x\leq x\varphi(x)$ with $\varphi$ the cumulative distribution function of the standard normal distribution (in particular, $\varphi$ outputs values in [0, 1]) we know that for any real number $x$, $\lvert\sigma(x)\rvert\leq\lvert x\rvert$.
+Using again the point (2) of Section D.3 and that $\bar{\xi}$ has unit norm, since $\sigma$ is applied entry-wise, we have
+$\lVert\sigma(W\bar{\xi})^{\top}\rVert=\lVert\sigma(W\bar{\xi})\rVert\leq\lVert W\bar{\xi}\rVert\leq\lVert W\rVert_{\mathrm{op}}.$
+In summary, taking the weight terms outside of the expectation, we have
+$\lVert\nabla_{U}\mathcal{L}(\theta)\rVert^{2}\leq\lVert W\rVert_{\mathrm{op}}^{2}\mathbb{E}_{x}\left[\lVert\mathscr{C}_{x}\rVert^{2}\right].$ (32)
+Conclusion.
+Putting (29), (30), (31) and (32) together and using (26) leads to
+$\lVert\nabla_{\theta}\mathcal{L}(\theta)\rVert^{2}\leq\left[\left[2n\sqrt{d}\lVert V\rVert_{\mathrm{op}}(1+\sqrt{\frac{2}{\pi}}\lVert W\rVert_{\mathrm{op}}^{2}\lVert U\rVert_{\mathrm{op}})\right]^{2}+d(1+2\lVert W\rVert_{\mathrm{op}}\lVert U\rVert_{\mathrm{op}})^{2}+4\lVert U\rVert_{\mathrm{op}}^{2}+\lVert W\rVert_{\mathrm{op}}^{2}\right]\cdot\mathbb{E}_{x}\left[\lVert\mathscr{C}_{x}\rVert^{2}\right].$ (33)
+Finally, denoting $\gamma_{j}=\hat{p}_{\theta}(y=j|x)-p(y=j|x)$ to ease notations, we notice that
+$\lVert\mathscr{C}_{x}\rVert^{2}$ $=\left\lVert\sum_{j=1}^{p}(\hat{p}_{\theta}(y=j|x)-p(y=j|x))E(j)\right\rVert^{2}$
+$=\left\lVert\sum_{j=1}^{p}\gamma_{j}E(j)\right\rVert^{2}$
+$=\sum_{l=1}^{d}\left(\sum_{j=1}^{p}\gamma_{j}E(j)_{l}\right)^{2}$
+$\leq\sum_{l=1}^{d}\left(\sum_{j=1}^{p}\lvert\gamma_{j}\rvert\lvert E(j)_{l}\rvert\right)^{2}$
+$\leq\left(\sum_{l=1}^{d}\max_{j\in[p]}\lvert E(j)_{l}\rvert^{2}\right)\left(\sum_{j=1}^{p}\lvert\gamma_{j}\rvert\right)^{2}$
+$\leq dB^{2}\left(\sum_{j=1}^{p}\lvert\gamma_{j}\rvert\right)^{2}.$ (using $\lVert E\rVert_{\infty}\leq B$)
+Recalling that the total variation distance verifies for any probability vectors $p,q$
+$d_{\mathrm{TV}}\left(p,q\right)=\frac{1}{2}\lVert p-q\rVert_{1}=\frac{1}{2}\sum_{\omega\in\Omega}\lvert p(\omega)-q(\omega)\rvert,$
+with $\lVert\cdot\rVert_{1}$ the $\ell_{1}$-norm. We notice that
+$\sum_{j=1}^{p}\lvert\gamma_{j}\rvert=\sum_{j=1}^{p}\lvert\hat{p}_{\theta}(y=j|x)-p(y=j|x)\rvert=\lVert\hat{p}_{\theta}(\cdot|x)-p(\cdot|x)\rVert_{1}=2\cdot d_{\mathrm{TV}}\left(\hat{p}_{\theta}(\cdot|x),p(\cdot|x)\right).$
+Recalling that the total variation distance takes values in $[0,1]$ and that $x\mapsto x^{2}\leq x\mapsto x$ on this interval, we know that $d_{\mathrm{TV}}\left(\hat{p}_{\theta}(\cdot|x),p(\cdot|x)\right)^{2}\leq d_{\mathrm{TV}}\left(\hat{p}_{\theta}(\cdot|x),p(\cdot|x)\right)$.
+It leads to
+$\mathbb{E}_{x}\left[\lVert\mathscr{C}_{x}\rVert^{2}\right]\leq 4dB^{2}\mathbb{E}_{x}\left[d_{\mathrm{TV}}\left(\hat{p}_{\theta}(\cdot|x),p(\cdot|x)\right)\right].$
+Rearranging the terms in front of the expectation in (33) finally leads to
+$\lVert\nabla_{\theta}\mathcal{L}(\theta)\rVert^{2}\leq\tilde{B}\cdot\mathbb{E}_{x}\left[d_{\mathrm{TV}}\left(\hat{p}_{\theta}(\cdot|x),p(\cdot|x)\right)\right]$
+where $\tilde{B}=4dB^{2}\left[\left[2n\sqrt{d}\lVert V\rVert_{\mathrm{op}}(1+\sqrt{\frac{2}{\pi}}\lVert W\rVert_{\mathrm{op}}^{2}\lVert U\rVert_{\mathrm{op}})\right]^{2}+d(1+2\lVert W\rVert_{\mathrm{op}}\lVert U\rVert_{\mathrm{op}})^{2}+4\lVert U\rVert_{\mathrm{op}}^{2}+\lVert W\rVert_{\mathrm{op}}^{2}\right]$.
+Noticing that
+$\tilde{B}=\mathcal{O}\left(\lVert V\rVert_{\mathrm{op}}\left[1+\lVert U\rVert_{\mathrm{op}}\lVert W\rVert_{\mathrm{op}}^{2}\right]^{2}+\left[1+\lVert U\rVert_{\mathrm{op}}\lVert W\rVert_{\mathrm{op}}\right]^{2}+\lVert U\rVert_{\mathrm{op}}^{2}+\lVert W\rVert_{\mathrm{op}}^{2}\right)$
+and taking the square root concludes the proof. ∎